### v1

# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

Abstract
Probabilistic electricity price forecasting (PEPF) is a key task for market participants in short-term electricity markets. The increasing availability of high-frequency data and the need for real-time decision-making in energy markets require online estimation methods for efficient model updating. We present an online, multivariate, regularized distributional regression model, allowing for the modeling of all distribution parameters conditional on explanatory variables. Our approach is based on the combination of the multivariate distributional regression and an efficient online learning algorithm based on online coordinate descent for LASSO-type regularization. Additionally, we propose to regularize the estimation along a path of increasingly complex dependence structures of the multivariate distribution, allowing for parsimonious estimation and early stopping. We validate our approach through one of the first forecasting studies focusing on multivariate probabilistic forecasting in the German day-ahead electricity market while using only online estimation methods. We compare our approach to online LASSO-ARX-models with adaptive marginal distribution and to online univariate distributional models combined with an adaptive Copula. We show that the multivariate distributional regression, which allows modeling all distribution parameters – including the mean and the dependence structure – conditional on explanatory variables such as renewable in-feed or past prices provide superior forecasting performance compared to modeling of the marginals only and keeping a static/unconditional dependence structure. Additionally, online estimation yields a speed-up by a factor of 80 to over 400 times compared to batch fitting.
Keywords: online learning, GAMLSS, LASSO, covariance estimation, cholesky-decomposition, low-rank approximation, multivariate distributional regression, probabilistic electricity price forecasting, day-ahead electricity market, EPF
1 Introduction
Short-term electricity markets play a key role in the integration of renewable energy sources and flexible generation in the electricity system. In Germany, the day-ahead auction is the major venue for physically delivered electricity. Trading volumes have grown with the increase of renewable generation capacity. To optimize decision-making and bidding strategies, market participants need accurate price forecasts. Additionally, electricity prices are multivariate time series characterized by high volatility, positive and negative spikes and skewness. Therefore, research and industry have moved towards probabilistic electricity price forecasting (PEPF) to account for their stochastic nature (see e.g. Nowotarski and Weron, 2018; Dexter Energy, 2024). However, the multivariate dimension of electricity price time series has received little attention for PEPF so far, while being of high importance for market participants in the context of the optimization of flexible assets and portfolio management (Löhndorf and Wozabal, 2023; Peña et al., 2024; Beykirch et al., 2022, 2024). At the same time, the increasing availability of high-frequency data and the need for real-time decision-making in energy markets require online estimation methods for efficient model updating. This work presents an online, multivariate distributional regression model, which we apply for probabilistic day-ahead electricity price forecasting in Germany. Our work is among the first to treat the 24-dimensional hourly electricity prices as multivariate distribution and the first to treat the problem in a strict online estimation setting, which makes the complex, high-dimensional distributional learning problem feasible on standard laptops. Our results show that modeling the dependence structure improves forecasting performance significantly compared to univariate approaches.
Figure 1: Correlation Matrix for day-ahead electricity prices $P_{d,h}$ in Germany. The lower triangle gives the Pearson hourly correlation $\rho$ for electricity prices. The upper triangle gives the hourly correlation of residuals $\varepsilon_{d,h}=P_{d,h}-\hat{\mu}_{d,h}$ for a standard LASSO-ARX model (see e.g. Nowotarski and Weron, 2018, and Eq. 18). The high degree of residual correlation, especially around the noon hours is clearly visible. All correlation coefficients are statistically significant to the $\alpha=0.01$ confidence level.
The need for multivariate PEPF
The literature on PEPF has evaluated a wide range of different statistical and machine learning methods, such as quantile regression, ARX-GARCH models (Nowotarski and Weron, 2018; Billé et al., 2023; Marcjasz et al., 2023), conformal prediction methods, (see e.g. Kath and Ziel, 2021; Zaffran et al., 2022; Brusaferri et al., 2024a; Lipiecki et al., 2024), distributional regression and neural network approaches (e.g. Brusaferri et al., 2024b; Marcjasz et al., 2023; Hirsch et al., 2024; Ziel et al., 2021). However, these works treat each delivery hour as independent, univariate time series as in Ziel and Weron (2018). Let us motivate the need for multivariate probabilistic forecasting approaches for the day-ahead electricity price by two simple plots. Figure 2 shows a time series plot for the 24 hourly day-ahead electricity prices in Germany. The left panel shows each delivery hour as individual, daily series, emphasizing the daily co-movement. The right panel shows the cross-section, i.e. the daily shape for the first 180 days of 2017. The temporal correlations along the dimension of the delivery hours $h=0,...,23$ is clearly visible. Additionally, Figure 1 shows the correlation matrix of the raw electricity prices, but also the residual correlation for standard LASSO-ARX models for the electricity price. We see a strong, statistically significant remaining residual cross-correlation, indicating that the resulting marginal error distributions, which are conditional on the mean, are not independent. On top of the statistical motivation, Beykirch et al. (2022, 2024) clearly describe the need for predicting joint distributions for the optimization of schedules and bidding curves in energy markets, further examples are provided by Peña et al. (2024); Löhndorf and Wozabal (2023).
Figure 2: Time series plot for day-ahead electricity prices $P_{d,h}$ in Germany. In the left panel, each color corresponds to one delivery hour $h=0,...,23$. The blue dotted line marks the split between test and training data set. The gray area corresponds to the time of the right panel, which shows the same data along the dimension of the delivery hour, where each line represents a delivery day $d$. The high volatility, occasional positive and negative spikes and co-movement of electricity prices are visible.
Literature on multivariate PEPF
Work on multivariate probabilistic forecasting for day-ahead electricity prices are sparse in the literature and the majority of the existing works, e.g. Maciejowska and Nitka (2024); Berrisch and Ziel (2024); Han (2023); Mashlakov et al. (2021) and Agakishiev et al. (2025), does not evaluate multivariate scoring rules such as the VS, DSS or ES, but focuses on the evaluation of the marginals of the multivariate distribution through the CRPS. This reduces the problem to modeling 24 marginal distributions, taking only lagged cross-information into account. To the best knowledge of the author, only two studies truly model and evaluate the multivariate dependence structure. First, Janke and Steinke (2020) approach the issue through implicit generative Copula models. Grothe et al. (2023) employ the Schaake shuffle, a post-processing method for point forecasts. On the contrary, in the fields of probabilistic weather, renewable production (Bjerregård et al., 2021; Sørensen et al., 2022; Kolkmann et al., 2024) and probabilistic load forecasting (Gioia et al., 2022; Browell et al., 2022) truly multivariate forecasting approaches have gained more attention.
Distributional Regression
The goal of distributional regression or “regression beyond the mean” (Kneib et al., 2023; Klein, 2024) is modeling not only the conditional expectation, but all distribution parameters of the assumed parametric response distribution conditional on explanatory variables. The most prominent model in this regard is the original GAMLSS (Generalized Additive Model for Location, Scale and Shape Rigby and Stasinopoulos, 2005), of which numerous extensions have been developed over the last years (Kock and Klein, 2023; Kneib et al., 2023; Muschinski et al., 2022) and distributional deep neural networks (DDNN, e.g. Klein et al., 2021, 2023; Rügamer et al., 2024). Through the direct modeling of the variable’s distribution, this method is well suited for the generation of probabilistic forecasts and has been successfully applied in energy markets (Muniain and Ziel, 2020; Gioia et al., 2022; Serinaldi, 2011; Brusaferri et al., 2024b; Marcjasz et al., 2023).
Online Learning
For environments with large amounts of continuously incoming data, such as energy markets, online learning describes the task of updating the model given new data, without falling back on previous samples. Formally, in the strict online setting, after having seen $N$ samples of our data set, we fit a model, predict for step $N+1$. Subsequently, we receive the realized values for $N+1$ and update our model, taking into account only the new row $N+1$. This approach allows an efficient processing of high-velocity data and results in greatly decreased computational effort. The principle is outlined as well in Figure 6. Online learning for LASSO-regularized regression for the mean has been introduced in Angelosante et al. (2009, 2010) and Messner and Pinson (2019). Univariate approaches suitable for probabilistic forecasting based stochastic gradient descent have been developed for specific distributions, (see e.g. Pierrot and Pinson, 2021), conformal prediction (see e.g. Zaffran et al., 2022; Brusaferri et al., 2024a) and the generic online distributional regression in Hirsch et al. (2024), however, in the multivariate case, the literature remains sparse and focused on unconditional distributions and Copulae (see e.g. Dasgupta and Hsu, 2007; Zhao et al., 2022; Landgrebe et al., 2020).
Contributions
We add to the literature by presenting a generic, online, regularized, multivariate distributional regression model, allowing to model all distribution parameters conditional on explanatory variables and validate the approach in a forecasting study for the day-ahead electricity market in Germany. Our paper is the first to tackle the issue of truly multivariate probabilistic energy in a strict online estimation setting. In detail, our contributions include
  • We present the online, multivariate distributional regression model based on the combination of the multivariate distributional regression (Muschinski et al., 2022; Kock and Klein, 2023; Gioia et al., 2022) and implement an efficient online learning algorithm based on the univariate work by Hirsch et al. (2024).
  • We propose a regularized estimation, using both, LASSO for each individual distribution parameter, but also a path-based estimation along increasingly complex dependence structures in the multivariate distribution, allowing for parsimonious estimation and early stopping.
  • Our case study explores the multivariate normal and multivariate $t$-distribution for the joint distribution of spot electricity prices, compared to both online LASSO-ARX-models (see e.g. Nowotarski and Weron, 2018) with constant marginal distributions, but also to the online univariate distributional model by Hirsch et al. (2024) combined with a Copula approach. We show that the multivariate distributional regression, which allows modeling all distributional parameters, i.e. the mean, but also the dependence structure, conditional on explanatory variables such as renewable in-feed or past prices provide superior forecasting performance compared to modeling of the marginals only respectively keeping a static/unconditional dependence structure.
  • We providing a high-performing Python implementation using just-in-time compilation and providing a familiar, scikit-learn-like API to facilitate the usage of our package for other researchers. Reproduction code can be found in the GitHub repository at https://github.com/simon-hirsch/online-mv-distreg. We plan to integrate our code in the ROLCH package by Hirsch et al. (2024).
Structure of the paper
The remainder of the paper is structured as usual: The following main Section 2 introduces the multivariate, online, regularized distributional regression model. Sections 3 introduce the forecasting study and the used data and 4 presents our results. Section 5 concludes the paper.
2 Online Multivariate Distribution Regression
We start the exposition on distributional regression by building from the univariate, batch case onto the multivariate setting and subsequently moving to the online setting. The following section introduces the necessary notation, briefly reviews online learning approaches for LASSO-regularized estimation present and the framework of distributional regression in a rather general way. Subsequently, we discuss possible options for achieving (almost) unconstrained parametrization for the scale and precision matrix of multivariate distributions (Section 2.2). Iteratively reweighed least squares for the estimation is introduced in Section 2.3 and the online algorithm is introduced in Section 2.4. Lastly, we discuss path-based regularization and early stopping in Section 2.5.
2.1 Preliminaries and Setting
Notation
We denote scalar float and integer values as lowercase letters (e.g. $a$), constants as large letters (e.g. $T$) vectors as bold, upright lower case letters (e.g. $\bm{\mathrm{v}}$) and matrices as bold upper case letters (e.g. $\bm{\mathrm{A}}$). The calligraphic $\mathcal{F}$ and $\mathcal{D}$ are reserved for (arbitrary) distributions, $\mathcal{N}$ denotes the normal distribution and $\mathcal{L}$ denotes the likelihood; other calligraphic letters (usually) denote index sets. Subscript values are usually indices in matrices, which we start with 0. Superscript indices (in square brackets) denote iterations and/or the number of samples received in the online setting.
Online Coordinate Descent for Regularized Linear Regression
Coordinate descent is the state-of-the-art method to estimate sparse and regularized regression problems of the form
$\bm{\mathrm{\beta}}=\arg\min_{\bm{\mathrm{\beta}}}\left\{\left\|\bm{\mathrm{y}}-\bm{\mathrm{X}}\bm{\mathrm{\beta}}\right\|_{2}+\lambda\|\bm{\mathrm{\beta}}\|_{1}\right\}$
where $\bm{\mathrm{X}}$ is the $N\times J$ design matrix, $\bm{\mathrm{y}}$ is the response variable, $\bm{\mathrm{\beta}}$ is the coefficient vector to be estimated and $\lambda$ is a parameter defining the strength of the regularization. Larger values of $\lambda$ lead to higher regularization. Angelosante et al. (2009, 2010) show that the problem can be reformulated using the Gramian matrices $\bm{\mathrm{G}}=\bm{\mathrm{X}}^{\top}\bm{\mathrm{W}}\bm{\mathrm{\Gamma}}\bm{\mathrm{X}}$ and $\bm{\mathrm{H}}=\bm{\mathrm{X}}^{\top}\bm{\mathrm{W}}\bm{\mathrm{\Gamma}}\bm{\mathrm{y}}$, potentially also accounting for weights $\bm{\mathrm{W}}=\operatorname{diag}(w_{1},...,w_{N})$ and exponential discounting $\bm{\mathrm{\Gamma}}=\operatorname{diag}((1-\gamma)^{N-1},....,(1-\gamma)^{1},(1-\gamma)^{0})$, where $\gamma\in(0,1)$ is a forget parameter. The LASSO problem can be solved by iteratively cycling through all coordinates $j\in J$ and solving
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left(\bm{\mathrm{H}}\left[j\right]-\bm{\mathrm{G}}\left[j,:\right]\bm{\mathrm{\beta}}+\bm{\mathrm{G}}\left[j,j\right]\hat{\beta}_{j},\lambda\right)}{\bm{\mathrm{G}}\left[j,j\right]}.$ (1)
where $S(x,\lambda)=\operatorname{sign}(x)\max(x-\lambda)$ is the so-called soft-threshold function. Coordinate descent is commonly solved on a decreasing grid of regularization strengths $\bm{\mathrm{\lambda}}$ on an exponential grid from $\lambda_{\text{max}}=\max{}\lvert\bm{\mathrm{G}}_{n+1}\rvert$. Algorithm 1 presents the full fitting process. A more detailed treatment of online coordinate descent can be found in Messner and Pinson (2019); Hirsch et al. (2024).
Input: New observations $\bm{\mathrm{x}}^{[n+1]},y^{[n+1]},w^{[n+1]}$ and stored $\bm{\mathrm{G}}^{[n]},\bm{\mathrm{H}}^{[n]}$.
1 Update $\bm{\mathrm{G}}^{[n+1]}=(1-\gamma)\bm{\mathrm{G}}^{[n]}+w_{n+1}(\bm{\mathrm{x}}^{[n+1]})^{\top}\bm{\mathrm{x}}^{[n+1]}$
2 Update $\bm{\mathrm{H}}^{[n+1]}=(1-\gamma)\bm{\mathrm{H}}^{[n]}+w_{n+1}(\bm{\mathrm{x}}^{[n+1]})^{\top}\bm{\mathrm{y}}^{[n+1]}$
3 Update $\lambda_{\text{max}}=\max{}\lvert\bm{\mathrm{G}}_{n+1}\rvert$ and initialize $\bm{\mathrm{\lambda}}$ as exponential grid.
4 for $\lambda\in\bm{\mathrm{\lambda}}$ do
5       Set starting coefficients $\bm{\mathrm{\beta}}_{\lambda}\leftarrow\bm{\mathrm{\beta}}_{\lambda[-1]}$
6       while not converged do
7             forall $j\in 1,...,J$ do
8                   Update $\hat{\beta}_{j,\lambda}$ according to Equation 1
9
10            Check convergence for $\widehat{\bm{\mathrm{\beta}}}_{n+1,\lambda}$ and proceed to next $\lambda$ if converged.
11
12
Output: $\widehat{\bm{\mathrm{\beta}}}_{n+1}=\left(\hat{\beta}_{j,\lambda},...\right)^{\top}$ for all $\lambda\in\bm{\mathrm{\lambda}}$
Algorithm 1 Online LASSO, see Angelosante et al. (2010) and Messner and Pinson (2019)
Univariate Distributional Regression Setting
Distributional regression aims to model the conditional distribution parameters of the univariate response vector $\bm{\mathrm{y}}=(y_{1},...,y_{N})\in\mathbb{R}^{N\times 1}$, conditional on the covariate or explanatory data in $\bm{\mathrm{X}}\in\mathbb{R}^{N\times J}$ by adopting a parametric distribution $\bm{\mathrm{y}}\sim\mathcal{F}(\Theta)$, where $\mathbf{\Theta}=(\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K})$ is a tuple of $K$ distribution parameters $\bm{\mathrm{\theta}}_{k}=(\theta_{k1},...,\theta_{kN})$. Each of the distribution parameters are linked to the covariate data through a known, twice differentiable link function $g_{k}(\cdot)$, leading to:
$g_{k}(\bm{\mathrm{\theta}}_{k})=\bm{\mathrm{\eta}}_{k}={\bm{\mathrm{X}}_{k}}\bm{\mathrm{\beta}}_{k}$ (2)
where $\bm{\mathrm{\beta}}_{k}$ is the coefficient vector to be estimated, relating the $J_{k}$ covariates in the design matrix $\bm{\mathrm{X}}_{k}=(\bm{\mathrm{x}}_{k1},...,\bm{\mathrm{x}}_{kJ})^{\top}$ to the distribution parameter $\bm{\mathrm{\theta}}_{k}$ through the link function $g_{k}(\cdot)$. Hence, we have:
$y_{i}\sim\mathcal{F}(\theta_{1i},...,\theta_{Ki})\quad\text{and}\quad\theta_{ki}=g^{-1}\left(\beta_{k}\bm{\mathrm{x}}_{ki}\right)$ (3)
and the probability density function $f(y_{i}\mid\theta_{1i},...,\theta_{Ki})$. The distributional regression framework therefore allows the modeling of all distribution parameters as linear regression equations of the design matrices $\bm{\mathrm{X}}_{k}$, which can a subset or all of the available the covariate data $\bm{\mathrm{X}}$. Commonly additive models are employed, where $\bm{\mathrm{\eta}}_{k}=f_{k1}(\bm{\mathrm{x}}_{k1})+...+f_{kJ}(\bm{\mathrm{x}}_{kJ})$ where the functions $f_{kj}(\cdot)$ can be linear terms, but also non-linear effects such as B-splines (Klein, 2024; Stasinopoulos et al., 2024). Note that while the functions $f_{kj}(\cdot)$ might be non-linear, they can be represented by a combination of linear regression coefficients and $B$ basis functions $b(\cdot)$, i.e. $f_{kj}(\cdot)=\sum_{i=1}^{B}\beta_{kji}b_{i}(x_{kj})$. Rigby and Stasinopoulos (2005) introduce iteratively reweighted least squares (IRLS), maximizing the penalized likelihood, to estimate $\bm{\mathrm{\beta}}_{k}$. It is important to note here that in the frequentist estimation, the IRLS algorithm is agnostic to the actual estimation technique (see e.g. p. 113 in Stasinopoulos et al., 2024). Different flavors of LASSO-type regularized estimation approaches have been introduced by Groll et al. (2019); Muniain and Ziel (2020); Ziel et al. (2021); O’Neill and Burke (2023). A regularized, incremental estimation approach using online coordinate descent has been proposed by Hirsch et al. (2024), which will form the basis for the multivariate approach proposed in this paper.
Multivariate Distributional Regression Setting
Moving to the multivariate setting, we are interested in learning the conditional distribution parameters of the $D$-dimensional response variable $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{1},...,\bm{\mathrm{y}}_{D})$, conditional on the covariate data $\bm{\mathrm{X}}$, by adopting a multivariate parametric distribution $\bm{\mathrm{Y}}_{i}\sim\mathcal{F}_{D}(\mathbf{\Theta}_{i})$, where $\mathbf{\Theta}_{i}=(\bm{\mathrm{\theta}}_{i1},...,\bm{\mathrm{\theta}}_{iK})$ is a tuple of $K$ scalar, vector or matrix-valued distribution parameters. Each of the coordinates $m$ of the distribution parameter $\bm{\mathrm{\theta}}_{k}$ can again be related to its linear predictors by
$g_{km}(\bm{\mathrm{\theta}}_{km})=\bm{\mathrm{\eta}}_{km}=\bm{\mathrm{X}}_{km}\bm{\mathrm{\beta}}_{km}.$ (4)
Let us appreciate here that this formulation is rather general. In practice, the different distribution parameters $\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K}$ can have many different shapes. Take, e.g. the multivariate $t$-distribution, parameterized using the Cholesky factor of the precision matrix $\bm{\mathrm{\Sigma}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{L}}$, denoted as $t_{D}(\bm{\mathrm{\theta}}_{1},\bm{\mathrm{\theta}}_{2},\bm{\mathrm{\theta}}_{3})\Leftrightarrow t_{D}(\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{\nu}})$. Then $\bm{\mathrm{\mu}}$ is a ${N\times D}$ matrix, $\bm{\mathrm{L}}$ is a ${N\times D\times D}$ cube (of which each vertical slice is a triangular matrix) and $\bm{\mathrm{\nu}}$ is a $N\times 1$ vector. Accordingly, the index set $\mathcal{M}_{k}$ of coordinates spans $\mathcal{M}_{1}=\{1,..,D\}$, $\mathcal{M}_{2}=\{(1,1),...,(D,D)\}$ and $\mathcal{M}_{3}=\{1\}$ and its cardinality is given by the product of the parameter’s dimensions beyond $N$. The general setting introduced here includes the Gaussian multivariate distributional regression introduced by Muschinski et al. (2022), the Copula-based multivariate distributional by Kock and Klein (2023) and the MCD-based additive covariance models by Gioia et al. (2022). The general estimation principle of repeatedly iterating through the distribution parameters until convergence translates in the multivariate case. However, we now introduce an additional inner cycle through all coordinates of the currently active distribution parameter. The exact estimation algorithm will be introduced in Section 2.3 and the following Section 2.2 briefly discusses different options to parameterize the covariance respectively precision matrix.
2.2 Parameterization of the Precision Matrix
Covariance Matrix Modeling Strategies
Covariance and precision matrices are commonly modeled in three distinct fashions: through the use of sparse or graphical estimators (e.g. the graphical LASSO, see Friedman et al., 2008), through covariance functions (see e.g. Browell et al., 2022) or the matrices’ elements arc modeled conditional on explanatory variables, usually through an unconstrained parametrization (see e.g. Pourahmadi, 2011; Muschinski et al., 2022; Salinas et al., 2019). Our approach falls into the third category. To save computational costs, we parameterize the distributions in terms of the inverse covariance matrix $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}$. This allows to avoid matrix inversion in the evaluation of the (log-) likelihood function. To ensure the positive definiteness of the scale matrix, we propose two unconstrained parameterizations of the precision matrix.
Cholesky-Decomposition
The CD has been introduced as suitable covariance parameterization by Pourahmadi (2011) in the context of GLMs. For the distributional regression framework, Muschinski et al. (2022); Kock and Klein (2023) employ the (modified) CD for the covariance matrix. For
$\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ (5)
Muschinski et al. (2022) parametrize the normal distribution in terms of $A^{-1}$ and Kock and Klein (2023) choose $A$. Additionally the modified Cholesky-decomposition (MCD) can be used, see e.g. Pourahmadi (2011); Muschinski et al. (2022); Gioia et al. (2022). For the CD to yield a positive semi-definite matrix, we require the diagonal of $\bm{\mathrm{A}}$ to be positive, which can be enforced by employing a log-link function. The lower diagonal of $\bm{\mathrm{A}}$ is unconstrained.
Low-Rank Approximation
The low-rank approximation (LRA) for the precision matrix has been proposed by Salinas et al. (2019) and März (2022) in the context of high-dimensional Gaussian processes and distributional gradient boosted trees. The LRA is defined as
$\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top},$ (6)
where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ and $\bm{\mathrm{V}}$ is a $D\times R$ matrix of rank $R$. The advantage of the LRA is that the dimensions of the parameters $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ scale linearly with the dimension $D$, however, the partial derivatives of the multivariate Gaussian and $t$-distribution with respect to the coordinates of $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ require inversion of the precision matrix. To ensure positive-definiteness for the LRA, we require the non-zero elements of $\bm{\mathrm{A}}$ to be positive, while $\bm{\mathrm{V}}$ is unconstrained. These requirements can easily be satisfied by choosing the log-link function for $\bm{\mathrm{A}}$.
2.3 Iterative Reweighted Least Squares for Distributional Regression
Overview
Rigby and Stasinopoulos (2005) introduce iteratively reweighted least squares for generalized additive models for location, shape and scale (GAMLSS). The RS algorithm consists of two nested loops, in which we cycle repeatedly through the distribution parameters and run a weighted fit of the score vector $u$ on the design matrix $\bm{\mathrm{X}}$ using the diagonal weight matrix $\bm{\mathrm{W}}$. The following paragraphs introduce the scoring vector and weights, the algorithm and the necessary modifications to move from a univariate case to the multivariate case.
Scoring and Weights
The score vector is defined as
$\bm{\mathrm{u}}=\frac{\partial\ell}{\partial\eta}$ (7)
where $\ell$ is the log-likelihood $\ell=\log(\mathcal{L})$ and $\eta=g(\bm{\mathrm{\theta}})$ is the linked predictor. The working vector for the Newton-Raphson or Fisher-Scoring algorithm is defined as
$\bm{\mathrm{z}}=g\big{(}\hat{\bm{\mathrm{\theta}}}\big{)}+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1}\Leftrightarrow\bm{\mathrm{z}}=\bm{\mathrm{\eta}}+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1}$ (8)
where the weights are defined as:
$\bm{\mathrm{W}}=-\frac{\partial^{2}\ell}{\partial\eta^{2}}$ or $\bm{\mathrm{W}}=-\mathbb{E}\left[\frac{\partial^{2}\ell}{\partial\eta^{2}}\right]$ (9)
for Newton-Raphson and Fisher’s scoring respectively. In the GLM, Fisher’s scoring and Newton-Raphson scoring coincide for the canonical link functions in the exponential family. However, for the scale and shape parameters, this is not necessarily the case anymore (for a detailed treatment of GLMs and estimation theory, see e.g. Lange et al., 2010). In the original GAMLSS, Rigby and Stasinopoulos (2005) use Fisher’s scoring. Our approach generally uses Newton-Raphson scoring for the multivariate case, since the derivation of the expected value of second derivatives can be intractable, especially for more complex parameterizations of the precision matrix.
Mix-and-Match Newton-Raphson Scoring
Newton-Raphson scoring requires the partial derivatives of the log-likelihood function with respect to the predictors. While many previous works on distributional regression employ Newton-Raphson scoring, each derive the partial derivatives for specific combinations of distribution function and link function only (see e.g. O’Neill and Burke, 2023; Muschinski et al., 2022). To facilitate the computational implementation in an mix-and-match fashion, we propose to use the first and second derivative of the log-likelihood with respect to the parameter and the first and second derivative of the link function and relate both to the necessary derivatives for Newton-Raphson scoring using the equalities given in the following Lemma 2.1, which allow for the simple utilization of arbitrary link functions and efficient calculation of working vector and weight matrices.
Distribution Location Scale $\bm{\Sigma}$ resp. Precision $\bm{\Omega}$ Shape
Param. Dim. Param. Dim. Param. Dim.
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ $N\times\text{triangular}(D\times D)$
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top}$ $N\times\text{diag}(D)$, $D\times r$
Multivariate-$t$ $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ $N\times\text{triangular}(D\times D)$ $\nu$ $N\times 1$
Multivariate-$t$ $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top}$ $N\times\text{diag}(D)$, $D\times r$ $\nu$ $N\times 1$
Table 1: Overview of multivariate distributions and scale matrix parametrization (Param.) implemented in the paper and the respective dimensions (Dim.) for input data $\bm{\mathrm{Y}}$ of shape $N\times D$. Note that the number of parameters for the CD-based parameterization grows quadratically in $D$, but the LRA-based parameterizations grow linear in $D$ for fixed $r$.
Lemma 2.1
Equipped with the first and second derivative of the log-likelihood with respect to the distribution parameter, ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$, as well as the first and second derivative of the link function $g(\cdot)$, we can retrieve the first and second derivative with respect to the predictor $\eta=g(\theta)$ as follows
$\frac{\partial\ell}{\partial\eta}$ $=\frac{\partial\ell}{\partial\theta}\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-1}\quad\quad\text{and}$ (10)
$\frac{\partial^{2}\ell}{\partial\eta^{2}}$ $=\left(\frac{\partial^{2}\ell}{\partial\theta^{2}}\frac{\partial g(\theta)}{\partial\theta}-\frac{\partial\ell}{\partial\theta}\frac{\partial^{2}g(\theta)}{\partial\theta^{2}}\right)\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-3}.$ (11)
The proof is straight-forward and utilizes the chain and quotient rules and can be found in Appendix A.3.
We provide the necessary first and second partial derivatives of the log-likelihood with respect to the distribution parameter’s coordinates, ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$, for all parameters for the multivariate normal and multivariate $t$-distribution given in Table 1. The derivation can be found in Appendix A.4 and Appendix A.5.
Link functions
Commonly, the log-link is used for scale parameters in the univariate and multivariate distributional regression models. However, the inverse transformation through the exponential is prone to yield extreme values (see e.g. Ziel, 2022). Narajewski and Ziel (2020) propose the use of the so-called logident link function, defined as
$g(x)=\operatorname{LogIdent}(x)=\begin{cases}\log(x)&\text{if $x<1$}\\ x-1&\text{else.}\end{cases}$ (12)
However, the function is not continuously twice differentiable. It can be made twice differentiable by a sigmoid spline on the non-differentiable part and then defined as
$g(x)=\operatorname{DifferentiableLogIdent}(x)=\begin{cases}\log(x)&\text{if $x<1$}\\ (1-f(x))\log(x)+f(x)(x-1)&\text{else.}\end{cases}$ (13)
where $f(x)=1/\exp(-k(x-1))$ is a sigmoid-function that ensures a smooth transition at $x=1$ for some constant $k$. Alternatively one can use the inverse softplus function as link function. The softplus and its inverse function are popular activation functions for neural networks (see e.g. Dubey et al., 2022; Dugas et al., 2000) and have been used in GLMs by Wiemann et al. (2024). The softplus and its inverse are defined as
$g(x)=\operatorname{InverseSoftPlus}(x)$ $=\log(\exp(x)-1)$ (14)
$g^{-1}(x)=\operatorname{SoftPlus}(x)$ $=\log(1+\exp(x))$ (15)
the inverse softplus function has (almost) linear behavior for large values of $x$ and log-like behavior for small $x$. Figure 3 compares the link functions
Figure 3: Comparison of the Log, Sqrt, LogIdent and InverseSoftPlus link and inverse functions. All map a distribution parameter to the positive real line $(0,\infty)$. The graceful, almost linear behavior of the inverse transformation of the InverseSoftPlus is clearly visible compared to the square root and log link.
2.4 Online Estimation Algorithm
High-level Overview
The IRLS algorithm consists of two nested loops. In the outer loop, we iterate through all distribution parameters. In the inner loop, we repeatedly run a weighted fit of the score vector $\bm{\mathrm{u}}$ on the design matrix $\bm{\mathrm{X}}$ using the weights $\bm{\mathrm{W}}$ until convergence. Note that in the inner loop, we run the weighted fit sequentially for all elements of the distribution parameter. Since the fit itself is agnostic to the regression technique (Stasinopoulos et al., 2024), we employ the online coordinate descent-based LASSO estimation here, as it has been proposed by Hirsch et al. (2024) for the univariate case already. Algorithm 2 gives an overview on the online estimation of multivariate distributional regression models. We define the index sets $\mathcal{K}=\{0,1,...,p-1\}$ for the number of parameters and $\mathcal{M}_{k}=\{0,1,..,M_{k}-1\}$ for the number of elements of each parameter as described in Section 2.1.
Input: $\bm{\mathrm{y}}^{[n+1]},\bm{\mathrm{X}}_{k,m}^{[n+1]}$ and the stored Gramian matrices $\bm{\mathrm{G}}_{km}^{[n]},\bm{\mathrm{H}}_{km}^{[n]}$.
1 Initialize the fitted values $\widehat{\bm{\mathrm{\theta}}}_{km}^{[n+1,0,0]}=\hat{\bm{\mathrm{\beta}}}_{km}^{[n]}(\bm{\mathrm{X}}_{k,m}^{[n+1]})^{\top}$ for $k,m\in\mathcal{K}\times\mathcal{M}$.
2 Evaluate the linear predictors ${\widehat{\bm{\mathrm{\eta}}}^{[n+1,0,0]}_{km}=g_{km}(\hat{\theta}_{km}^{[n+1,0,0]})}$ for $k,m\in\mathcal{K}\times\mathcal{M}$.
3 for $i=0,...$ until convergence do
4       forall $k\in\mathcal{K}$ do
5             Start the inner cycle and iterate over all elements of the distribution parameter.
6             for $r=0,1,...$ until convergence do
7                   forall $m\in\mathcal{M}_{k}$ do
8                         Evaluate $u_{km}^{[n+1,i,r]}$, $w_{km}^{[n+1,i,r]}$ and $z_{km}^{[n+1,i,r]}$ using Equations (7), (8) and (9).
9                         Update $\bm{\mathrm{G}}_{km}^{[n+1,i,r]}\leftarrow\gamma\bm{\mathrm{G}}_{km}^{[n]}+w_{km}^{[n+1,i,r]}\left((\bm{\mathrm{X}}_{km}^{[n+1]})^{\top}(\bm{\mathrm{X}}_{km}^{[n+1]})\right)$
10                         Update $\bm{\mathrm{H}}_{km}^{[n+1,i,r]}\leftarrow\gamma\bm{\mathrm{H}}_{km}^{[n]}+w_{km}^{[n+1,i,r]}\left((\bm{\mathrm{X}}_{km}^{[n+1]})^{\top}z_{km}^{[n+1,i,r]}\right)$
11                         Update $\hat{\bm{\mathrm{\beta}}}_{km\lambda}^{[n+1,i,r+1]}\leftarrow\hat{\bm{\mathrm{\beta}}}_{km\lambda}^{[n]}$ based on $\bm{\mathrm{G}}_{km}^{[n+1,i,r]}$ and $\bm{\mathrm{H}}_{km}^{[n+1,i,r]}$ using the online LASSO (see Algorithm 1) or recursive least squares.
12                         Select the optimal $\lambda$ using IC and set $\widehat{\bm{\mathrm{\beta}}}_{km}^{n+1,i,r+1]}\leftarrow\hat{\bm{\mathrm{\beta}}}_{km\lambda^{\text{opt}}}^{[n+1,i,r+1]}$.
13                         Calculate the updated $\hat{\bm{\mathrm{\eta}}}_{km}^{[n+1,i,r+1]}$ and $\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i,r+1]}$
14                         Evaluate the convergence.
15
16            End the inner cycle on the convergence of $\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i,r]}$.
17             Set $\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i+1,0]}\leftarrow\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i,r]}$ and set $\hat{\bm{\mathrm{\eta}}}_{km}^{[n+1,i+1,0]}\leftarrow\hat{\bm{\mathrm{\eta}}}_{km}^{[n+1,i,r]}$ and set $\hat{\bm{\mathrm{\theta}}}_{km}^{[n+1,i+1,0]}\leftarrow\hat{\bm{\mathrm{\theta}}}_{km}^{[n+1,i,r]}$.
18             End the outer cycle if the change in the penalized likelihood is sufficiently small.
19
20
Output: $\widehat{\bm{\mathrm{\beta}}}_{k,n+1}$ and $\widehat{\bm{\mathrm{\Theta}}}^{[n+1]}=(\widehat{\bm{\mathrm{\theta}}}_{0}^{[n+1]},...,\widehat{\bm{\mathrm{\theta}}}_{p}^{[n+1]})$ and the updated $\bm{\mathrm{G}}_{km}^{[n+1]}$ and $\bm{\mathrm{H}}_{km}^{[n+1]}$.
Algorithm 2 Online regularized multivariate distributional regression.
Update of the Gramian and Weights
For each inner iteration $i$, the update of the Gramian matrices starts at the Gramian matrices of $\bm{\mathrm{G}}_{km}^{[n]}$ and $\bm{\mathrm{H}}_{km}^{[n]}$ and the new information enters the Gramian matrices through the update of the weights and the working vector. However, the weights are also updated iteratively along each inner and outer iteration $i$ and $r$ due to the Newton-Raphson step towards the optimal coefficients. The weights can only be updated for the current update step $n+1$, while previous weights remain fixed. In a pure batch case, all weights are updated within each Newton-Raphson step. This introduces an approximation error for the online case, which can be controlled by the forget parameter $\gamma$ as shown in Hirsch et al. (2024).
Model selection
For each element of the distribution parameter, we estimate a regularization path. This raises the issue of model selection, i.e. the selection of the optimal regularization parameter $\lambda_{mk}^{\text{opt}}$. We propose to use information criteria (IC), as it is well-aligned to the likelihood-based framework of distributional regression. Define a generalized IC as
$\operatorname{IC}=-2\ell\left(\bm{\mathrm{Y}}\mid\widehat{\bm{\mathrm{\Theta}}}\right)+\nu_{0}K+\nu_{1}K\log(N)+\nu_{2}K\log\left(\log(N)\right)$ (16)
where $\ell$ is the log-likelihood under the model, $K$ is the number of parameters in the model and $N$ the number of seen observations. Let us note that we can recover Akaikes Information Criterion (AIC), the Bayesian Information Criterion (BIC) and the Hannan-Quinn Information Criterion (HQC) by setting $\nu_{0},\nu_{1},\nu_{2}$ accordingly. The optimal regularization parameter is then selected as $\lambda_{mk}^{\text{opt}}=\operatorname{argmin}_{\lambda}\operatorname{IC}$. Since the evaluation of the likelihood can be costly for high-dimensional data, we propose to employ the first derivative of the log-likelihood, i.e. calculate
$\ell\left(\bm{\mathrm{Y}}\mid\widehat{\bm{\mathrm{\Theta}}}^{[\lambda_{i}]}\right)\approx\ell\left(\bm{\mathrm{Y}}\mid\widehat{\bm{\mathrm{\Theta}}}^{[\lambda_{0}]}\right)+\frac{\partial\ell}{\partial\theta}\left(\widehat{\theta}_{km}^{[\lambda_{0}]}-\widehat{\theta}_{km}^{[\lambda_{i}]}\right)$ (17)
where the superscript ${[\lambda_{i}]}$ denotes the model with the regularization parameter $\lambda_{i}$. The approximation is valid for small changes in the regularization parameter and avoids the costly re-evaluation of the likelihood.
Step-size and Damping
The algorithm goes iteratively along all coordinates of the distribution parameter. The coordinates of the distribution parameters might impact each other, e.g. in the matrix multiplication of the CD-based scale matrix (see also the definition of the derivatives in A.4 and A.5). At the same time, we initialize the fitted values $\hat{\theta}_{m}$ as constant values. To stabilize the estimation, we propose to update the values in the very first iteration $i$ by a “dampened” version, i.e. taking
$\hat{\eta}_{m}^{[0,i]}\leftarrow g_{m}^{-1}\left((i+1)\hat{\theta}_{m}^{[0,i]}+\hat{\theta}_{m}^{[0,i-1]})/(i+1)\right)$
Hence, the predictions from the first iteration will be the average of the first fitted values and the initialization. This feature is mainly important for the scale matrix, whose coordinates are usually not orthogonal and less so for the location and (scalar) tail parameters.
Parallelization
Since the partial derivatives are not information orthogonal, the options for parallelization remain limited unfortunately. For the multivariate normal and $t$-distribution used in this paper, only the estimation of the location parameter can be parallelized in any case, as well as the estimation of the coordinates of the LRA matrix $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ for the normal distribution. For the $t$-distribution, the estimation of $\bm{\mathrm{A}}$ can only be parallelized for sufficiently high degrees of freedom. As parallelization would incur further open questions with respect to individual or joint regularization and model selection and the location parameter generally converges rather fast, we have not implemented parallel computation yet.
2.5 Path-based Regularized Estimation for the Scale Matrix
Idea
Often, some structure can be imposed on the covariance matrix, i.e. in spatial or temporal data, which has a clear dependence pattern along the diagonal. In these cases, the covariance matrix can be regularized by systematically setting far off-diagonal elements to zero (Gabriel, 1962; Zimmerman and Núñez-Antón, 1997; Zimmerman et al., 1998). While both, the CD-based and the LRA-based scale matrix parametrization lend themselves to this type of regularization, the approach is mainly popular with the Cholesky-based parameterization due to the relationship between the elements of the CD and the temporal correlation for longitudinal data under the name AD-$r$ regularization. However, such regularization is commonly applied a-priori and not in a data-driven fashion, see e.g. Muschinski et al. (2022); Zimmerman and Núñez-Antón (1997). On the other side, in coordinate descent estimation of regularized problems such as LASSO, path-based estimation starting from a strongly regularized solution towards an (almost) not regularized solution has proven itself as efficient solution approach. In this section, we aim to combine these two principles by introducing path-based estimation for the regularized scale matrix. On a high level, our algorithm starts with an “independence-parameterization” of the scale matrix and subsequently adds more non-zero elements and thus complexity to the parameterization of the scale matrix. Figure 4 illustrates how the path-based estimation uses increasingly complex specifications for the scale matrix $\bm{\mathrm{\Sigma}}$ respectively $\bm{\mathrm{\Omega}}$. Formally, for some regularization parameter $\alpha$, we set the elements of the scale to zero for
  • the CD-based parameterization if the indices $i,j$ are such that $|i-j|>\alpha$,
  • the LRA-based parameterization if the indices $d,r$ of $\bm{\mathrm{V}}$ are such that $r\geq\alpha$
and present the schematic overview in Algorithm 3. Note that we can use warm-starting for all previously fitted elements of the scale matrix, however, due to the non-orthogonality of the elements, we need to re-estimate all elements of the scale matrix in each iteration.
[…]
Figure 4: Path-based estimation along increasingly complex scale matrix parameterizations. The top panel shows the AD-$r$ regression for a Cholesky-based parameterization. The lower panel shows the estimation along the LRA-based parameterization, where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{d})$ is not regularized and the $D\times r$ matrix $\bm{\mathrm{V}}$ is filled column-wise with non-zero elements. Own Illustration.
1 for $\alpha=0,...,D$ do
2       Fit the online distributional regression Algorithm 2 for regularization level $\alpha$.
3       Evaluate the log-likelihood for the current regularization level $\alpha$.
4       Early Stop if the log-likelihood (or information criteria) does not increase sufficiently.
5
Output: Estimates for all $\alpha$.
Algorithm 3 Path-based scale-regularization for online multivariate distributional regression.
Implementation, Stability and Early Stopping
Note that both approaches can be used for parameterizations using the covariance and the precision matrix. The path-based estimation allows for re-using the previous iterations’ coefficients to achieve fast convergence in the OCD. For the CD-based parameterization, we increasingly add more off-diagonals to the lower-diagonal matrix. For the LRA-based parametrization, we add more and more columns to the low-rank matrix $\bm{\mathrm{V}}$. Let us note a few observations:
  • For a (small) fixed maximum regularization size, the number of parameters in the CD-based parameterization grows (almost) linear in $D$, alleviating the disadvantage of quadratic complexity.
  • For the multivariate $t$-distribution, independence is only achieved as $\nu\rightarrow\infty$. We therefore set a high initial guess ($\nu=10\operatorname{e}^{6}$) for the first outer iteration of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{\Omega}}$ to ensure numerical stability for the first iteration and subsequently choose a lower initial guess for the first iteration of $\nu$, since the Newton-Raphson algorithm relies on appropriate start values and tends to alternate between extrema otherwise (see e.g. Casella and Bachmann, 2021; Kornerup and Muller, 2006, on the impact of initial values for Newton-Raphson algorithms.). [^1] We have found the algorithm to iterate between $\nu=2$ and $\nu>10\text{e}^{10}$ for too large start values for the degrees of freedom. The proposed approach however has proved stable through the full simulation study with highly volatile electricity prices.
  • We can employ the path-based estimation to early stop the estimation, if the log-likelihood respectively an information criterion does not increase sufficiently by adding more non-zero elements. This allows for both, implicit regularization and decreased estimation time. However, once we early stop, we cannot increase the complexity of the parametrization in the online estimation but need to treat this as fixed.
Currently, the Algorithm will add only full off-diagonals (CD) respectively columns (LRA). The implementation however could also work for block-wise schemes (see e.g. the adaptive block structure in Cai and Yuan, 2012) or user-defined regularization patterns. The development of smart selection schemes for the next coordinates of the covariance matrix to include would be much beneficial for the speed of the algorithm.
3 Forecasting Study
Electricity Market Design
For electricity produced on day $t$ and hour $h$, the short-term electricity market in Germany is split in three major parts: The daily day-ahead auction on $t-1$ at 12:00 hours for 24 hourly delivery periods $h\in\{0,...,23\}$, the afternoon auction with quarter-hourly delivery periods on $t-1$, at 15:00 hours and the continuous intraday market. In 2024, two additional 15-minute auctions have been introduced at $t-1$, 22:00 hours and 10:00 on $t$. The daily procedure for the day-ahead auction, which is the focus of this paper, is shown in Figure 5. The market is organized by EPEX SPOT and Nordpool in the joint single day-ahead coupling (SDAC) as a pay-as-cleared auction through the EUPHEMIA algorithm, resembling the merit-order model for the electricity market (Billé et al., 2023; Hirsch et al., 2024; Viehmann, 2017).
[…]
Figure 5: Structure of the day-ahead electricity market in Germany. Own illustration based on information on EPEX SPOTs website and Viehmann (2017).
[…]
Figure 6: Repeated Batch Learning vs. Online Learning for the forecasting study. Own illustration.
Online Forecasting Study
At each day $t$, before the day-ahead auction at 12:00, we aim to generate forecasts for day $t+1$. Prior to forecasting, we update our models by taking into account the realized prices and forecasts for delivery day $t$. Figure 6 shows the structure of the online forecasting study. We are among the first studies to enforce a strict online setting for the forecasting study. Hence, on each day $t$, models are only updated using information revealed to the forecaster on this day. This is in contrast to the repeated batch learning approach, where the model is re-estimated on the full data set after each day.
Data and Notation
We use the same data set as in Marcjasz et al. (2023); Hirsch et al. (2024); Brusaferri et al. (2024a), which consists of electricity prices for the German day-ahead market from 2015-01-01 to 2021-01-01. In line with previous works, we use the data until 2018-12-26 as initial training set, leaving 736 observations and therefore more than 2 years, as it is best practice (Lago et al., 2021), for out-of-sample testing. Additionally, the data set contains day-ahead renewable production forecasts, load forecasts and prices for fundamental commodities. The data is available at the Github repository of Marcjasz et al. (2023). We denote the electricity price for day $t$ and hour $h\in 0,...,23$ as $y_{th}$ and therefore have $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{23})$ as the 24-dimensional ($D=H=24$) response matrix. We therefore, in this application study, we have $T$ corresponding to $N$ and $H$ corresponding to $D$ in the general notation. All fundamental features are briefly described in Table 2.
Variable Description Resolution Source
$\operatorname{RES}_{t,h}$ Day-ahead Renewable Energy Production Forecast Hourly ENTSO-E
$\overline{\operatorname{RES}}_{t}$ Day-ahead baseload RES Forecast $\frac{1}{H}\sum_{h=1}^{H}\operatorname{RES}_{t,h}$ Daily ENTSO-E
$\operatorname{Load}_{t,h}$ Day-ahead Electricity Load Forecast Hourly ENTSO-E
$\overline{\operatorname{Load}}_{t}$ Day-ahead baseload Load Forecast $\frac{1}{H}\sum_{h=1}^{H}\operatorname{Load}_{t,h}$ Daily ENTSO-E
$\operatorname{EUA}_{t}$ EU Emission Allowances Daily Refinitiv
$\operatorname{Gas}_{t}$ Natural Gas Prices Daily Refinitiv
$\operatorname{Coal}_{t}$ Coal Prices Daily Refinitiv
$\operatorname{Oil}_{t}$ Oil Prices Daily Refinitiv
$\operatorname{WD}_{t}$ Weekday dummies Daily Calender
Table 2: Variables from the data set of Marcjasz et al. (2023).
Model definition
We propose modeling the multivariate distribution of the day-ahead electricity prices in increasing complexity. All models are updated online, i.e. using only the new data for each day. We differentiate between an adaptive estimation, which is updating a single, unconditional distributional parameter and the full conditional estimation linking the distribution parameter to explanatory variables.
Model Mean / Location Marginal Distribution Dependence Structure
LARX Online conditional Adaptive but unconditional Adaptive but unconditional
Distr. Regression + Copula Online conditional Online conditional Adaptive but unconditional
Multivariate Dist. Regression Online conditional Online conditional Online conditional
Table 3: Increasingly complex model structures.
Table 3 shows the increasing model complexity. We start with the established LARX models and naively estimate the unconditional residual distribution (denoted as $\operatorname{LARX}+\mathcal{N}(0,\sigma)$ and $\operatorname{LARX}+\mathcal{N}(0,\Sigma)$). We increase complexity by moving to full distributional model for the marginals and adding an adaptive estimation of the dependence structure using the Gaussian copula (denoted as oDistReg+Copula). Lastly, we estimate the full multivariate distribution in a conditional way using the proposed multivariate online distributional regression approach (denoted as $\mathcal{F}$, parameterization, method)). We describe the full model in the following and note that we additionally describe the hyper parameters in Appendix A.2. For all three complexity levels, we include a reference model that assumes independence to showcase the value-add of including the dependence structure. We model the mean/location for all regression models by
$g_{\mu}(\mu_{t,h})$ $=\beta_{\mu,0,h}+\sum_{l=1}^{L=7}+\beta_{\mu,l,h}y_{t-l,h}+\sum_{h\in\{0,...,23\}/h}\beta_{\mu,8+h,h}y_{t-1,h}$ (18)
$+\beta_{\mu,31,h}\operatorname{min}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,32,h}\operatorname{max}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,33,h}\operatorname{Q}_{10}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,34,h}\operatorname{Q}_{90}(\bm{\mathrm{y}}_{t-1})$
$+\beta_{\mu,35,h}\operatorname{Load}_{t,h}+\beta_{\mu,36,h}\operatorname{RES}_{t,h}+\beta_{\mu,37,h}\overline{\operatorname{Load}}_{t}+\beta_{\mu,38,h}\overline{\operatorname{RES}}_{t}$
$+\beta_{\mu,39,h}\operatorname{EUA}_{t}+\beta_{\mu,40,h}\operatorname{Gas}_{t}+\beta_{\mu,41,h}\operatorname{Coal}_{t}+\beta_{\mu,42,h}\operatorname{Oil}_{t}+\sum_{w}^{W=6}\beta_{\mu,42+w,h}\operatorname{WD}_{t,h}$
We model the scale parameters for univariate distributional models, as well as the elements of the Cholesky-factor $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$, and the elements of the diagonal matrix $\bm{\mathrm{A}}$ in the LRA-based scale matrices by
$g_{\theta}(\theta_{t,h,h})$ $=\beta_{\theta,0,h}+\beta_{\theta,1,h}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\beta_{\theta,2,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)+\beta_{\theta,3,h}\overline{\operatorname{Load}}_{t}+\beta_{\theta,4,h}\overline{\operatorname{RES}}_{t}$ (19)
$+\beta_{\theta,5,h}\operatorname{Load}_{t,h}+\beta_{\theta,6,h}\operatorname{RES}_{t,h}+\beta_{\theta,7,h}\operatorname{EUA}_{t}+\beta_{\theta,8,h}\operatorname{Gas}_{t}+\beta_{\theta,9,h}\operatorname{Coal}_{t}+\beta_{\theta,10,h}\operatorname{Oil}_{t}$
where $\operatorname{SignedSquare}(a)=\operatorname{sign}(a)\sqrt{|a|}$ is the signed square root and $\bm{\mathrm{\Sigma}}^{[t-1:t-7]}$ is the rolling empirical covariance matrix of $\bm{\mathrm{y}}_{t}$ for the last 7 days. For the LRA-based parameterization, we choose $r=2$ and model the elements of $\bm{\mathrm{V}}$ as
$g_{v}(v_{t,h,0})$ $=\beta_{v,0,h}+\beta_{v,1,h}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\beta_{v,2,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)+\beta_{v,3,h}\operatorname{min}(\bm{\mathrm{y}}_{t-1})$ (20)
$+\beta_{v,4,h}\operatorname{max}(\bm{\mathrm{y}}_{t-1})+\beta_{v,5,h}\operatorname{Q}_{10}(\bm{\mathrm{y}}_{t-1})+\beta_{v,6,h}\operatorname{Q}_{90}(\bm{\mathrm{y}}_{t-1})+\beta_{v,7,h}\overline{\operatorname{Load}}_{t}+\beta_{v,8,h}\overline{\operatorname{RES}}_{t}$
$+\beta_{v,9,h}\operatorname{Load}_{t,h}+\beta_{v,10,h}\operatorname{RES}_{t,h}+\beta_{v,11,h}\operatorname{EUA}_{t}+\beta_{v,12,h}\operatorname{Gas}_{t}+\beta_{v,13,h}\operatorname{Coal}_{t}+\beta_{v,14,h}\operatorname{Oil}_{t}$
$g_{v}(v_{t,h,1})=\sum_{w}^{W=6}\beta_{v,14+w,h}\operatorname{WD}_{t,h}$ (21)
that is, the first rank takes most of the fundamental variables, while the second rank contains the weekday binary variables. The degrees of freedom are modeled as
$g_{\nu}(\nu_{t})$ $=\beta_{\nu,0}+\beta_{\nu,1}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\sum_{w}^{W=6}\beta_{\nu,1+w,h}\operatorname{WD}_{t,h}+\beta_{\nu,8}\overline{\operatorname{Load}}_{t}+\beta_{\nu,9}\overline{\operatorname{RES}}_{t}$ (22)
$+\beta_{\nu,10}\operatorname{EUA}_{t}+\beta_{\nu,11}\operatorname{Gas}_{t}+\beta_{\nu,12}\operatorname{Coal}_{t}+\beta_{\nu,13}\operatorname{Oil}_{t}.$
The univariate models are therefore a slight simplification compared to the models used in Hirsch et al. (2024), however thereby the multivariate distributional regression models and the Copula-based approaches are better comparable. Lastly, let us remark on the online tracking of the Gaussian copula. The probability density function (PDF) for the Gaussian copula is given by:
$\ell(\bm{\mathrm{u}}\mid\bm{\mathrm{\Sigma}})=\frac{1}{|\tilde{\bm{\mathrm{\Sigma}}}|^{1/2}}\exp\left(-\frac{1}{2}\bm{\mathrm{n}}^{\top}(\tilde{\bm{\mathrm{\Sigma}}}^{-1}-\bm{\mathrm{I}})\bm{\mathrm{n}}\right)\prod_{d=0}^{D}p(y_{d}\mid\bm{\mathrm{\theta_{d}}})$ (23)
where $\bm{\mathrm{u}}$ are the pseudo-observations on the $\mathcal{U}(0,1)$ space, $\bm{\mathrm{n}}=\Phi^{-1}(\bm{\mathrm{u}})$, $\Phi$ is the CDF of the standard normal distribution and $\tilde{\bm{\mathrm{\Sigma}}}$ is the covariance matrix $\bm{\mathrm{\Sigma}}$ scaled to the correlation matrix, $\bm{\mathrm{I}}$ is the identity matrix and $p(x_{d}\mid\bm{\mathrm{\theta}}_{d})$ is the likelihood of the observation $y_{d}$ under the (conditional) marginal distribution (see Kock and Klein, 2023; Arbenz, 2013). We fit the Copula model by the transforming the in-sample data to the uniform space $\bm{\mathrm{u}}$ by the probability integral transformation (PIT) and subsequently transforming to the $\mathcal{N}(0,1)$ space $\bm{\mathrm{n}}$, on which we can fit the dependence structure. We update the scale matrix of the Gaussian copula by taking
$\widehat{\bm{\mathrm{\Sigma}}}^{[t+1]}=\frac{t-1}{t}\widehat{\bm{\mathrm{\Sigma}}}^{[t]}+\frac{1}{t}\left(\bm{\mathrm{n}}^{[t+1]}(\bm{\mathrm{n}}^{[t+1]})^{\top}\right)$ (24)
where $\bm{\mathrm{n}}$ are the PIT-transformed in-sample values and the superscript $[t]$ denotes the observations available in the online learning (see e.g. Dasgupta and Hsu, 2007). Samples are drawn from the Gaussian copula in the usual manner. We use the same principle to track the residual covariance structure for the LARX models under the normality assumption. We employ a second model, where we sparsify the estimated dependence matrix of the Gaussian copula by the graphical LASSO (Friedman et al., 2008).
Scoring Rules
We employ four well-established multivariate probabilistic scoring scores: The Energy Score (ES), the Dawid-Sebastiani Score (DSS), the Variogram Score (VS) and the Log-Score (LS). Additionally, we employ the root mean square error RMSE (RMSE), mean absolute error (MAE) and the continuous ranked probability score (CRPS). We test for statistically significant score differences using the well-established Diebold-Mariano test. The following paragraphs introduce the scores and are largely based on Gneiting et al. (2007); Gneiting and Raftery (2007); Nowotarski and Weron (2018); Marcotte et al. (2023); Ziel and Berk (2019) as well as the references mentioned for the individual scores. Denote the true price vector as $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{H})$ of shape $T\times H$ and an ensemble forecast $\bm{\mathrm{F}}$ of shape $T\times H\times M$ of $M=2500$ samples. The RMSE is defined as
$\operatorname{RMSE}=\sqrt{\frac{1}{TH}\sum_{t=0}^{T}\sum_{h=0}^{H}\left(\bm{\mathrm{y}}_{t}-\hat{\bm{\mathrm{\mu_{t}}}}\right)^{2}}$ (25)
where $\hat{\bm{\mathrm{\mu}}}_{t}=\frac{1}{M}\sum_{m=0}^{M}\bm{\mathrm{F}}_{t,h,m}$ is the mean prediction vector. The MAE is defined as
$\operatorname{MAE}=\frac{1}{TH}\sum_{t=0}^{T}\sum_{h=0}^{H}\left|\bm{\mathrm{y}}_{t}-\operatorname{median}\left(\bm{\mathrm{F_{t}}}\right)\right|$ (26)
where $\operatorname{median}\left(\bm{\mathrm{F_{t}}}\right)$ denotes the median trajectory for each day $t$. The CRPS is estimated from the forecast ensemble by using the probability-weighted moment estimator of Zamo and Naveau (2018):
$\operatorname{CRPS}_{t}=\frac{1}{M}\sum_{m=0}^{M}|\bm{\mathrm{F}}_{t,h,m}-y_{t.h}|+\frac{1}{M}\sum_{m=0}^{M}\bm{\mathrm{F}}_{t,h,m}+\frac{1}{M(M-1)}\sum_{m=0}^{M}m\bm{\mathrm{F}}_{t,h,m}$ (27)
The CRPS is strictly proper scoring rule for the marginal distribution. Note that many works on energy price forecasting report the average pinball loss (APS) as CRPS, which needs to be rescaled $\operatorname{CRPS}=2\cdot\operatorname{APS}$ to be comparable. The energy score (ES, Gneiting and Raftery, 2007) is defined as
$\operatorname{ES}_{t}=\frac{1}{M}\sum_{m=0}^{M}\lVert\bm{\mathrm{y}}_{t}-\bm{\mathrm{F}}_{t,m}\rVert_{2}^{2}-\frac{1}{M^{2}}\sum_{i=0}^{M}\sum_{j=i+1}^{M}\lVert\bm{\mathrm{F}}_{t,i}-\bm{\mathrm{F}}_{t,j}\rVert_{2}^{2}.$ (28)
The energy score is a strictly proper scoring rule, however, Alexander et al. (2024); Pinson and Tastu (2013); Marcotte et al. (2023) argue that the ES is rather insensitive to misspecified dependence structures. We aggregate the ES by taking the average: $\operatorname{ES}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{ES}_{t}$. The Log-Score (LS) is defined as
$\operatorname{LS}_{t}=-\log\left(\mathcal{L}(\bm{\mathrm{y}}_{t}\mid\hat{\bm{\mathrm{\theta}}}^{\mathcal{D}}_{t})\right)$ (29)
where $\mathcal{L}$ is the underlying likelihood or probability density function of the distribution $\mathcal{D}$ and $\hat{\bm{\mathrm{\theta}}}^{\mathcal{D}}_{t}$ is the estimated parameter vector. Again, we aggregate the LS by simple averaging over all points in the test set $\operatorname{LS}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{LS}_{t}$. It is a strictly proper scoring rule. The Dawid-Sebastiani-Score (DSS, 1999) is defined as
$\operatorname{DSS}_{t}=\log\left(\det(\widehat{\bm{\mathrm{\Sigma}}}_{F})\right)+(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}})\widehat{\bm{\mathrm{\Sigma}}}_{F}^{-1}(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}})$ (30)
where $\hat{\Sigma}_{F}$ denotes the empirical covariance of the forecast ensemble $\bm{\mathrm{F}}$ and $\hat{\bm{\mathrm{\mu}}}$ denotes the mean ensemble as above. We aggregate the $\operatorname{DSS}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{DSS}_{t}$ by simple averaging. The DSS is a proper scoring rule for the first and second moment and strictly proper for the Gaussian predictive distribution, since it is a linear transformation of Gaussian log-likelihood. The Variogram Score (VS, Scheuerer and Hamill, 2015) is defined as
$\operatorname{VS}^{p}_{t}=\sum_{i=0}^{H}\sum_{j=0}^{H}\left(\frac{1}{M}\sum_{m=0}^{M}|\bm{\mathrm{F}}_{t,i,m}-\bm{\mathrm{F}}_{t,j,m}|^{p}-|y_{t,i}-y_{t.j}|^{p}\right)^{2}$ (31)
and is a proper scoring rule. We aggregate the VS by taking the average and normalize the score by dividing by $H^{2}$ and taking the square root, i.e. $\operatorname{VS}=\frac{1}{T}\sum_{t=0}^{T}\sqrt{\frac{1}{H^{2}}\operatorname{VS}_{t}}$ to make the scales of the score comparable. The scoring rules used are implemented in the Python package scoringrules (Zanetta and Allen, 2024).
Diebold-Mariano-Test
Conclusions on the performance of forecasting models cannot be drawn from looking at aggregate scores alone, but need to be drawn by evaluating whether the differential between the loss series of two models is statistically significantly from zero (Diebold and Mariano, 2002; Diebold, 2015). For the DM-test, we evaluate the differential of two score series $\Delta\bm{\mathrm{s}}^{A,B}=\bm{\mathrm{s}}^{A}-\bm{\mathrm{s}}^{B}$, where $\bm{\mathrm{s}}^{\mathcal{A}}=(s_{0}^{A},...,s_{T}^{A})$ are the scores for each scoring rule at $t$ for model $A$ respectively $B$. We provide two one-sided and hence complimentary tests.
4 Results
This section describes the results from the forecasting study. Exemplary simulations and predicted covariance matrices are shown in Figures 7 and 8. Some time-varying behavior of the covariance matrix over the week, especially in the morning hours is visible. We present aggregate scoring rules in Table 4 and significance testing using the DM-test in Figure 9 and discussed in the following paragraph. Computation times are discussed in the last paragraph and given in Table 5.
Figure 7: Illustrative predicted covariance matrices for one week in the test sample.
Figure 8: Illustrative simulations drawn from two models.
Scores
Table 4 gives the results for the scoring rules for each model and Figure 9 provides one-side Diebold-Mariano-tests for all pairwise model comparisons. As a mental guidance for the increasing complexity of the online regression models, remember Section 3: The first two models have an adaptive, but unconditional estimation for the scale parameter/matrix under a Gaussian assumption. The Copula-based models employ online, conditional estimation for all marginal distributional parameters and an adaptive, but unconditional estimation for the dependence structure, while the multivariate distributional regression models yield an online estimation conditional multivariate distribution. Let us note a few main results from Table 4 here:
  • The baseline LARX-models yield the best performance in terms of the RMSE. This is common theme in probabilistic forecasting using distributional models due to the fact that the likelihood-based estimation down-weights observations with high (estimated) variance, thereby reducing the precision in the mean estimation to improve the distributional fit overall. Similar results can be observed e.g. in Marcjasz et al. (2023); Hirsch et al. (2024). The LARX model with the multivariate Gaussian distribution also yields, across all various metrics, very robust results.
  • The two univariate models, the $\operatorname{LARX}+\mathcal{N}(0,\sigma)$ and the univariate distributional regression model without Copula yield, not surprisingly, weak scores for the multivariate scoring rules (VS, ES, DSS and LS), while being competitive in the univariate scores with their respective counterparts. This result underscores that univariate models cannot capture the dependence structure and therefore miss a crucial element for probabilistic forecasting of electricity prices.
  • The univariate online distributional regression including the copula yields strong results for VS and the ES. However, for the DSS and the LS, the Copula-based models do not perform as well. An issue here might be necessary two-step approach, which naturally introduces some friction in the estimation of the scale matrix if the marginal assumption does not fit perfectly.
  • For the multivariate online distributional regression models, we see that the models using the Cholesky-based parametrization provide better performance. This is likely due to the fact that the regularized CD is closer to the natural, time-based structure of the (conditional) covariance than the LRA. This is also visible in Figure 8, which visually compares trajectories from the two scale matrix parameterizations.
  • Overall, the estimation using LASSO compared to OLS increases the forecasting performance for the multivariate distributional regression models significantly. The comparison of the independence configuration of the distributional regression models and their unrestricted counterparts show that the CD-based models yield higher RMSE errors, most likely through cross-transmission of errors in the covariance matrix.
  • The $p$-values of the DM-test in Figure 9 largely confirm the statistical significance of the aforementioned results. We note the strong performance of the Copula-based models for the ES and the statistically significant superior performance of the multivariate distributional regression for the DSS and LS.
Overall, our results highlight that neglecting the dependence structure by relying solely on marginal, univariate models yields subpar probabilistic forecasting performance. We note that for the truly multivariate approaches, using both Copula-based combinations of univariate models and the fully multivariate distributional regression yield statistically significant performance improvements. However, the CRPS and Energy Score are not improved by the multivariate models. This might be due to the fact that the dependence structure in the day-ahead electricity prices does not change significantly over time, as well as error transmission in modelling the conditional scale matrix.
Model RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS
LARX + N(0, $\sigma$) 7.346 4.530 3.406 1.017 6.373 20.935 119.170 81.510
LARX + N(0, $\Sigma$) 7.346 4.532 3.406 0.885 5.712 20.455 84.487 64.149
oDistReg 7.463 4.393 3.305 1.012 6.546 20.701 114.483 76.754
oDistReg+GC 7.454 4.394 3.323 0.851 5.548 20.252 103.057 61.536
oDistReg+spGC 7.449 4.391 3.320 0.851 5.548 20.238 102.207 61.476
oMvDistReg(t, CD, OLS, ind) 8.027 4.447 3.456 0.851 5.495 21.351 137.906 70.410
oMvDistReg(t, LRA, OLS, ind) 8.014 4.440 3.440 0.855 5.518 21.276 134.802 70.083
oMvDistReg(t, CD, OLS) 8.345 4.723 3.559 0.861 5.574 21.762 89.068 55.861
oMvDistReg(t, LRA, OLS) 8.165 4.605 3.556 0.881 5.656 21.980 135.251 70.280
oMvDistReg(t, CD, LASSO) 8.323 4.664 3.494 0.870 5.637 21.420 81.496 55.356
oMvDistReg(t, LRA, LASSO) 8.858 5.190 3.929 0.992 6.433 24.045 128.334 73.295
Table 4: Scoring Rules for the full out of sample period of 736 days. The best score in each column is marked bold. Note that the $\operatorname{LARX}+\mathcal{N}(0,\sigma)$, the oDistReg and the oMvDistreg(…, ind) models do not model the dependence structure.
Figure 9: Diebold-Mariano Test Matrix. A $p$-value $p<0.05$ implies that the forecasts given by a model on the column are significantly better than forecasts by a model on the row.
Computation times
Table 5 gives computation times for all experiments. The initial fit for the multivariate distributional regression model takes a few minutes, the update algorithm can be executed in seconds. On a standard laptop, the experiments can be run in about 2 hours. Even though we did not run a repeated batch fitting, an estimate for the benefit of online vs. repeated batch fitting can be achieved by multiplying the initial fit duration with 736 days of out of sample and comparing this to the total time of the online study:
Speedup=Initial $\text{Speedup}=\frac{\text{Initial Fit}\times T}{\text{Total Time}}.$
By this (albeit simple) measure, the online learning improves computation by a factor of 80 to 400. Thereby our online update algorithm makes the approach practically viable for researchers and data scientists without access to specialized high-performance computation centers. These estimates are in line with benefits reported in Hirsch et al. (2024) for the univariate online distributional regression case in an explicit comparison.
Model Initial Fit Avg. Update Std. Update Total Time Est. Speedup
LARX + N(0, $\sigma$) 1.74 0.02 0.04 13.70 $\times$93
LARX + N(0, $\Sigma$) 1.74 0.02 0.04 13.70 $\times$93
oDistReg 25.08 0.12 0.07 114.26 $\times$161
oDistReg+GC 25.08 0.12 0.07 114.71 $\times$160
oDistReg+spGC 25.28 0.29 0.07 241.68 $\times$76
oMvDistReg(t, CD, OLS, ind) 46.44 0.09 0.07 113.68 $\times$300
oMvDistReg(t, LRA, OLS, ind) 55.32 0.11 0.03 132.70 $\times$306
oMvDistReg(t, CD, OLS) 146.94 0.31 0.01 377.33 $\times$286
oMvDistReg(t, LRA, OLS) 188.24 1.96 0.17 1628.10 $\times$85
oMvDistReg(t, CD, LASSO) 290.06 3.07 0.03 2549.02 $\times$83
oMvDistReg(t, LRA, LASSO) 864.99 0.68 0.04 1363.32 $\times$466
Table 5: Computation times. All timings are in seconds. The out-of-sample data for the forecasting study consists of 736 days. We update the 24-dimensional distributional regression model each model on each day. All experiments are run on a standard laptop (Intel Core i7 (16 Threads, 4.9 GHz), 32GB RAM). Estimated speead-ups are calculated by taking Speedup=(Initial $\text{Speedup}=(\text{Initial Fit}\times T)/\text{Total Time}$.
5 Discussion and Conclusion
Summary and Contribution
Distributional learning algorithms such as GAMLSS and deep distributional networks have been used successfully for probabilistic electricity price forecasting (PEPF, see e.g. Muniain and Ziel, 2020; Hirsch et al., 2024; Marcjasz et al., 2023). However, even for univariate distributions, these models are computationally expensive. At the same time, the literature on probabilistic electricity price forecasting has largely focused on modeling the hourly marginal distributions only, leaving the dependence structure neglected. Against this background, we develop an online estimation algorithm for multivariate distributional regression models, making the use of these algorithms feasible even for high-dimensional problems such as the 24-dimensional distribution of electricity prices on a standard laptop. We benchmark our implementation in a forecasting study for the German day-ahead electricity market and thereby provide the first study exclusively focused on online learning for multivariate PEPF.
Main Results
Our results show that modeling the dependence structure improves forecasting performance and that multivariate distributional regression models yield superior results in terms of the Variogram, Log-Score (LS) and Dawid-Sebastiani Score (DSS) compared to simple benchmark models (LARX). However, we note that multivariate distributional regression models perform better in terms of the LS and DSS than univariate distributional models using a Gaussian copula, while performing worse in terms of the CRPS and Energy Score. This behavior is likely driven by error transmission in the conditional dependence structure, but might also hint at the possibility that the dependence structure in the day-ahead electricity does not change significantly over time. We also find that the Cholesky-based parametrization (which has been used as well in Gioia et al., 2022; Kock and Klein, 2023; Muschinski et al., 2022) of the scale matrix yields better results than the LRA-based parametrization proposed by Salinas et al. (2019). Lastly, we find that the LASSO-regularized estimation of the scale matrix improves forecasting performance significantly. On the computational side, our study can be estimated in roughly 2-3 hours on a standard laptop, providing estimated speed-ups between 80 and 400+ times compared to repeated batch fitting.
Implementation
We implement our algorithm in a fairly generic manner, allowing e.g. for different distributional assumptions and keeping a familiar, sklearn-like API to facilitate the usage by other researchers and data scientists (Pedregosa et al., 2011). We employ just-in-time compilation using numba to further improve the computation speed (Lam et al., 2015). Currently, the code is available on GitHub at https://github.com/simon-hirsch/online-mv-distreg and will be contributed to the ROLCH package (Hirsch et al., 2024).
Future Research
Our research opens multiple avenues for future research. First, further research on the driving forces of the dependence structure in the German electricity market is necessary to improve the forecasting performance and guide decision-making processes in electricity trading. Modeling the dependence structure in electricity markets is a rather open field and has implications beyond forecasting, concerning also risk and portfolio management and asset optimization (Peña et al., 2024; Löhndorf and Wozabal, 2023; Beykirch et al., 2022, 2024). From an algorithmic perspective, we note that while our algorithm is already quite fast, further improvements in the computation speed might be possible by using a CG-type scoring algorithm (Rigby and Stasinopoulos, 2005; Green, 1984; Cole and Green, 1992) and parallelizing over the elements of the distribution parameter. A further open issue is model selection - while the regularized online estimation is fast, the models are still quite complex and can be prone to overfitting. Lastly, due to the generic nature of our implementation, the usage for other high-dimensional forecasting problems such as probabilistic wind, solar and load forecasting can be explored.
Acknowledgments
Simon Hirsch is employed as an industrial PhD student by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. Simon Hirsch is grateful for helpful discussions with Florian Ziel and Daniel Gruhlke.
Declaration of Interest
The author declare that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Generative AI Statement
During the preparation of this work the author used generative AI tools such as ChatGTP and GitHub Copilot in order to improve the quality of the language and of the code. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication
Appendix A Appendix
A.1 Abbreviations
AIC Akaike Information Criterion
APS Average Pinball Score
BIC Bayesian Information Criterion
CD Cholesky-Decomposition
CDF Cumulative Density Function
DDNN Distributional Deep Neural Networks
DSS Dawid-Sebastiani Score
EPF Electricity Price Forecasting
ES Energy Score
GAMLSS Generalized Additive Models for Location, Scale and Shape
GLM Generalized Linear Model
HQC Hannan-Quinn Criterion
IC Information Criterion
IRLS Iteratively Reweighted Least Squares
LASSO Least Absolute Shrinkage and Selection Operator
LARX LASSO-estimated AutoRegressive Model with eXogenous variables
LRA Low-Rank Approximation
LS Log-Score (= negative log-likelihood)
MAE Mean Absolute Error
OCD Online Coordinate Descent
OLS Ordinary Least Squares
PDF Probability Density Function
PEPF Probabilistic Electricity Price Forecasting
PIT Probability Integral Transformation
RMSE Root Mean Squared Error
RS Rigby & Stasinopolous (Algorithm)
VS Variogram Score
Table 6: Abbreviations used in the Paper.
A.2 Hyperparameters for the Multivariate Distributional Regression Model
To align with the reproducible research best practices, as described in e.g Lago et al. (2021), we publish the reproduction code on GitHub at https://github.com/simon-hirsch/online-mv-distreg, allowing for full reproducibility of all experiments. Additionally, we take the following paragraph to describe the hyperparameters of the model:
  • Information criteria and model selection: We use the AIC for the multivariate distributional regression model and run the online coordinate descent on an exponential grid of 50 $\lambda$ values. We employ fast model selection based on the first derivatives for the CD-based models.
  • Link functions: We use the identity link for the location for all models. For the CD-based distributional models, we use the InverseSoftPlusLink. For the LRA-based models, we employ the SqrtLink for the diagonal matrix $\bm{\mathrm{A}}$ as initial experiments showed a more robust convergence behavior and the IdentityLink for the matrix $\bm{\mathrm{V}}$. For the degrees of freedom $\nu$, we employ the LogShiftTwoLink, which ensures that $\nu>2$ and hence the covariance matrix is positive definite.
  • Early stopping: We employ early stopping for the path-based regularization of the scale matrix if the AIC does not improve, as described in Section 2.5. We limit the number of off-diagonals for the CD-based parameterization to max 6, however note that the algorithm breaks after fitting 2-3 off-diagonals. We do not limit the number of columns fitted in the LRA-based model and note that the algorithm breaks after fitting the full rank-2 matrix $\bm{\mathrm{V}}$.
  • Number of iterations, step-size and dampening: We dampen the estimation in the first iteration for the scale parameters only. We generally allow for a maximum of 30 inner and 10 outer iterations in the initial fit and the update steps.
A.3 Derivation of Equation 10 and 11 for Newton-Raphson Scoring
We aim to calculate ${\partial\ell}/{\partial\eta}$ and ${\partial^{2}\ell}/{\partial\eta^{2}}$ for the calculation of the score and weight vectors (see Eq. 10 and Eq. 11) using the partial derivatives of the log-likelihood with respect to the distribution parameter (resp. the coordinate of the distribution parameter in case of matrix-valued parameters), ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$. For continuous, twice differentiable link functions $\eta=g(\theta)$, we have
$\frac{\partial\ell}{\partial\eta}$ $=\frac{\partial\ell}{\partial\theta}\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-1}$ (32)
and
$\frac{\partial^{2}\ell}{\partial\eta^{2}}$ $=\frac{\partial\left(\dfrac{\partial\ell}{\partial\theta}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}\right)}{\partial\eta}=\frac{\partial\left(\dfrac{\partial\ell}{\partial\theta}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}\right)}{\partial\theta}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}$
and by the quotient rule, we have
$\frac{\partial^{2}\ell}{\partial\eta^{2}}=\left(\frac{\dfrac{\partial^{2}\ell}{\partial\theta^{2}}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)-\dfrac{\partial\ell}{\partial\theta}\left(\dfrac{\partial^{2}g(\theta)}{\partial\theta^{2}}\right)}{\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{2}}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}\right)$ (33)
and the simplification
$\frac{\partial^{2}\ell}{\partial\eta^{2}}=\left(\frac{\partial^{2}\ell}{\partial\theta^{2}}\frac{\partial g(\theta)}{\partial\theta}-\frac{\partial\ell}{\partial\theta}\frac{\partial^{2}g(\theta)}{\partial\theta^{2}}\right)\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-3}$ (34)
concludes the derivation $\blacksquare$
A.4 Partial derivatives of the multivariate Gaussian Distribution
The probability density function of the multivariate normal distribution of dimension $D$ is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Sigma}})=\frac{1}{(2\pi)^{D/2}|\bm{\mathrm{\Sigma}}|^{1/2}}\exp\left(-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)$ (35)
with the location or mean vector $\bm{\mathrm{\mu}}$ and the scale respectively covariance matrix $\bm{\mathrm{\Sigma}}$. We parameterize the PDF in terms of the inverse scale matrix $\bm{\mathrm{\Omega}}=\bm{\mathrm{\Sigma}}^{-1}$:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Sigma}})=\frac{1}{(2\pi)^{D/2}}|\bm{\mathrm{\Omega}}|^{1/2}\exp\left(-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)$ (36)
and calculate the log-likelihood as $\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}})=\log(f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}}))$, which reads:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Sigma}})=-\frac{D}{2}\log(2\pi)-\frac{1}{2}\log(|\bm{\mathrm{\Sigma}}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (37)
and parameterize the inverse covariance matrix through the Cholesky-decomposition $\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$, which yields:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1})=-\frac{D}{2}\log(2\pi)-\log(|\bm{\mathrm{A}}^{-1}|)-\frac{1}{2}\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$ (38)
where $\bm{\mathrm{z}}=\bm{\mathrm{A}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ and $z=\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$. The first derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{A}}^{-1}$ are given in Muschinski et al. (2022) and read
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Omega}}_{ik}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})$ (39)
$\frac{\partial\ell}{\partial(\bm{\mathrm{A}}^{-1})_{ij}}$ $=\frac{1}{(\bm{\mathrm{A}}^{-1})_{ij}}-(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{A}}^{-1})_{kj}$ (40)
and the second derivatives are given by
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\bm{\mathrm{\Omega}}_{ii}$ (41)
$\frac{\partial\ell^{2}}{\partial(\bm{\mathrm{A}}^{-1})_{ij}^{2}}$ $=-\frac{1}{(\bm{\mathrm{A}}^{-1})^{2}_{ij}}-(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (42)
For the low-rank approximation, we parameterize Equation 36 in terms of the low-rank approximation $\bm{\mathrm{\Omega}}=\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}$, which yields:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{U}},\bm{\mathrm{V}})=-\frac{D}{2}\log(2\pi)-\log(|(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})^{-1}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (43)
we note that the derivatives with respect to the elements of the mean vector $\bm{\mathrm{\mu}}_{i}$ remain the same:
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Omega}}_{ik}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})$ (44)
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\bm{\mathrm{\Omega}}_{ii}$ (45)
and the derivatives with respect to the elements of $\bm{\mathrm{D}}_{ii}$ are given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{1}{2}\left(\bm{\mathrm{\Sigma}}_{ii}-(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})^{2}\right)$ (46)
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{D}}_{ii}^{2}}$ $=-\frac{1}{2}\Sigma_{ii}^{2}.$ (47)
The partial derivatives with respect to the elements of $\bm{\mathrm{V}}$ are given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{V}}_{ij}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}$ (48)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{V}}_{ij}^{2}}$ $=\bm{\mathrm{\Sigma}}_{ij}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{ii}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{qk}\bm{\mathrm{V}}_{kj}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{iq}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}$ (49)
$-\left((\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}-\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}\right)^{2}\right)$
which concludes the derivation of the partial derivatives $\blacksquare$
A.5 Partial derivatives of the multivariate t-distribution
The probability density function (PDF) of the multivariate $t$-distribution of dimension $D$ is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\Sigma,\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}|\bm{\mathrm{\Sigma}}|^{1/2}}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}$
with the location vector $\mu$, the shape matrix $\bm{\mathrm{\Sigma}}$ and the degrees of freedom $\nu$. We parameterize the PDF in terms of the inverse shape matrix $\bm{\mathrm{\Omega}}=\bm{\mathrm{\Sigma}}^{-1}$:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|\bm{\mathrm{\Omega}}|^{1/2}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}.$ (50)
We start with the partial derivatives for the CD-based parametrization. We have the Choleksy-decomposition $\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$, which yields:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|(\bm{\mathrm{A}}^{-1})|\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}.$
Let us introduce some notation to simply the following derivatives. Define:
$\bm{\mathrm{z}}$ $=\bm{\mathrm{A}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (51)
$z$ $=\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$ (52)
The log-likelihood is given by $\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\log(f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu))$ and reads:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|(\bm{\mathrm{A}}^{-1})|\right)+\log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{z}}^{T}\bm{\mathrm{z}})\right)^{-(\nu+D)/2}\right)$ (53)
For the partial derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{A}}^{-1}$, we notice that $\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$ can be treated as a function of these elements and employ the chain rule. We see that:
$\frac{\partial(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})}{\partial\bm{\mathrm{\mu}}_{i}}$ $=2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})$ (54)
$\frac{\partial^{2}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-2\bm{\mathrm{\Omega}}_{ij}$ (55)
$\frac{\partial(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})}{\partial(\bm{\mathrm{A}}^{-1})_{ij}}$ $=2(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\sum_{m=1}^{M=j}(\bm{\mathrm{y}}_{m}-\bm{\mathrm{\mu}}_{m})(\bm{\mathrm{A}}^{-1})_{mj}$ (56)
$\frac{\partial(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{2}}{\partial(\bm{\mathrm{A}}^{-1})_{ij}^{2}}$ $=(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (57)
The chain rule for the last term of Equation 53 yields:
$\left[\log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})\right)^{-(\nu+D)/2}\right)\right]^{\prime}$ $=\frac{(D+\nu)}{2((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{\prime}$ (58)
$\left[\log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})\right)^{-(\nu+D)/2}\right)\right]^{\prime\prime}$ $=-\frac{(D+\nu)(((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{\prime\prime}-((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{\prime})^{2})}{2((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)^{2}}$ (59)
and plugging in the according partial derivatives in Equations 54 to 57 and applying integration by parts for the remainder of Equation 53, we have:
$\frac{\partial l}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)$ (60)
$\frac{\partial^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)(-2\bm{\mathrm{\Omega}}_{ij})-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2(z+\nu)^{2}}$ (61)
$\frac{\partial\ell}{\partial(\bm{\mathrm{A}}^{-1})_{ij}}$ $=\frac{1}{(\bm{\mathrm{A}}^{-1})_{ij}}\bm{1}_{i=j}+\frac{(D+\nu)}{2(z+\nu)}\left(2(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\sum_{m=1}^{M=i}(\bm{\mathrm{y}}_{m}-\bm{\mathrm{\mu}}_{m})(\bm{\mathrm{A}}^{-1})_{mj}\right)$ (62)
$\frac{\partial^{2}l}{\partial(\bm{\mathrm{A}}^{-1}){ij}^{2}}$ $=-\frac{1}{(\bm{\mathrm{A}}^{-1})_{ij}^{2}}\bm{1}_{i=j}-\frac{(D+\nu)\left((z+\nu)(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)^{2}}$ (63)
Where $\bm{1}$ is the indicator function for $i=j$, since the partial derivative of $\log\left(|\bm{\mathrm{A}}^{-1}|\right)$ are only relevant for the partial derivatives of the diagonal elements of $\bm{\mathrm{A}}^{-1}$. For the partial derivatives with respect to the degrees of freedom $\nu$, integration by parts yields:
$\frac{\partial l}{\partial\nu}$ $=-\frac{-\nu\operatorname{digamma(\frac{D+\nu}{2})}+D+\nu\operatorname{digamma}(\frac{\nu}{2})}{2\nu}+\frac{1}{2}\left(\frac{z(D+\nu)}{\nu(\nu+z)}-\log\left(\frac{(\nu+z)}{\nu}\right)\right)$ (64)
$\frac{\partial^{2}l}{\partial\nu^{2}}$ $=\frac{1}{4}\left(\frac{2k}{\nu^{2}}+\operatorname{trigamma(\frac{D+\nu}{2})}-\operatorname{trigamma}(\frac{\nu}{2})\right)+\frac{z(\nu z-D(2\nu+z))}{2\nu^{2}(\nu+z)^{2}}.$ (65)
For the low-rank approximation, we follow a similar notation. The LRA is given by $\bm{\mathrm{\Omega}}=\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}$ and hence the PDF is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{D}},\bm{\mathrm{V}},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}|^{1/2}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}$ (66)
and the log-likelihood is given by
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})|\right)+\\ \log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}\right)$ (67)
we follow a similar strategy as above and see that the partial derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and with respect to the degrees of freedom $\nu$ are the same as above:
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)$ (68)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)(-2\bm{\mathrm{\Omega}}_{ij})-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2(z+\nu)^{2}}$ (69)
$\frac{\partial l}{\partial\nu}$ $=-\frac{-\nu\operatorname{digamma(\frac{D+\nu}{2})}+D+\nu\operatorname{digamma}(\frac{\nu}{2})}{2\nu}+\frac{1}{2}\left(\frac{z(D+\nu)}{\nu(\nu+z)}-\log\left(\frac{(\nu+z)}{\nu}\right)\right)$ (70)
$\frac{\partial^{2}l}{\partial\nu^{2}}$ $=\frac{1}{4}\left(\frac{2k}{\nu^{2}}+\operatorname{trigamma(\frac{D+\nu}{2})}-\operatorname{trigamma}(\frac{\nu}{2})\right)+\frac{z(\nu z-D(2\nu+z))}{2\nu^{2}(\nu+z)^{2}}.$ (71)
where $z$ is now defined as
$z=(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}}).$ (72)
For the partial derivatives with respect to the elements of $\bm{\mathrm{D}}$, we note that the partial derivatives of the second term are given by:
$\frac{\partial}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{1}{2}\bm{\mathrm{\Sigma}}_{ii}$ (73)
$\frac{\partial}{\partial(\bm{\mathrm{D}}_{ii})^{2}}$ $=-\frac{1}{2}\left(\bm{\mathrm{\Sigma}}_{ii}\right)^{2}$ (74)
and the partial deriviatives of the third term are given by
$\frac{\partial}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{D+\nu}{2(z+\nu)}(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (75)
$\frac{\partial}{\partial\bm{\mathrm{D}}_{ii}^{2}}$ $=0$ (76)
and the second defaults to 0. The partial derivatives of the log-likelihood with respect to the elements of $\bm{\mathrm{D}}$ are hence given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{1}{2}\bm{\mathrm{\Sigma}}_{ii}-\frac{D+\nu}{2(z+\nu)}(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (77)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{D}}_{ii}^{2}}$ $=-\frac{1}{2}\left(\bm{\mathrm{\Sigma}}_{ii}\right)^{2}-\frac{D+\nu}{2(z+\nu)^{2}}\left((\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}\right)^{2}.$ (78)
For the partial derivatives with respect to the elements of $\bm{\mathrm{V}}$, we have a more complex formulation for the second term involving the determinant:
$\frac{\partial}{\partial V_{ij}}$ $=$ (79)
$\frac{\partial}{\partial V_{ij}^{2}}$ $=\bm{\mathrm{\Sigma}}_{ij}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{ii}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{qk}\bm{\mathrm{V}}_{kj}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{iq}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}$ (80)
and the partial derivatives of the third term are given by
$\frac{\partial}{\partial V_{ij}}$ $=\frac{D+\nu}{(z+\nu)}\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}$ (81)
$\frac{\partial}{\partial V_{ij}^{2}}$ $=\frac{D+\nu}{(z+\nu)^{2}}(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (82)
and hence integration by parts again gives us, similiar to the partial derivatives for the multivariate normal distribution in Equation 48 and 49:
$\frac{\partial\ell}{\partial\bm{\mathrm{V}}_{ij}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}+\frac{D+\nu}{(z+\nu)}\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}$ (83)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{V}}_{ij}^{2}}$ $=\bm{\mathrm{\Sigma}}_{ij}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{ii}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{qk}\bm{\mathrm{V}}_{kj}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{iq}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}$ (84)
$-\frac{D+\nu}{(z+\nu)^{2}}\left((\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}-\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}\right)^{2}\right)$
which concludes the derivation of the partial derivatives for the multivariate $t$-distribution $\blacksquare$