### v2

# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

Abstract
Probabilistic electricity price forecasting (PEPF) is vital for short-term electricity markets, yet the multivariate nature of day-ahead prices — spanning 24 consecutive hours — remains underexplored. At the same time, real-time decision-making requires methods that are both accurate and fast. We introduce an online algorithm for multivariate distributional regression models, allowing an efficient modelling of the conditional means, variances, and dependence structures of electricity prices. The approach combines multivariate distributional regression with online coordinate descent and LASSO-type regularization, enabling scalable estimation in high-dimensional covariate spaces. Additionally, we propose a regularized estimation path over increasingly complex dependence structures, allowing for early stopping and avoiding overfitting. In a case study of the German day-ahead market, our method outperforms a wide range of benchmarks, showing that modeling dependence improves both calibration and predictive accuracy. Furthermore, we analyse the trade-off between predictive accuracy and computational costs for batch and online estimation and provide an high-performing open-source Python implementation in the ondil package.
keywords:
online learning , multivariate distributional regression , probabilistic electricity price forecasting , LASSO regularization , day-ahead electricity market
[…]
1 Introduction
Short-term electricity markets play a key role in the integration of renewable energy sources and flexible generation in the electricity system. In Germany, the day-ahead auction is the major venue for physically delivered electricity. Trading volumes have grown with the increase of renewable generation capacity. To optimize decision-making and bidding strategies, market participants need accurate price forecasts. Since electricity prices are characterized by high volatility, positive and negative spikes and skewness, research and industry have moved towards probabilistic electricity price forecasting (PEPF) to account for their stochastic nature (see e.g. Nowotarski and Weron, 2018; Dexter Energy, 2024). With 24 prices per day, electricity prices are multivariate time series with a potentially complex dependency structure. However, the multivariate dimension of electricity price time series has received little attention for PEPF so far, while being of high importance for market participants in the context of the optimization of flexible assets and portfolio management (Löhndorf and Wozabal, 2023; Peña et al., 2024; Beykirch et al., 2022, 2024). At the same time, the increasing availability of high-frequency data and the need for real-time decision-making in energy markets require online estimation methods for efficient model updating. This work presents an online, multivariate distributional regression model, which we apply for probabilistic day-ahead electricity price forecasting in Germany. Our work is among the first to treat the 24-dimensional hourly electricity prices as multivariate distribution and the first to treat the problem in a strict online estimation setting, which makes the complex, high-dimensional distributional learning problem feasible on standard laptops. Our results show that modeling the dependence structure improves forecasting performance significantly compared to univariate approaches.
Figure 1: Time series plot for day-ahead electricity prices $P_{d,h}$ in Germany. In the left panel, each color corresponds to one delivery hour h=0,,23h=0,...,23. The blue dotted line marks the split between test and training data set. The gray area corresponds to the time of the right panel, which shows the same data along the dimension of the delivery hour, where each line represents a delivery day $d$. The high volatility, occasional positive and negative spikes and co-movement of electricity prices are visible.
Mean $\mu$ Std $\sigma$ MAD Min Max
Pre-2021 34.6 16.6 8.8 -130.1 200.0
2021 96.8 73.7 26.3 -69.0 620.0
2022 235.5 142.8 86.2 -19.0 871.0
2023 95.2 47.6 23.1 -500.0 524.3
Table 1: Summary statistics for day-ahead electricity prices $P_{d,h}$ in Germany for the years 2015 to 2023. The mean $\mu$, standard deviation $\sigma$, median absolute deviation (MAD), minimum and maximum are given.
The literature on PEPF has evaluated a wide range of different statistical and machine learning methods, such as quantile regression, ARX-GARCH models (Nowotarski and Weron, 2018; Billé et al., 2023; Marcjasz et al., 2023), conformal prediction methods, (see e.g. Kath and Ziel, 2021; Zaffran et al., 2022; Brusaferri et al., 2024a; Lipiecki et al., 2024), distributional regression and neural network approaches (e.g. Brusaferri et al., 2024b; Marcjasz et al., 2023; Hirsch et al., 2024; Ziel et al., 2021). However, these works treat each delivery hour as independent, univariate time series as in Ziel and Weron (2018). Let us motivate the need for multivariate probabilistic forecasting approaches for the day-ahead electricity price by two simple plots. Figure 1 shows a time series plot for the 24 hourly day-ahead electricity prices in Germany. The left panel shows each delivery hour as individual, daily series, emphasizing the daily co-movement. The right panel shows the cross-section, i.e. the daily shape for the first 180 days of 2017. The temporal correlations along the dimension of the delivery hours h=0,,23h=0,...,23 is clearly visible. We also like to highlight the high volatility during the winter 2021 and the Russian invasion of Ukraine 2022 and 2023. The long-run average prices have shifted from roughly 35 EUR/MWh in prior to 2021 to over 235 EUR/MWh in 2022. In the same reign, the standard deviation has increased from roughly 16 EUR/MWh to over 85 EUR/MWh. Summary statistics are given in Table 1, where the years 2021, 2022 and 2023 are shown separately to highlight the changes in the market. Additionally, Figure 2 shows the correlation matrix of the raw electricity prices (lower triangular) and the residual correlation for standard LASSO-ARX models for the electricity price (upper triangular). We see a strong, statistically significant remaining residual cross-correlation, indicating that the resulting marginal error distributions, which are conditional on the mean, are not independent. On top of the statistical motivation, Beykirch et al. (2022, 2024) clearly describe the need for predicting joint distributions for the optimization of schedules and bidding curves in energy markets, further examples are provided by Peña et al. (2024); Löhndorf and Wozabal (2023).
Figure 2: Correlation Matrix for day-ahead electricity prices $P_{d,h}$ in Germany. The lower triangle gives the Pearson hourly correlation $\rho$ for electricity prices. The upper triangle gives the hourly correlation of residuals $\varepsilon_{d,h}=P_{d,h}-\hat{\mu}_{d,h}$ for a standard LASSO-ARX model (see e.g. Nowotarski and Weron, 2018, and Eq. 15). The high degree of residual correlation, especially around the noon hours is clearly visible. All correlation coefficients are statistically significant to the $\alpha=0.01$ confidence level.
Work on multivariate probabilistic forecasting for day-ahead electricity prices are sparse in the literature and the majority of the existing works, e.g. Maciejowska and Nitka (2024); Berrisch and Ziel (2024); Han (2023); Mashlakov et al. (2021) and Agakishiev et al. (2025), do not evaluate multivariate scoring rules such as the VS, DSS or ES, but focus on the evaluation of the marginals of the multivariate distribution through the CRPS. This reduces the problem to modeling 24 marginal distributions, taking only lagged cross-information into account. To the best knowledge of the author, only two studies truly model and evaluate the multivariate dependence structure. First, Janke and Steinke (2020) approach the issue through implicit generative Copula models. Grothe et al. (2023) employ the Schaake shuffle, a post-processing method for point forecasts. On the contrary, in the fields of probabilistic weather, renewable production (Bjerregård et al., 2021; Sørensen et al., 2022; Kolkmann et al., 2024) and probabilistic load forecasting (Gioia et al., 2022; Browell et al., 2022) truly multivariate forecasting approaches have gained more attention.
The goal of distributional regression or “regression beyond the mean” (Kneib et al., 2023; Klein, 2024) is modeling not only the conditional expectation, but all distribution parameters of the assumed parametric response distribution conditional on explanatory variables. The most prominent model in this regard is the original Generalized Additive Model for Location, Scale and Shape (GAMLSS, Rigby and Stasinopoulos, 2005), of which numerous extensions have been developed over the last years (Kock and Klein, 2023; Kneib et al., 2023; Muschinski et al., 2022) and distributional deep neural networks (DDNN, e.g. Klein et al., 2021, 2023; Rügamer et al., 2024). Through the direct modeling of the variable’s distribution, this method is well suited for the generation of probabilistic forecasts and has been successfully applied in energy markets (Muniain and Ziel, 2020; Gioia et al., 2022; Serinaldi, 2011; Brusaferri et al., 2024b; Marcjasz et al., 2023). A drawback of fully distributional models is the computational effort and the need for efficient estimation approaches for (multivariate) distributional regression models has been recognized by Umlauf et al. (2025) and Gioia et al. (2025), who propose efficient batch estimation approaches.
For environments with large amounts of continuously incoming data, such as energy markets, online learning describes the task of updating the model given new data, without falling back on previous samples. Formally, in the strict online setting, after having seen $N$ samples of our data set, we fit a model, predict for step $N+1$. Subsequently, we receive the realized values for $N+1$ and update our model, taking into account only the new row $N+1$. This approach allows an efficient processing of high-velocity data and results in greatly decreased computational effort. The principle is outlined as well in Figure 6. Online learning for LASSO-regularized regression for the mean has been introduced in Angelosante et al. (2009, 2010) and Messner and Pinson (2019). Univariate approaches suitable for probabilistic forecasting based stochastic gradient descent have been developed for specific distributions, (see e.g. Pierrot and Pinson, 2021), conformal prediction (see e.g. Zaffran et al., 2022; Gibbs and Candes, 2021; Gibbs and Candès, 2024) and the generic online distributional regression in Hirsch et al. (2024). However, in the multivariate case, the literature remains sparse and focused on unconditional distributions and Copulae (see e.g. Dasgupta and Hsu, 2007; Zhao et al., 2022; Landgrebe et al., 2020).
We add to the literature by presenting a generic, online, regularized, multivariate distributional regression model, allowing to model all distribution parameters conditional on explanatory variables and validate the approach in a forecasting study for the day-ahead electricity market in Germany. Our paper is one the first to tackle the issue of truly multivariate probabilistic electricity price forecasting. The contribution of this paper is therefore threefold:
  • 1.
    Methodological Contribution: We develop a regularized online estimation method for multivariate distributional regression based on the univariate work by Hirsch et al. (2024). Our algorithm allows for two layers of regularization:
    • (a)
      By leveraging online coordinate descent and LASSO-type penalties for each individual distribution parameter, we allow for high-dimensional covariate spaces.
    • (b)
      By exploiting structure in scale matrix of the multivariate distribution, we develop a path-based estimation along increasing complex dependency structures, allowing for parsimonious estimation and early stopping. We validate the trade-off between model complexity and predictive accuracy in our application study.
    Our algorithm is generic and can be applied to any parametric multivariate distribution, as long as the (log-)likelihood function and its derivatives are available. We implement the multivariate normal and $t$-distributions with three different parameterizations of the scale matrix. Further distributions can be easily added.
  • 2.
    Applied Contribution: We apply the method to probabilistic forecasting of the joint distribution of spot electricity prices. We benchmark multivariate distributional regression models to LASSO models, conformal prediction, GARCH, univariate distributional regressions approaches, partly combined with Copula constructions.
    • (a)
      We provide the first comprehensive study on multivariate probabilistic forecasting of day-ahead electricity prices, evaluating the full multivariate distribution through proper scoring rules such as the Variogram score (VS), energy score (ES), Dawid-Sebastiani score (DSS) and Log-Score (LS) for a wide range of models and including the volatile years 2021 to 2023 in our test set.
    • (b)
      We show that the multivariate distributional regression, which allows modeling all distributional parameters, i.e. the mean, but also the dependence structure, conditional on explanatory variables such as renewable in-feed or past prices provide superior forecasting performance compared to modeling of the marginals only respectively keeping a static/unconditional dependence structure. Furthermore, we discuss the need for multivariate distributional forecasts for accurate prediction bands of the 24-hour path of day-ahead prices.
    • (c)
      We showcase the computational advantage of online estimation by benchmarking with repeated batch estimation on various re-estimation frequencies. These results provide an “efficient frontier” of computational costs against predictive accuracy.
  • 3.
    Software: We provide a high-performing Python implementation using just-in-time compilation and providing a familiar, scikit-learn-like API to facilitate the usage of our package for other researchers. We contributed our code to the ondil package by Hirsch et al. (2024).
Thereby, our contribution is valuable both from a methodological and applied perspective in research and industry alike. Reproduction code can be found on GitHub. [^1] See: https://github.com/simon-hirsch/online-mv-distreg.
The remainder of the paper is structured as usual: The following main Section 2 introduces the multivariate, online, regularized distributional regression model. Section 3 introduces the forecasting study, the used data and discusses multivariate forecast evaluation in detail. Section 4 presents our results and the computational costs. Section 5 concludes the paper.
2 Online Multivariate Distribution Regression
The following subsections introduce the building blocks for the multivariate, online distributional regression algorithm. The general structure of the algorithm and also the flow of the following section is outlined in Figure 3. First, we like to introduce some intuition to distributional regression. Distributional regression or “regression beyond the mean” (Kneib et al., 2023) aims at modeling not only the conditional expectation, but all distribution parameters of the assumed parametric response distribution conditioned on explanatory variables. Generally, we aim to model
$\bm{\mathrm{y}}\sim\mathcal{F}(\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K})\quad\text{where}\quad g_{k}(\bm{\mathrm{\theta}}_{k})=\bm{\mathrm{X}}_{k}\bm{\mathrm{\beta}}_{k}$
that is, we model the parameters $\bm{\mathrm{\theta}}_{k}$ of the distribution $\mathcal{F}$ of the response variable $\bm{\mathrm{y}}$ as regression based on the covariates in $\bm{\mathrm{X}}$. The link function $g(\cdot)$ ensures that the distribution parameters are in their domain. Naturally, this distributional model is aligned with our goal of probabilistic forecasting.
Figure 3: Online Distributional Regression. For univariate distributions (top panel), online distributional regression has been introduced by (Hirsch et al., 2024) by a combination of the iteratively reweighted least squares algorithm with online coordinate descent. We extend the approach to multivariate distributions by utilizing unconstrained scale matrix parameterizations and introducing a path-based regularization along increasingly complex dependency structures.
Section 2.1 gives a formal introduction to the univariate and multivariate case. To achieve an unconstrained estimation, we employ different parameterizations of the scale matrix in the multivariate case (Section 2.2). For the univariate case, a regularized online estimation has been introduced by Hirsch et al. (2024) based on a combination of the iteratively reweighted least squares algorithm for the estimation of distributional regression models (Section 2.3) and online coordinate descent (Section 2.4). We introduce our main contribution, the multivariate extension of this algorithm in Section 2.5 and further develop a path-based estimation along a sequence of the increasingly complex scale matrices, which allows for a regularized estimation and early stopping in Section 2.6.
We denote scalar float and integer values as lowercase letters (e.g. $a$), constants as large letters (e.g. $T$) vectors as bold, upright lower case letters (e.g. $\bm{\mathrm{v}}$) and matrices as bold upper case letters (e.g. $\bm{\mathrm{A}}$). The calligraphic $\mathcal{F}$ and $\mathcal{D}$ are reserved for (arbitrary) distributions, $\mathcal{N}$ denotes the normal distribution and $\mathcal{L}$ denotes the likelihood; other calligraphic letters (usually) denote index sets. Subscript values are usually indices in matrices, which we start with 0. Superscript indices (in square brackets) denote iterations and/or the number of samples received in the online setting.
2.1 Distributional Regression Setting
Starting from the univariate case, distributional regression aims to model the conditional distribution parameters of the response vector $\bm{\mathrm{y}}=(y_{1},...,y_{N})\in\mathbb{R}^{N\times 1}$, conditional on the covariate or explanatory data in $\bm{\mathrm{X}}\in\mathbb{R}^{N\times J}$ by adopting a parametric distribution $\bm{\mathrm{y}}\sim\mathcal{F}(\Theta)$, where $\mathbf{\Theta}=(\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K})$ is a tuple of $K$ distribution parameters $\bm{\mathrm{\theta}}_{k}=(\theta_{k1},...,\theta_{kN})$. Each of the distribution parameters are linked to the covariate data through a known, twice differentiable link function $g_{k}(\cdot)$, leading to:
$g_{k}(\bm{\mathrm{\theta}}_{k})=\bm{\mathrm{\eta}}_{k}={\bm{\mathrm{X}}_{k}}\bm{\mathrm{\beta}}_{k}$ (1)
where $\bm{\mathrm{\beta}}_{k}$ is the coefficient vector to be estimated, relating the $J_{k}$ covariates in the design matrix $\bm{\mathrm{X}}_{k}=(\bm{\mathrm{x}}_{k1},...,\bm{\mathrm{x}}_{kJ})^{\top}$ to the distribution parameter $\bm{\mathrm{\theta}}_{k}$ through the link function $g_{k}(\cdot)$. Hence, we have:
$y_{i}\sim\mathcal{F}(\theta_{1i},...,\theta_{Ki})\quad\text{and}\quad\theta_{ki}=g^{-1}\left(\beta_{k}\bm{\mathrm{x}}_{ki}\right)$ (2)
and the probability density function $f(y_{i}\mid\theta_{1i},...,\theta_{Ki})$. The distributional regression framework therefore allows the modeling of all distribution parameters as linear regression equations of the design matrices $\bm{\mathrm{X}}_{k}$, which can a subset or all of the available the covariate data $\bm{\mathrm{X}}$. Commonly additive models are employed, where $\bm{\mathrm{\eta}}_{k}=f_{k1}(\bm{\mathrm{x}}_{k1})+...+f_{kJ}(\bm{\mathrm{x}}_{kJ})$ where the functions $f_{kj}(\cdot)$ can be linear terms, but also non-linear effects such as B-splines (Klein, 2024; Stasinopoulos et al., 2024). Note that while the functions $f_{kj}(\cdot)$ might be non-linear, they can be represented by a combination of linear regression coefficients and $B$ basis functions $b(\cdot)$, i.e. $f_{kj}(\cdot)=\sum_{i=1}^{B}\beta_{kji}b_{i}(x_{kj})$. Rigby and Stasinopoulos (2005) introduce iteratively reweighted least squares (IRLS), maximizing the penalized likelihood, to estimate $\bm{\mathrm{\beta}}_{k}$. It is important to note here that in the frequentist estimation, the IRLS algorithm is agnostic to the actual estimation technique (see e.g. p. 113 in Stasinopoulos et al., 2024). Different flavors of LASSO-type regularized estimation approaches have been introduced by Groll et al. (2019); Muniain and Ziel (2020); Ziel et al. (2021); O’Neill and Burke (2023). A regularized, incremental estimation approach using online coordinate descent has been proposed by Hirsch et al. (2024), which will form the basis for the multivariate approach proposed in this paper.
Moving to the multivariate setting, we are interested in learning the conditional distribution parameters of the $D$-dimensional response variable $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{1},...,\bm{\mathrm{y}}_{D})$, conditional on the covariate data $\bm{\mathrm{X}}$, by adopting a multivariate parametric distribution $\bm{\mathrm{Y}}_{i}\sim\mathcal{F}_{D}(\mathbf{\Theta}_{i})$, where $\mathbf{\Theta}_{i}=(\bm{\mathrm{\theta}}_{i1},...,\bm{\mathrm{\theta}}_{iK})$ is a tuple of $K$ scalar, vector or matrix-valued distribution parameters. Each of the coordinates $m$ of the distribution parameter $\bm{\mathrm{\theta}}_{k}$ can again be related to its linear predictors by
$g_{km}(\bm{\mathrm{\theta}}_{km})=\bm{\mathrm{\eta}}_{km}=\bm{\mathrm{X}}_{km}\bm{\mathrm{\beta}}_{km}.$ (3)
This formulation is rather general. We note two important points here:
  • 1.
    The distribution parameters are subject to constraints. This especially holds for the scale (respectively precision) matrix, which is often required to be positive semi-definite. To ensure this holds, unconstrained parameterizations such as the (modified) Cholesky-decomposition or the low-rank approximation are often used (Pourahmadi, 2011; Muschinski et al., 2022; Salinas et al., 2019). The parameterization of the scale matrix will be discussed in Section 2.2.
  • 2.
    In practice, the different distribution parameters $\bm{\mathrm{\theta}}_{1},...,\bm{\mathrm{\theta}}_{K}$ can have many different shapes. Take, e.g. the multivariate $t$-distribution, parameterized using the Cholesky factor of the precision matrix $\bm{\mathrm{\Sigma}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{L}}$, denoted as $t_{D}(\bm{\mathrm{\theta}}_{1},\bm{\mathrm{\theta}}_{2},\bm{\mathrm{\theta}}_{3})\Leftrightarrow t_{D}(\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{\nu}})$. Then $\bm{\mathrm{\mu}}$ is a $N\times D$} matrix, $\bm{\mathrm{L}}$ is a $N\times D\times D$} cube (of which each vertical slice is a triangular matrix) and $\bm{\mathrm{\nu}}$ is a $N\times 1$ vector. Accordingly, the index set $\mathcal{M}_{k}$ of coordinates spans $\mathcal{M}_{1}=\{1,..,D\}$, $\mathcal{M}_{2}=\{(1,1),...,(D,D)\}$ and $\mathcal{M}_{3}=\{1\}$ and its cardinality is given by the product of the parameter’s dimension beyond $N$.
The general setting introduced here includes the Gaussian multivariate distributional regression introduced by Muschinski et al. (2022), the Copula-based multivariate distributional by Kock and Klein (2023) and the MCD-based additive covariance models by Gioia et al. (2022). The general estimation principle of repeatedly iterating through the distribution parameters until convergence translates to the multivariate case. However, in this setting we introduce an additional inner cycle through all coordinates of the currently active distribution parameter. The exact estimation algorithm will be introduced in Section 2.3 and the following Section 2.2 briefly discusses different options to parameterize the covariance respectively precision matrix.
2.2 Parameterization of the Precision Matrix
To ensure the positive definiteness of the scale matrix, we propose three unconstrained parameterizations. To save computational costs, we parameterize the distributions in terms of the inverse covariance matrix $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}$. This allows to avoid matrix inversion in the evaluation of the (log-) likelihood function. We briefly review first the (modified) Cholesky decompositions (CD resp. MCD), which have been proposed in the context of multivariate distributional regression by e.g. Pourahmadi (2011); Muschinski et al. (2022); Kock and Klein (2023). Additionally, we employ a low-rank approximation of the precision matrix, which has been used by Salinas et al. (2019) in the context of high-dimensional Gaussian processes and by März (2022) and O’Malley et al. (2023) in the context of distributional gradient boosted random forests.
The Cholesky-decomposition (CD) of the covariance matrix $\bm{\mathrm{\Sigma}}$ and the precision matrix $\bm{\mathrm{\Omega}}$ are defined as:
$\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ (4)
Muschinski et al. (2022) parametrize the normal distribution in terms of $\bm{\mathrm{A}}^{-1}$ and Kock and Klein (2023) choose $\bm{\mathrm{A}}$. Additionally the modified Cholesky-decomposition (MCD) can be used:
$\bm{\mathrm{\Sigma}}=(\bm{\mathrm{L}}^{-1})^{\top}\bm{\mathrm{D}}\bm{\mathrm{L}}^{-1}$ and $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ (5)
For the CD to yield a positive definite matrix, we require the diagonal of $\bm{\mathrm{A}}$ to be positive, which can be enforced by employing a log-link function. The lower diagonal of $\bm{\mathrm{A}}$ is unconstrained. The same holds for the MCD, where $\bm{\mathrm{D}}$ is a diagonal matrix with positive entries and $\bm{\mathrm{L}}$ is a unit lower triangular matrix with ones on the diagonal. Note that $\bm{\mathrm{A}}^{-1}=\bm{\mathrm{D}}^{-1/2}\bm{\mathrm{L}}$. The low-rank approximation (LRA) is defined as
$\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top},$ (6)
where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ and $\bm{\mathrm{V}}$ is a $D\times R$ matrix of rank $R$. The advantage of the LRA is that the dimensions of the parameters $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ scale linearly with the dimension $D$, however, the partial derivatives of the multivariate Gaussian and $t$-distribution with respect to the coordinates of $\bm{\mathrm{A}}$ and $\bm{\mathrm{V}}$ require inversion of the precision matrix. To ensure positive-definiteness for the LRA, we require the non-zero elements of $\bm{\mathrm{A}}$ to be positive, while $\bm{\mathrm{V}}$ is unconstrained. These requirements can easily be satisfied by choosing the log-link or the square root link function for $\bm{\mathrm{A}}$.
2.3 Iterative Reweighted Least Squares for Distributional Regression
Rigby and Stasinopoulos (2005) introduce iteratively reweighted least squares for the estimation of GAMLSS models. The RS algorithm consists of two nested loops, in which we cycle repeatedly through the distribution parameters and run a weighted fit of the score vector $\bm{\mathrm{z}}$ on the design matrix $\bm{\mathrm{X}}$ using the diagonal weight matrix $\bm{\mathrm{W}}$. The following paragraphs introduce the scoring vector and weights, the algorithm and the necessary modifications to move from a univariate case to the multivariate case.
The score vector is defined as
$\bm{\mathrm{u}}=\frac{\partial\ell}{\partial\eta},$ (7)
where $\ell$ is the log-likelihood $\ell=\log(\mathcal{L})$ and $\eta=g(\bm{\mathrm{\theta}})$ is the linked predictor. The working vector for the Newton-Raphson or Fisher-Scoring algorithm is defined as
$\bm{\mathrm{z}}=g\big(\hat{\bm{\mathrm{\theta}}}\big)+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1}\Leftrightarrow\bm{\mathrm{z}}=\bm{\mathrm{\eta}}+\frac{\partial\ell}{\partial\eta}\bm{\mathrm{W}}^{-1},$ (8)
where the weights are defined as:
$\bm{\mathrm{W}}=-\frac{\partial^{2}\ell}{\partial\eta^{2}}$ or $\bm{\mathrm{W}}=-\mathbb{E}\left[\frac{\partial^{2}\ell}{\partial\eta^{2}}\right]$ (9)
for Newton-Raphson and Fisher’s scoring respectively. Algorithm 1 shows the high-level algorithm for fitting distributional regression models using IRLS. To move from the univariate to the multivariate case, we need to cycle through all coordinates $m\in\mathcal{M}_{k}$ of the currently active distribution parameter $\bm{\mathrm{\theta}}_{k}$ in the inner loop and calculate the score vector $\bm{\mathrm{u}}_{km}$, working vector $\bm{\mathrm{z}}_{km}$ and weight matrix $\bm{\mathrm{W}}_{km}$ for each coordinate and run the weighted regression on $\bm{\mathrm{z}}_{km}$ on $\bm{\mathrm{X}}$. To allow for high-dimensional covariate spaces for each coordinate of the distribution parameter, we employ online coordinate descent and LASSO penalties for the weighted regression step, as introduced in Section 2.4. Section 2.5 will give a detailed view on the online version, also including the necessary details on multivariate distributions, model selection, and online updates.
1for outer iterations $i=0,$... until convergence do
2   Iterate through all distribution parameters.
3 forall $k\in\mathcal{K}$ do
4      Fit one distribution parameter.
5    for inner iterations $r=0,1,$... until convergence do
6         Evaluate three steps:
7
  1. 1.
    Evaluate $\bm{\mathrm{u}}_{k}$, $\bm{\mathrm{W}}_{k}$ and $\bm{\mathrm{z}}_{k}$ using Equations (7), (8) and (9).
  2. 2.
    Weighted regression of $\bm{\mathrm{z}}_{k}$ on $\bm{\mathrm{X}}_{k}$ using $\bm{\mathrm{W}}_{k}$ and yield $\hat{\bm{\mathrm{\beta}}}_{k}$
  3. 3.
    Evaluate convergence and end if converged.
8        -0.5em
9
10  Evaluate convergence and end if converged.
Output: $\widehat{\bm{\mathrm{\beta}}}_{k}$ and $\widehat{\bm{\mathrm{\Theta}}}=(\widehat{\bm{\mathrm{\theta}}}_{0},...,\widehat{\bm{\mathrm{\theta}}}_{p})$
Algorithm 1 High-level description of the IRLS algorithm for estimating distributional regression models (Rigby and Stasinopoulos, 2005; Stasinopoulos et al., 2024).
In the GLM, Fisher’s scoring and Newton-Raphson scoring coincide for the canonical link functions in the exponential family. However, for the scale and shape parameters, this is not necessarily the case anymore (for a detailed treatment of GLMs and estimation theory, see e.g. Lange et al., 2010). In the original GAMLSS, Rigby and Stasinopoulos (2005) use Fisher’s scoring. Our approach generally uses Newton-Raphson scoring for the multivariate case, since the derivation of the expected value of second derivatives can be intractable, especially for more complex parameterizations of the precision matrix. Newton-Raphson scoring requires the partial derivatives of the log-likelihood function with respect to the predictors. While many previous works on distributional regression employ Newton-Raphson scoring, each derive the partial derivatives for specific combinations of distribution function and link function only (see e.g. O’Neill and Burke, 2023; Muschinski et al., 2022). To facilitate the computational implementation in an mix-and-match fashion, we propose to use the first and second derivative of the log-likelihood with respect to the parameter and the first and second derivative of the link function and relate both to the necessary derivatives for Newton-Raphson scoring using the equalities given in the following Lemma 2.1, which allow for the simple utilization of arbitrary link functions and efficient calculation of working vector and weight matrices.
Lemma 2.1
Equipped with the first and second derivative of the log-likelihood with respect to the distribution parameter, ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$, as well as the first and second derivative of the link function $g(\cdot)$, we can retrieve the first and second derivative with respect to the predictor $\eta=g(\theta)$ as follows
$\frac{\partial\ell}{\partial\eta}$ $=\frac{\partial\ell}{\partial\theta}\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-1}\quad\quad\text{and}$ (10)
$\frac{\partial^{2}\ell}{\partial\eta^{2}}$ $=\left(\frac{\partial^{2}\ell}{\partial\theta^{2}}\frac{\partial g(\theta)}{\partial\theta}-\frac{\partial\ell}{\partial\theta}\frac{\partial^{2}g(\theta)}{\partial\theta^{2}}\right)\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-3}.$ (11)
The proof is straight-forward and utilizes the chain and quotient rules and can be found in A.2. We provide the necessary first and second partial derivatives of the log-likelihood with respect to the distribution parameter’s coordinates, ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$, for all parameters for the multivariate normal and multivariate $t$-distribution given in Table 2. The derivation can be found in A.3 and A.4.
Distribution Location Scale $\bm{\Sigma}$ resp. Precision $\bm{\Omega}$ Shape
Param. Dim. Param. Dim. Param. Dim.
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ $N\times\text{triangular}(D\times D)$ - -
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ $N\times\text{triangular}(D\times D)$, $N\times\text{diag}(D)$ - -
Multivariate Gaussian $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top}$ $N\times\text{diag}(D)$, $D\times r$ - -
Multivariate-$t$ $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$ $N\times\text{triangular}(D\times D)$ $\nu$ $N\times 1$
Multivariate-$t$ $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ $N\times\text{triangular}(D\times D)$, $N\times\text{diag}(D)$ $\nu$ $N\times 1$
Multivariate-$t$ $\bm{\mathrm{\mu}}$ $N\times D$ $\bm{\mathrm{\Omega}}=\bm{\mathrm{A}}+\bm{\mathrm{V}}\bm{\mathrm{V}}^{\top}$ $N\times\text{diag}(D)$, $D\times r$ $\nu$ $N\times 1$
Table 2: Overview of multivariate distributions and scale matrix parametrization (Param.) implemented in the paper and the respective dimensions (Dim.) for input data $\bm{\mathrm{Y}}$ of shape $N\times D$. Note that the number of parameters for the CD-based parameterization grows quadratically in $D$, but the LRA-based parameterizations grow linear in $D$ for fixed $r$.
2.4 Online Coordinate Descent for LASSO
Coordinate descent is the state-of-the-art method to estimate sparse and regularized regression problems of the form
$\bm{\mathrm{\beta}}=\arg\min_{\bm{\mathrm{\beta}}}\left\{\left\|\bm{\mathrm{y}}-\bm{\mathrm{X}}\bm{\mathrm{\beta}}\right\|_{2}+\lambda\|\bm{\mathrm{\beta}}\|_{1}\right\}$
where $\bm{\mathrm{X}}$ is the $N\times J$ design matrix, $\bm{\mathrm{y}}$ is the response variable, $\bm{\mathrm{\beta}}$ is the coefficient vector to be estimated and $\lambda$ is a parameter defining the strength of the regularization. Larger values of $\lambda$ lead to higher regularization. Angelosante et al. (2009, 2010) show that the problem can be reformulated using the Gramian matrices $\bm{\mathrm{G}}=\bm{\mathrm{X}}^{\top}\bm{\mathrm{W}}\bm{\mathrm{\Gamma}}\bm{\mathrm{X}}$ and $\bm{\mathrm{h}}=\bm{\mathrm{X}}^{\top}\bm{\mathrm{W}}\bm{\mathrm{\Gamma}}\bm{\mathrm{y}}$, potentially also accounting for weights $\bm{\mathrm{W}}=\operatorname{diag}(w_{1},...,w_{N})$ and exponential discounting $\bm{\mathrm{\Gamma}}=\operatorname{diag}((1-\gamma)^{N-1},....,(1-\gamma)^{1},(1-\gamma)^{0})$, where $\gamma\in(0,1)$ is a forget parameter. The LASSO problem can be solved by iteratively cycling through all coordinates $j\in J$ and solving
$\hat{\beta}_{j}\xleftarrow{}\frac{S\left(\bm{\mathrm{h}}\left[j\right]-\bm{\mathrm{G}}\left[j,:\right]\bm{\mathrm{\beta}}+\bm{\mathrm{G}}\left[j,j\right]\hat{\beta}_{j},\lambda\right)}{\bm{\mathrm{G}}\left[j,j\right]}$ (12)
where $S(x,\lambda)=\operatorname{sign}(x)\max(x-\lambda)$ is the so-called soft-threshold function. Coordinate descent is commonly solved on a decreasing grid of regularization strengths $\bm{\mathrm{\lambda}}$ on an exponential grid from $\lambda_{\text{max}}=\max{}\lvert\bm{\mathrm{G}}_{n+1}\rvert$. Algorithm 2 presents the full fitting process. The formulation can be easily extended to solve ElasticNet and Ridge Regression problems, however, in this work we restrict ourselves to the LASSO estimation, as the sparse solution allows for easy use of information criteria (IC) for model selection. The implementation in the ondil Python package supports the ElasticNet as well as various extensions such as box-constrained coefficients and early stopping. A more detailed treatment of online coordinate descent can be found in Messner and Pinson (2019) and Hirsch et al. (2024).
Input: New observations $\bm{\mathrm{x}}^{[n+1]},y^{[n+1]},w^{[n+1]}$ and stored $\bm{\mathrm{G}}^{[n]},\bm{\mathrm{h}}^{[n]}$.
1 Update $\bm{\mathrm{G}}^{[n+1]}=(1-\gamma)\bm{\mathrm{G}}^{[n]}+w_{n+1}(\bm{\mathrm{x}}^{[n+1]})^{\top}\bm{\mathrm{x}}^{[n+1]}$
2 Update $\bm{\mathrm{h}}^{[n+1]}=(1-\gamma)\bm{\mathrm{h}}^{[n]}+w_{n+1}(\bm{\mathrm{x}}^{[n+1]})^{\top}\bm{\mathrm{y}}^{[n+1]}$
3 Update $\lambda_{\text{max}}=\max{}\lvert\bm{\mathrm{G}}_{n+1}\rvert$ and initialize $\bm{\mathrm{\lambda}}$ as exponential grid.
4 for $\lambda\in\bm{\mathrm{\lambda}}$ do
5   Set starting coefficients $\bm{\mathrm{\beta}}_{\lambda}\leftarrow\bm{\mathrm{\beta}}_{\lambda[-1]}$
6 while not converged do
7    forall $j\in 1,...,J$ do
8         Update $\hat{\beta}_{j,\lambda}$ according to Equation 12
9
10     Check convergence for $\widehat{\bm{\mathrm{\beta}}}_{n+1,\lambda}$ and proceed to next $\lambda$ if converged.
11
12
[…]
Algorithm 2 Online LASSO, see Angelosante et al. (2010) and Messner and Pinson (2019)
2.5 Online Estimation Algorithm
As outlined in Section 2.3 and Algorithm 1, the IRLS algorithm consists of two nested loops. In the outer loop, we iterate through all distribution parameters. In the inner loop, we repeatedly run a weighted fit of the score vector $\bm{\mathrm{u}}$ on the design matrix $\bm{\mathrm{X}}$ using the weights $\bm{\mathrm{W}}$ until convergence. Note that in the inner loop, we run the weighted fit sequentially for all elements of the distribution parameter. Since the fit itself is agnostic to the regression technique (Stasinopoulos et al., 2024), we employ the online coordinate descent-based LASSO estimation here, as it has been proposed by Hirsch et al. (2024) for the univariate case already. Algorithm 3 gives an overview on the online estimation of multivariate distributional regression models. While its general structure follows the high-level overview in Algorithm 1, we provide a detailed treatment on the online update using OCD and the implications due to the multivariate case. We define the index sets $\mathcal{K}=\{1,...,p\}$ for the number of parameters and $\mathcal{M}_{k}=\{1,..,M_{k}\}$ for the number of elements of each parameter as described in Section 2.4.
Input: $\bm{\mathrm{y}}^{[n+1]},\bm{\mathrm{X}}_{k,m}^{[n+1]}$ and the stored Gramian matrices $\bm{\mathrm{G}}_{km}^{[n]},\bm{\mathrm{h}}_{km}^{[n]}$.
1 Initialize the fitted values $\widehat{\bm{\mathrm{\theta}}}_{km}^{[n+1,0,0]}=\hat{\bm{\mathrm{\beta}}}_{km}^{[n]}(\bm{\mathrm{X}}_{k,m}^{[n+1]})^{\top}$ for $k,m\in\mathcal{K}\times\mathcal{M}$.
2 Evaluate the linear predictors $\widehat{\bm{\mathrm{\eta}}}^{[n+1,0,0]}_{km}=g_{km}(\hat{\theta}_{km}^{[n+1,0,0]})$} for $k,m\in\mathcal{K}\times\mathcal{M}$.
3 for $i=0,$... until convergence do
4 forall $k\in\mathcal{K}$ do
5      Start the inner cycle and iterate over all elements of the distribution parameter.
6    for $r=0,1,$... until convergence do
7       forall $m\in\mathcal{M}_{k}$ do
8            Evaluate $u_{km}^{[n+1,i,r]}$, $w_{km}^{[n+1,i,r]}$ and $z_{km}^{[n+1,i,r]}$ using Equations (7), (8) and (9).
9            Update $\bm{\mathrm{G}}_{km}^{[n+1,i,r]}\leftarrow\gamma\bm{\mathrm{G}}_{km}^{[n]}+w_{km}^{[n+1,i,r]}\left((\bm{\mathrm{X}}_{km}^{[n+1]})^{\top}(\bm{\mathrm{X}}_{km}^{[n+1]})\right)$
10            Update $\bm{\mathrm{h}}_{km}^{[n+1,i,r]}\leftarrow\gamma\bm{\mathrm{h}}_{km}^{[n]}+w_{km}^{[n+1,i,r]}\left((\bm{\mathrm{X}}_{km}^{[n+1]})^{\top}z_{km}^{[n+1,i,r]}\right)$
11            Update $\hat{\bm{\mathrm{\beta}}}_{km\lambda}^{[n+1,i,r+1]}\leftarrow\hat{\bm{\mathrm{\beta}}}_{km\lambda}^{[n]}$ based on $\bm{\mathrm{G}}_{km}^{[n+1,i,r]}$ and $\bm{\mathrm{h}}_{km}^{[n+1,i,r]}$ using the online LASSO (see Algorithm 2) or recursive least squares.
12            Select the optimal $\lambda$ using IC and set $\widehat{\bm{\mathrm{\beta}}}_{km}^{n+1,i,r+1]}\leftarrow\hat{\bm{\mathrm{\beta}}}_{km\lambda^{\text{opt}}}^{[n+1,i,r+1]}$.
13            Calculate the updated $\hat{\bm{\mathrm{\eta}}}_{km}^{[n+1,i,r+1]}$ and $\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i,r+1]}$
14            Evaluate the convergence.
15
16     End the inner cycle on the convergence of $\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i,r]}$.
17      Set $\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i+1,0]}\leftarrow\hat{\bm{\mathrm{\beta}}}_{km}^{[n+1,i,r]}$ and set $\hat{\bm{\mathrm{\eta}}}_{km}^{[n+1,i+1,0]}\leftarrow\hat{\bm{\mathrm{\eta}}}_{km}^{[n+1,i,r]}$ and set $\hat{\bm{\mathrm{\theta}}}_{km}^{[n+1,i+1,0]}\leftarrow\hat{\bm{\mathrm{\theta}}}_{km}^{[n+1,i,r]}$.
18      End the outer cycle if the change in the penalized likelihood is sufficiently small.
19
Output: $\widehat{\bm{\mathrm{\beta}}}_{k,n+1}$ and $\widehat{\bm{\mathrm{\Theta}}}^{[n+1]}=(\widehat{\bm{\mathrm{\theta}}}_{0}^{[n+1]},...,\widehat{\bm{\mathrm{\theta}}}_{p}^{[n+1]})$ and the updated $\bm{\mathrm{G}}_{km}^{[n+1]}$ and $\bm{\mathrm{h}}_{km}^{[n+1]}$.
Algorithm 3 Online regularized multivariate distributional regression.
For each inner iteration $i$, the update of the Gramian matrices starts at the Gramian matrices of $\bm{\mathrm{G}}_{km}^{[n]}$ and $\bm{\mathrm{h}}_{km}^{[n]}$ and the new information enters the Gramian matrices through the update of the weights $\bm{\mathrm{W}}$ and the working vector $\bm{\mathrm{z}}$. However, the weights are also updated iteratively along each inner and outer iteration $i$ and $r$ due to the Newton-Raphson step towards the optimal coefficients. The weights can only be updated for the current update step $n+1$, while previous weights remain fixed. In a pure batch case, all weights are updated within each Newton-Raphson step. This introduces an approximation error for the online case, which can be controlled by the forget parameter $\gamma$ as shown in Hirsch et al. (2024).
For each element of the distribution parameter, we estimate a regularization path. This raises the issue of model selection, i.e. the selection of the optimal regularization parameter $\lambda_{mk}^{\text{opt}}$. We propose to use information criteria (IC), as it is well-aligned to the likelihood-based framework of distributional regression. Define a generalized IC as
$\operatorname{IC}=-2\ell\left(\bm{\mathrm{Y}}\mid\widehat{\bm{\mathrm{\Theta}}}\right)+\nu_{0}K+\nu_{1}K\log(N)+\nu_{2}K\log\left(\log(N)\right)$ (13)
where $\ell$ is the log-likelihood under the model, $K$ is the number of parameters in the model and $N$ the number of seen observations. We can recover Akaikes Information Criterion (AIC), the Bayesian Information Criterion (BIC), and the Hannan-Quinn Information Criterion (HQC) by setting $\nu_{0},\nu_{1},\nu_{2}$ accordingly. The optimal regularization parameter is then selected as $\lambda_{mk}^{\text{opt}}=\operatorname{argmin}_{\lambda}\operatorname{IC}$. Since the evaluation of the likelihood can be costly for high-dimensional data, we propose to employ the first derivative of the log-likelihood, i.e. calculate
$\ell\left(\bm{\mathrm{Y}}\mid\widehat{\bm{\mathrm{\Theta}}}^{[\lambda_{i}]}\right)\approx\ell\left(\bm{\mathrm{Y}}\mid\widehat{\bm{\mathrm{\Theta}}}^{[\lambda_{0}]}\right)+\frac{\partial\ell}{\partial\theta}\left(\widehat{\theta}_{km}^{[\lambda_{0}]}-\widehat{\theta}_{km}^{[\lambda_{i}]}\right)$ (14)
where the superscript $\lambda_{i}]$} denotes the model with the regularization parameter $\lambda_{i}$. The approximation is valid for small changes in the regularization parameter and avoids the costly re-evaluation of the likelihood.
The algorithm goes iteratively along all coordinates of the distribution parameter. The coordinates of the distribution parameters might impact each other, e.g. in the matrix multiplication of the CD-based scale matrix (see also the definition of the derivatives in A.3 and A.4). At the same time, we initialize the fitted values $\hat{\theta}_{m}$ as constant values. To stabilize the estimation, we propose to update the values in the very first iteration $i$ by a “dampened” version, i.e. taking
$\hat{\eta}_{m}^{[0,i]}\leftarrow g_{m}^{-1}\left((i+1)\hat{\theta}_{m}^{[0,i]}+\hat{\theta}_{m}^{[0,i-1]})/(i+1)\right)$
Hence, the predictions from the first iteration will be the average of the first fitted values and the initialization. This feature is mainly important for the scale matrix, whose coordinates are usually not orthogonal and less so for the location and (scalar) tail parameters.
Since the partial derivatives are not information orthogonal, the options for parallelization remain limited unfortunately. For the multivariate normal and $t$-distribution used in this paper, only the estimation of the location parameter can be parallelized, as well as the estimation of the coordinates of the LRA matrix $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ for the normal distribution. For the $t$-distribution, the estimation of $\bm{\mathrm{A}}$ can only be parallelized for sufficiently high degrees of freedom. One plausible option for parallelization with non-orthogonal parameters would be using a step size smaller than one, i.e. using a convex combination of the previous and the newly estimated coefficients. However, this would introduce an additional hyperparameter. As parallelization would incur further open questions with respect to individual or joint regularization and model selection and the location parameter generally converges rather fast, we have not implemented parallel computation yet and leave it for future research.
2.6 Path-based Regularized Estimation for the Scale Matrix
Using LASSO-type regularization, the algorithm can handle high-dimensional covariate spaces for each coordinate of the distribution parameter. However, for high-dimensional response variables, i.e. large $D$, the number of parameters in the scale matrix grows quadratically in $D$ for the CD-based parameterizations and linearly in $D$ for the LRA-based parameterization (see Table 2). To alleviate this issue and allow for parsimonious modelling for large $D$, we propose a path-based estimation approach that starts with a simple (highly regularized) structure of the scale matrix and gradually increases its complexity. This approach is inspired by regularization techniques in time series analysis and path-based estimation methods in high-dimensional statistics (such as the graphical LASSO, Friedman et al., 2008). We exploit that in many cases, some structure can be imposed on the scale matrix, i.e. in spatial or temporal data, which has a clear dependence pattern along the diagonal. In these cases, the scale matrix can be regularized by systematically setting off-diagonal elements to zero (Gabriel, 1962; Zimmerman and Núñez-Antón, 1997; Zimmerman et al., 1998). While both, the CD-based and the LRA-based scale matrix parametrization lend themselves to this type of regularization, the approach is mainly popular with the Cholesky-based parameterizations due to the relationship between the elements of the CD and the temporal correlation for longitudinal data under the name AD-$r$ regularization. However, such regularization is commonly applied a-priori and not in a data-driven fashion, see e.g. Muschinski et al. (2022); Zimmerman and Núñez-Antón (1997). On the other side, in coordinate descent estimation of regularized problems such as (graphical) LASSO, path-based estimation starting from a strongly regularized solution towards an (almost) not regularized solution has proven itself as an efficient solution approach. In this section, we aim to combine these two principles by introducing path-based estimation for the regularized scale matrix.
On a high level, our algorithm starts with an “independence-parameterization” of the scale matrix and subsequently adds more non-zero elements and thus complexity to the parameterization of the scale matrix. Figure 4 illustrates how the path-based estimation uses increasingly complex specifications for the scale matrix $\bm{\mathrm{\Sigma}}$ respectively $\bm{\mathrm{\Omega}}$. Formally, for some regularization parameter $\alpha$, we set the elements of the scale to zero for
  • 1.
    the Cholesky-based parameterizations if the indices of the diagonal matrices $\bm{\mathrm{A}}$ or $\bm{\mathrm{L}}$, $i,j$ are such that $|i-j|>\alpha$,
  • 2.
    the LRA-based parameterization if the indices $d,r$ of $\bm{\mathrm{V}}$ are such that $r\geq\alpha$,
and present the schematic overview in Algorithm 4. Note that we can use warm-starting for all previously fitted elements of the scale matrix, however, due to the non-orthogonality of the elements, we need to re-estimate all elements of the scale matrix in each iteration.
[…]
Figure 4: Path-based estimation along increasingly complex scale matrix parameterizations. The top panel shows the AD-$r$ regression for a Cholesky-based parameterization $((\bm{\mathrm{L}}^{[\alpha]})^{-1})^{\top}(\bm{\mathrm{L}}^{[\alpha]})^{-1}$. The superscript $\alpha]$ denotes the iteration regularization. Black elements are the state after the initial fit assuming independence. Red elements of the vectors are added in iteration $\alpha$. Blue elements have been added in previous iterations. The lower panel shows the estimation along the LRA-based parameterization $\bm{\mathrm{A}}+(\bm{\mathrm{V}}^{[\alpha]})(\bm{\mathrm{V}}^{[\alpha]})^{\top}$, where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{d})$ is not regularized and the $D\times r$ matrix $\bm{\mathrm{V}}$ is filled column-wise with non-zero elements. Own Illustration.
1for $\alpha=0,...,D$ do
2   Fit the online distributional regression Algorithm 3 for regularization level $\alpha$.
3   Evaluate the log-likelihood for the current regularization level $\alpha$.
4 Early Stop if the log-likelihood (or information criteria) does not increase sufficiently.
5
Output: Estimates for all $\alpha$.
Algorithm 4 Path-based scale-regularization for online multivariate distributional regression.
Note that both approaches can be used for parameterizations using the covariance and the precision matrix. The path-based estimation allows for re-using the previous iterations’ coefficients to achieve fast convergence in the OCD. For the CD-based parameterization, we increasingly add more off-diagonals to the lower-diagonal matrix. For the LRA-based parametrization, we add more and more columns to the low-rank matrix $\bm{\mathrm{V}}$. Let us note a few observations:
  • 1.
    For a (small) fixed maximum regularization size, the number of parameters in the CD and MCD-based parameterization grows (almost) linearly in $D$, alleviating the disadvantage of quadratic complexity.
  • 2.
    For the multivariate $t$-distribution, independence is only achieved as $\nu\rightarrow\infty$. We therefore set a high initial guess ($\nu=10^{6}$) for the first outer iteration of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{\Omega}}$ to ensure numerical stability for the first iteration and subsequently choose a lower initial guess for the first iteration of $\nu$, since the Newton-Raphson algorithm relies on appropriate start values and tends to alternate between extrema otherwise (see e.g. Casella and Bachmann, 2021; Kornerup and Muller, 2006, on the impact of initial values for Newton-Raphson algorithms.) [^2] We have found the algorithm to iterate between $\nu=2$ and $\nu>10^{10}$ for too large start values for the degrees of freedom. The proposed approach however has proved stable through the full simulation study with highly volatile electricity prices.
  • 3.
    We can employ the path-based estimation to early stop the estimation, if the log-likelihood or an information criterion does not increase sufficiently by adding more non-zero elements. This allows for both, implicit regularization and decreased estimation time. However, once we early stop, we cannot increase the complexity of the parametrization in the online estimation but need to treat this as fixed.
Currently, the Algorithm will add only full off-diagonals (for Cholesky-based approaches) respectively columns (LRA). The implementation however could also work for block-wise schemes (see e.g. the adaptive block structure in Cai and Yuan, 2012) or user-defined regularization patterns. The development of smart selection schemes for the next coordinates of the covariance matrix to include would be beneficial for the speed of the algorithm. We provide an analysis of the in-sample selection of $\alpha$ and the out-of-sample performance for various $\alpha$ in the forecasting study in Section 4.3 and leave the development of more advanced selection schemes for future research.
3 Forecasting Study
3.1 Day-ahead Electricity Market and Data
For electricity produced on day $t$ and hour $h$, the short-term electricity market in Germany is split in three major parts: The daily day-ahead auction on t1t-1 at 12:00 hours for 24 hourly delivery periods $h\in\{0,...,23\}$, the afternoon auction with quarter-hourly delivery periods on t1t-1, at 15:00 hours and the continuous intraday market. The daily procedure for the day-ahead auction, which is the focus of this paper, is shown in Figure 5. The market is organized by EPEX SPOT and Nordpool in the joint single day-ahead coupling (SDAC) as a pay-as-cleared auction through the EUPHEMIA algorithm, resembling the merit-order model for the electricity market (Billé et al., 2023; Hirsch et al., 2024; Viehmann, 2017).
Auction t1t-1, 12:00012Delivery periods h=0,,23h=0,...,23 on day tt222223Auction $t$, 12:00012Delivery periods h=0,,23h=0,...,23 on day t+1t+1222223Auction $t+1$, 12:00012Delivery periods h=0,,23h=0,...,23 on day t+2t+2222223Results at 12:42, t1t-1Results at 12:42, $t$
Figure 5: Structure of the day-ahead electricity market in Germany. Own illustration based on information on EPEX SPOTs website and Viehmann (2017).
At each day $t$, before the day-ahead auction at 12:00, we aim to generate forecasts for day $t+1$. Prior to forecasting, we update our models by taking into account the realized prices and forecasts for delivery day $t$. Figure 6 contrasts online learning with the expanding and rolling window batch learning, two popular schemes for forecasting studies. Note that in the online learning scheme, we only use the new observation for updating the model, while in the batch learning schemes, we re-use all or a fixed window of previous observations. The rolling window scheme is popular in the EPF community (see e.g. Lago et al., 2021; Nowotarski and Weron, 2018) while online learning schemes are a rather recent development in EPF (e.g. Berrisch and Ziel, 2024; Zaffran et al., 2022; Hirsch et al., 2024).
Initial set i=0,,Ni=0,...,NBatch Expanding WindowDataset i=0,,N+1i=0,...,N+1Dataset i=0,,N+2i=0,...,N+2Dataset i=0,,N+3i=0,...,N+3Dataset i=0,,N+4i=0,...,N+4Initial set i=0,,Ni=0,...,NBatch Rolling WindowDataset i=1,,N+1i=1,...,N+1Dataset i=2,,N+2i=2,...,N+2Dataset i=3,,N+3i=3,...,N+3Dataset i=4,,N+4i=4,...,N+4Initial set i=0,,Ni=0,...,NOnline LearningNN+1NN+2New observations added row-by-row.NN+3NN+4
Figure 6: Repeated Batch Learning vs. Online Learning for the forecasting study. Own illustration.
We use the same data set as in Lipiecki et al. (2024), which consists of electricity prices for the German day-ahead market from 2015-01-01 to 2024-01-01 exclusive. In line with previous works, we use the data until 2018-12-26 as initial training set, leaving 1831 observations and therefore more than 4 years, as it is best practice (Lago et al., 2021), for out-of-sample testing. This dataset has been employed in various studies and to enable comparability, we further split the dataset in two sub-samples, that is 2018-12-27 to 2020-12-31 (736 days) and 2021-01-01 to 2023-12-31 (1095 days). Additionally, the data set contains day-ahead renewable production forecasts, load forecasts and prices for fundamental commodities. All features are briefly described in Table 3. We use an incremental mean-variance scaling to post-process the electricity prices. We denote the electricity price for day $t$ and hour $h\in 0,...,23$ as $P_{th}$. We have $y_{th}=(P_{th}-\tilde{\mu}_{th})/\tilde{\sigma}_{th}$, where $\tilde{\mu}_{th}=1/t\sum_{i}^{t}P_{ih}$ and $\tilde{\sigma}_{th}=\sqrt{1/t\sum_{i}^{t}(P_{ih}-\tilde{\mu}_{ih})^{2}}$ are the mean and standard deviation up to observation $t,h$ and therefore have $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{23})$ as the 24-dimensional ($D=H=24$) response matrix. We find that this stabilizes the estimation of covariance matrices and the normalization can be re-applied after the estimation, i.e. the covariance of $\bm{\mathrm{p}}_{t}$ corresponds to $\operatorname{diag}(\tilde{\bm{\mathrm{\sigma}}}_{t})\bm{\mathrm{\Sigma}}\operatorname{diag}(\tilde{\bm{\mathrm{\sigma}}}_{t})$, where $\bm{\mathrm{\Sigma}}$ is estimated based on $\bm{\mathrm{y}}_{t}$. Incremental updates of the mean and variance are straight-forward in an online learning setting using Welford’s method (Welford, 1962). Note that, in this application study, we have $T$ corresponding to $N$ and $H$ corresponding to $D$ in the general notation.
Variable Description Resolution Source
$\operatorname{ResLoad}_{t,h}$ Day-ahead residual load forecast Hourly ENTSO-E
$\overline{\operatorname{ResLoad}}_{t}$ Day-ahead baseload residual load forecast $\frac{1}{H}\sum_{h=1}^{H}\operatorname{ResLoad}_{t,h}$ Daily ENTSO-E
$\operatorname{EUA}_{t}$ EU emission allowances Daily Refinitiv
$\operatorname{Gas}_{t}$ Natural gas prices Daily Refinitiv
$\operatorname{Coal}_{t}$ Coal prices Daily Refinitiv
$\operatorname{Oil}_{t}$ Oil prices Daily Refinitiv
$\operatorname{WD}_{t}$ Weekday dummies Daily Calender
Table 3: Variables from the data set of Marcjasz et al. (2023) and Lipiecki et al. (2024).
3.2 Model Definition and Benchmarks
We propose modeling the multivariate distribution of the day-ahead electricity prices in increasing complexity. If possible, models are updated online, i.e. using only the new data for each day. We differentiate between an adaptive estimation, which is updating a single, unconditional distributional parameter and the full conditional estimation linking the distribution parameter to explanatory variables.
Figure 7: Model Taxonomy. We differentiate between modelling the marginal distribution and the dependence structure.
Figure 7 shows the increasing model complexity. We start with the established LARX models and naively estimate the unconditional residual distribution (denoted as $\operatorname{LARX}+\mathcal{N}(0,\sigma)$ and $\operatorname{LARX}+\mathcal{N}(0,\Sigma)$). Additionally, we employ two additional univariate benchmarks. First, we use a conformal prediction (CP) approach, which is based on the LARX model and uses a combination of the split-conformal prediction (SCP) and adaptive conformal prediction (ACI, see Gibbs and Candes, 2021; Gibbs and Candès, 2024). We use absolute residual scores and a calibration set of 200 data points. Note that conformal prediction is generally a univariate post-processing method, since it is centered on issuing prediction intervals and therefore not suitable for the generation of ensemble forecasts. Multivariate conformal prediction approaches exists, but are in their infancy and focussed on a notion of multivariate quantiles. [^3] As a simple workaround, we have tried to combine CP with a copula-based approach using a PIT transformation. However, the conformal predictive density approximated using 199 quantiles was not sufficient to generate useful ensembles in the inverse transformation, i.e. the step from the simulated copula on $\mathcal{U}(0,1)$ to the original domain failed. Second, we use a GARCH(1,1) model with a Gaussian distributional assumption (denoted as LARX+GARCH(1,1)). We estimated the GARCH model on the residuals of the LARX model using the arch package in Python (Sheppard et al., 2024). We increase complexity by moving to full distributional model for the marginals and adding an adaptive estimation of the dependence structure using the Gaussian copula (denoted as oDistReg+Copula). Lastly, we estimate the full multivariate distribution in a conditional way using the proposed multivariate online distributional regression approach (denoted as oMvDistReg($\mathcal{F}$, parameterization, method)). We describe the full model in the following and note that we additionally describe the hyper parameters in A.5. For all three complexity levels, we include a reference model that assumes independence to showcase the value-add of modelling the dependence structure. We model the mean/location for all regression models, also for the LARX models , by
$g_{\mu}(\mu_{t,h})$ $=\beta_{\mu,0,h}+\sum_{l=1}^{L=7}\beta_{\mu,l,h}y_{t-l,h}+\sum_{i\in\{0,...,23\}\setminus h}\beta_{\mu,8+i,h}y_{t-1,i}+\sum_{w=1}^{W=6}\beta_{\mu,30+w,h}\operatorname{WD}_{t,h}$ (15)
$+\beta_{\mu,37,h}\operatorname{min}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,38,h}\operatorname{max}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,39,h}\operatorname{Q}_{10}(\bm{\mathrm{y}}_{t-1})+\beta_{\mu,40,h}\operatorname{Q}_{90}(\bm{\mathrm{y}}_{t-1})$
$+\beta_{\mu,41,h}\operatorname{ResLoad}_{t,h}+\beta_{\mu,42,h}\operatorname{EUA}_{t}+\beta_{\mu,43,h}\operatorname{Gas}_{t}+\beta_{\mu,44,h}\operatorname{Coal}_{t}+\beta_{\mu,45,h}\operatorname{Oil}_{t}.$
We model the scale parameters for univariate distributional models, as well as the elements of the Cholesky-factor $\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$, and the elements of the diagonal matrix $\bm{\mathrm{A}}$ in the LRA-based scale matrices by
$g_{\theta}(\theta_{t,h,h})$ $=\beta_{\theta,0,h}+\beta_{\theta,1,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)^{-1}+\beta_{\theta,2,h}\operatorname{ResLoad}_{t,h}$ (16)
$+\beta_{\theta,4,h}\operatorname{EUA}_{t}+\beta_{\theta,5,h}\operatorname{Gas}_{t}+\beta_{\theta,6,h}\operatorname{Coal}_{t}+\beta_{\theta,7,h}\operatorname{Oil}_{t},$
where $\operatorname{SignedSquare}(a)=\operatorname{sign}(a)\sqrt{|a|}$ is the signed square root and $\bm{\mathrm{\Sigma}}^{[t-1:t-7]}$ is the rolling empirical covariance matrix of $\bm{\mathrm{y}}_{t}$ for the last 7 days. Note that we use the inverse of the signed square, since we are working on the preicision matrix. For the univariate models, we replace this accordingly with the empirical rolling standard deviation. For the LRA-based parameterization, we choose $r=2$ and model the elements of $\bm{\mathrm{V}}$ as
$g_{v}(v_{t,h,0})$ $=\beta_{v,0,h}+\beta_{v,1,h}\operatorname{SignedSquare}\left(\bm{\mathrm{\Sigma}}_{h,h}^{[t-1:t-7]}\right)+\beta_{v,2,h}\overline{\operatorname{ResLoad}}_{t}$ (17)
$+\beta_{v,3,h}\operatorname{EUA}_{t}+\beta_{v,4,h}\operatorname{Gas}_{t}+\beta_{v,5,h}\operatorname{Coal}_{t}+\beta_{v,6,h}\operatorname{Oil}_{t},$
$g_{v}(v_{t,h,1})=\sum_{w}^{W=6}\beta_{v,14+w,h}\operatorname{WD}_{t,h}$ (18)
that is, the first rank takes most of the fundamental variables, while the second rank contains the weekday binary variables. The degrees of freedom are modeled as
$g_{\nu}(\nu_{t})$ $=\beta_{\nu,0}+\beta_{\nu,1}\operatorname{mean}(\bm{\mathrm{y}}_{t-1})+\sum_{w}^{W=6}\beta_{\nu,1+w,h}\operatorname{WD}_{t,h}+\beta_{\nu,8}\overline{\operatorname{ResLoad}}_{t}$ (19)
$+\beta_{\nu,9}\operatorname{EUA}_{t}+\beta_{\nu,10}\operatorname{Gas}_{t}+\beta_{\nu,11}\operatorname{Coal}_{t}+\beta_{\nu,12}\operatorname{Oil}_{t}.$
The univariate models are therefore a slight simplification compared to the models used in Hirsch et al. (2024), however thereby the multivariate distributional regression models and the Copula-based approaches are better comparable. Lastly, let us remark on the online tracking of the Gaussian copula. The probability density function (PDF) for the Gaussian copula is given by:
$\ell(\bm{\mathrm{u}}\mid\bm{\mathrm{\Sigma}})=\frac{1}{|\tilde{\bm{\mathrm{\Sigma}}}|^{1/2}}\exp\left(-\frac{1}{2}\bm{\mathrm{n}}^{\top}(\tilde{\bm{\mathrm{\Sigma}}}^{-1}-\bm{\mathrm{I}})\bm{\mathrm{n}}\right)\prod_{d=0}^{D}p(y_{d}\mid\bm{\mathrm{\theta_{d}}})$ (20)
where $\bm{\mathrm{u}}$ are the pseudo-observations on the $\mathcal{U}(0,1)$ space, $\bm{\mathrm{n}}=\Phi^{-1}(\bm{\mathrm{u}})$, $\Phi$ is the CDF of the standard normal distribution and $\tilde{\bm{\mathrm{\Sigma}}}$ is the covariance matrix $\bm{\mathrm{\Sigma}}$ scaled to the correlation matrix, $\bm{\mathrm{I}}$ is the identity matrix and $p(x_{d}\mid\bm{\mathrm{\theta}}_{d})$ is the likelihood of the observation $y_{d}$ under the (conditional) marginal distribution (see Kock and Klein, 2023; Arbenz, 2013). We fit the Copula model by the transforming the in-sample data to the uniform space $\bm{\mathrm{u}}$ by the probability integral transformation (PIT) and subsequently transforming to the $\mathcal{N}(0,1)$ space $\bm{\mathrm{n}}$, on which we can fit the dependence structure. We update the scale matrix of the Gaussian copula by taking
$\widehat{\bm{\mathrm{\Sigma}}}^{[t+1]}=\frac{t-1}{t}\widehat{\bm{\mathrm{\Sigma}}}^{[t]}+\frac{1}{t}\left(\bm{\mathrm{n}}^{[t+1]}(\bm{\mathrm{n}}^{[t+1]})^{\top}\right)$ (21)
where $\bm{\mathrm{n}}$ are the PIT-transformed in-sample values and the superscript $[t]$ denotes the observations available in the online learning (see e.g. Dasgupta and Hsu, 2007). Samples are drawn from the Gaussian copula in the usual manner. We use the same principle to track the residual covariance structure for the LARX models under the normality assumption. We employ a second model, where we sparsify the estimated dependence matrix of the Gaussian copula by the graphical LASSO (Friedman et al., 2008).
3.3 Forecast Evaluation and Scoring Rules
Forecast evaluation should follow the well-known principle of sharpness subject to calibration Gneiting et al. (2007). We check the calibration of the forecasts by calculating joint prediction bands (JPB) from the simulations. JPB aim to cover the true price vector with certain probability $1-\alpha$ and can be thought of a multivariate generalization of marginal prediction intervals (Staszewska-Bystrova, 2011; Lütkepohl et al., 2015). JPBs have been used in energy market forecasting by Serafin et al. (2022); Chen et al. (2025). As it is standard in forecasting probabilistic forecasting, we evaluate both marginal and multivariate quality of the forecasts. Therefore, we employ the root mean square error RMSE (RMSE), mean absolute error (MAE) and the continuous ranked probability score (CRPS), which focus on mean/point prediction and the marginal distribution. We employ four well-established multivariate probabilistic scoring scores: The Energy Score (ES), the Dawid-Sebastiani Score (DSS), the Variogram Score (VS) and the Log-Score (LS). Let us briefly review some recent results with respect to the four multivariate scores:
  • 1.
    The ES, DSS and LS are all able to reliably detect mis-specifications in the mean structure of the multivariate distribution (Marcotte et al., 2023), while the VS (by design) is not sensitive. However, Pinson and Tastu (2013); Alexander et al. (2024) discuss the discrimination ability of the ES. While widely used for multivariate forecast evaluation, the ES has been shown to have low discrimination ability with respect to misspecified covariance structures, especially in double-digit and larger dimensions and Alexander et al. (2024) recommend to use the VS in addition to the ES.
  • 2.
    In a similar vein, Marcotte et al. (2023) show that the ES has low reliability, i.e. statistical power to discriminate between a correctly and incorrectly specified dependence model for multivariate forecasts compared to the DSS and the LS, with the VS being somewhat in the middle. Their analysis also reflects the role of the number of test samples (which, with more than 1800 out-of-sample days, is not a concern in our study) and the number of sample paths $M$. Interestingly, they find that the ES and the VS are complimentary, i.e. the VS is more reliable in cases where the ES is not and vice versa, and they find the reliability of the VS to increase non-monotonically with respect to the dimension $D$ and the number of samples $M$.
  • 3.
    In contrast, Ziel and Berk (2019) find that the ES, in combination with the Diebold-Marino test, is able to detect the correct model specification in a simulation study. Nevertheless, they still recommend the use of multiple scores.
Additionally, we test for statistically significant score differences using the well-established Diebold-Mariano test. The following paragraphs introduce JPBs and the scoring rules and are largely based on Gneiting et al. (2007); Gneiting and Raftery (2007); Nowotarski and Weron (2018); Marcotte et al. (2023); Ziel and Berk (2019) as well as the references mentioned for the individual scores. Denote the true price vector as $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{H})$ of shape $T\times H$ and the ensemble forecast as $\bm{\mathrm{F}}$ of shape $T\times H\times M$ of $M=2500$ samples.
A joint prediction band for the $1-\alpha$ coverage is defined by the lower bound $\bm{\mathrm{l}}_{t}=(l_{t,0},...,l_{t,23})$ and upper bound $\bm{\mathrm{u}}_{t}=(u_{t,0},...,u_{t,23})$, such that
$\mathbb{P}_{t}\left(\bm{\mathrm{l}}_{t}\leq\bm{\mathrm{y}}_{t}\leq\bm{\mathrm{u}}_{t}\;\forall\;h\in\{0,...,23\}\right)=1-\alpha,$
that is, the true price trajectory $\bm{\mathrm{y}}_{t}$ is fully covered by the JPB with probability $1-\alpha$. This is in difference to marginal prediction intervals based on predicted quantiles, which consider element-wise coverage. The difference between marginal, quantile-based prediction bands and joint prediction bands is shown in Figure 8. There are multiple algorithms to construct such bands and following Serafin et al. (2022); Chen et al. (2025), we use the neighouring paths method described by Staszewska-Bystrova (2011); Lütkepohl et al. (2015). The methods is based on iteratively removing paths from the ensemble forecast $\bm{\mathrm{F}}$, such that the envelope of the remaining paths covers the true price trajectory $\bm{\mathrm{y}}_{t}$ with probability $1-\alpha$. It should be noted that JPB are in general not unique, i.e. there are multiple sets of lower and upper bounds that cover the true trajectory with the desired probability. We compare the mis-coverage of the prediction interval, which is defined as
$\operatorname{MC}_{1-\alpha}=\frac{1}{T}\sum^{T}_{t=0}\bm{1}\left(\bm{\mathrm{l}}_{t}\leq\bm{\mathrm{y}}_{t}\leq\bm{\mathrm{u}}_{t}\;\forall\;h\in\{0,...,23\}\right)-(1-\alpha),$ (22)
and gives a measure of the multivariate calibration of the joint prediction. Additionally, we evaluate the mean width of the prediction bands, which is defined as
$\operatorname{JPBW}_{1-\alpha}=\frac{1}{TH}\sum^{T}_{t=0}\sum^{H}_{h=0}\left(\bm{\mathrm{u}}_{t,h}-\bm{\mathrm{l}}_{t,h}\right),$ (23)
which gives a measure of the efficiency of the prediction bands and can be interpreted as the area of the prediction band divided by the dimensions.
Figure 8: Illustrative 50%-prediction bands based on marginal quantiles and the iterative method for JPBs presented in Lütkepohl et al. (2015) and Staszewska-Bystrova (2011). In the upper panel, the simulations disregard any dependece between the 24 delivery hours, otherwise the marginal distributions are identical. Hence, the quantile prediction bands are similar (up to simulation noise), while for the joint prediction bands, modelling the dependence decreases the width of the prediction bands. Own illustration.
The RMSE is defined as
$\operatorname{RMSE}=\sqrt{\frac{1}{TH}\sum_{t=0}^{T}\sum_{h=0}^{H}\left(\bm{\mathrm{y}}_{t}-\hat{\bm{\mathrm{\mu_{t}}}}\right)^{2}}$ (24)
where $\hat{\bm{\mathrm{\mu}}}_{t}=\frac{1}{M}\sum_{m=0}^{M}\bm{\mathrm{F}}_{t,h,m}$ is the mean prediction vector. The MAE is defined as
$\operatorname{MAE}=\frac{1}{TH}\sum_{t=0}^{T}\sum_{h=0}^{H}\left|\bm{\mathrm{y}}_{t}-\operatorname{median}\left(\bm{\mathrm{F_{t}}}\right)\right|$ (25)
where $\operatorname{median}\left(\bm{\mathrm{F_{t}}}\right)$ denotes the median trajectory for each day $t$. The CRPS is estimated from the forecast ensemble by using the probability-weighted moment estimator of Zamo and Naveau (2018):
$\operatorname{CRPS}_{t}=\frac{1}{M}\sum_{m=0}^{M}|\bm{\mathrm{F}}_{t,h,m}-y_{t.h}|+\frac{1}{M}\sum_{m=0}^{M}\bm{\mathrm{F}}_{t,h,m}+\frac{1}{M(M-1)}\sum_{m=0}^{M}m\bm{\mathrm{F}}_{t,h,m}.$ (26)
The CRPS is strictly proper scoring rule for the marginal distribution. Note that many works on energy price forecasting report the average pinball loss (APS) as CRPS, which needs to be rescaled $\operatorname{CRPS}=2\cdot\operatorname{APS}$ to be comparable. The energy score (ES, Gneiting and Raftery, 2007) is defined as
$\operatorname{ES}_{t}=\frac{1}{M}\sum_{m=0}^{M}\lVert\bm{\mathrm{y}}_{t}-\bm{\mathrm{F}}_{t,m}\rVert_{2}^{2}-\frac{1}{M^{2}}\sum_{i=0}^{M}\sum_{j=i+1}^{M}\lVert\bm{\mathrm{F}}_{t,i}-\bm{\mathrm{F}}_{t,j}\rVert_{2}^{2}.$ (27)
The energy score is a strictly proper scoring rule. We aggregate the ES by taking the average: $\operatorname{ES}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{ES}_{t}$. The Log-Score (LS) is defined as
$\operatorname{LS}_{t}=-\log\left(\mathcal{L}(\bm{\mathrm{y}}_{t}\mid\hat{\bm{\mathrm{\theta}}}^{\mathcal{D}}_{t})\right),$ (28)
where $\mathcal{L}$ is the underlying likelihood or probability density function of the distribution $\mathcal{D}$ and $\hat{\bm{\mathrm{\theta}}}^{\mathcal{D}}_{t}$ is the estimated parameter vector. Again, we aggregate the LS by simple averaging over all points in the test set $\operatorname{LS}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{LS}_{t}$. It is a strictly proper scoring rule. The Dawid-Sebastiani-Score (DSS, 1999) is defined as
$\operatorname{DSS}_{t}=\log\left(\det(\widehat{\bm{\mathrm{\Sigma}}}_{F})\right)+(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}})\widehat{\bm{\mathrm{\Sigma}}}_{F}^{-1}(\bm{\mathrm{y_{t}}}-\bm{\mathrm{\hat{\mu}_{t}}}),$ (29)
where $\hat{\Sigma}_{F}$ denotes the empirical covariance of the forecast ensemble $\bm{\mathrm{F}}$ and $\hat{\bm{\mathrm{\mu}}}$ denotes the mean ensemble as above. We aggregate the $\operatorname{DSS}=\frac{1}{T}\sum_{t=0}^{T}\operatorname{DSS}_{t}$ by simple averaging. The DSS is a proper scoring rule for the first and second moment and strictly proper for the Gaussian predictive distribution, since it is a linear transformation of Gaussian log-likelihood. The Variogram Score (VS, Scheuerer and Hamill, 2015) is defined as
$\operatorname{VS}^{p}_{t}=\sum_{i=0}^{H}\sum_{j=0}^{H}\left(\frac{1}{M}\sum_{m=0}^{M}|\bm{\mathrm{F}}_{t,i,m}-\bm{\mathrm{F}}_{t,j,m}|^{p}-|y_{t,i}-y_{t.j}|^{p}\right)^{2}$ (30)
and is a proper scoring rule. We aggregate the VS by taking the average and normalize the score by dividing by $H^{2}$, i.e. $\operatorname{VS}=\frac{1}{TH^{2}}\sum_{t=0}^{T}{\operatorname{VS}_{t}}$ to make the scales of the score comparable. The scoring rules used are implemented in the Python package scoringrules (Zanetta and Allen, 2024).
Conclusions on the performance of forecasting models cannot be derived by looking at aggregate scores alone, but need to be drawn by evaluating whether the differential between the loss series of two models is statistically significantly from zero (Diebold and Mariano, 2002; Diebold, 2015). For the DM-test, we evaluate the differential of two score series $\Delta\bm{\mathrm{s}}^{\mathcal{A},\mathcal{B}}=\bm{\mathrm{s}}^{\mathcal{A}}-\bm{\mathrm{s}}^{\mathcal{B}}$, where $\bm{\mathrm{s}}^{\mathcal{A}}=(s_{0}^{\mathcal{A}},...,s_{T}^{\mathcal{A}})$ are the scores for each scoring rule at $t$ for model $\mathcal{A}$ respectively $\mathcal{B}$. We provide two one-sided and hence complimentary tests. To ensure validity of the DM-test, we check stationarity of the differential series $\Delta\bm{\mathrm{s}}^{\mathcal{A},\mathcal{B}}$ by the augmented Dickey-Fuller test (ADF, Dickey and Fuller, 1979; Cheung and Lai, 1995).
4 Results
We present the results of our forecasting study in three parts. First, we illustrate the differences between the models by showing example simulations and analyzing the time-varying dependence structure. Second, we present the forecasting accuracy of all models using the scoring rules presented above. Third, we analyze the role of overfitting and regularization in our proposed path-based estimation approach. Lastly, we discuss the computational costs of the different models compare the online and batch estimation. Figure 9 shows illustrative simulations drawn from two models. Figure 10 shows exemplary predicted covariance matrices.
Figure 9: Illustrative simulations drawn from two models.
Figure 10: Illustrative predicted covariance matrices for one week in the test sample.
4.1 Forecasting the Time-varying Dependence Structure
Seven exemplary predicted covariance matrices are shown in Figure 10 and some time-varying behavior of the covariance matrix over the week, especially in the morning hours is visible. This section aims to further analyze the time-varying dependence structure. Figure 11 plots the evolution of the predicted volatility (standard deviation) and correlation matrix for the oMvDistReg($t$, MCD ,OLS) model over time. The lower three panels show the 1st, 2nd and 3rd off-diagonal of the correlation matrix over time. We see that the correlation is the lowest around the hours 5-7, which corresponds to start of the morning ramp. The correlation between to hours $h$ and $h+i$ decays stronger for days with lower levels of volatility and vice versa. There are some weekly patterns visible: The dependence between the hours 5-7 and 16-19 tapers off more strongly on working days. This is likely driven by stronger shapes in prices and residual demand. Overall, we see that the dependence structure is not constant over time, which underlines the importance of including a time-varying dependence structure in multivariate electricity price forecasting.
Figure 11: Evolution of the predicted volatility (standard deviation) and correlation matrix for the oMvDistReg($t$,MCD,OLS) model over time. The lower three panels show the 1st, 2nd and 3rd off-diagonal of the correlation matrix over time. We see that the correlation is the lowest around the hours 5-6, which corresponds to start of the morning ramp. The correlation between to hours $h$ and $h+i$ decays stronger for days with lower levels of volatilty and vice versa. There are some weekly patterns visible.
4.2 Forecasting Accuracy
Panel A: Full sample (2018-12-27 to 2023-12-31) RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS $\text{MC}_{0.95}$ $\text{JPBW}_{0.95}$ LARX+N(0, $\sigma$) 31.177 18.214 13.677 4.709 887.202 81.201 176.367 110.038 -0.134 131.129 LARX+N(0, $\Sigma$) 31.182 18.219 13.678 3.396 654.953 78.603 124.345 84.057 -0.161 118.914 LARX+Adaptive CP 31.268 18.374 14.402 -0.655 53.751 LARX+GARCH(1,1) 31.181 18.218 12.992 4.536 850.636 78.312 159.545 101.790 -0.040 140.459 oDistReg 36.040 19.104 13.980 4.722 1046.549 84.445 160.422 99.266 0.015 332.441 oDistReg+GC 33.901 18.511 13.501 3.591 793.156 78.372 124.869 77.934 -0.001 294.100 oDistReg+spGC 33.964 18.522 13.508 3.597 793.049 78.417 124.646 77.964 0.004 294.163 oMvDistReg(t, CD, OLS, ind) 38.898 20.207 15.410 3.787 771.899 89.011 174.092 95.446 -0.046 191.574 oMvDistReg(t, CD, OLS) 46.742 23.493 17.463 3.747 797.964 97.948 126.357 76.013 -0.017 273.946 oMvDistReg(t, CD, LASSO) 46.793 23.474 17.612 3.938 872.181 98.819 127.551 75.598 0.002 306.158 oMvDistReg(t, MCD, OLS, ind) 38.937 20.196 15.379 3.861 792.223 88.771 175.635 95.274 -0.034 196.792 oMvDistReg(t, MCD, OLS) 46.450 23.376 17.255 3.596 747.954 96.739 121.365 74.907 -0.037 234.422 oMvDistReg(t, MCD, LASSO) 46.581 23.345 17.261 3.699 787.099 96.803 124.217 74.789 0.017 298.814 oMvDistReg(t, LRA, OLS, ind) 39.197 20.292 15.710 5.189 >10,000>10,000 93.714 175.884 95.602 -0.049 201.897 oMvDistReg(t, LRA, OLS) 41.642 21.292 16.520 3.940 1161.734 94.398 174.901 93.993 -0.072 176.183 oMvDistReg(t, LRA, LASSO) 41.725 21.328 16.419 3.911 963.429 93.844 169.019 93.780 -0.047 188.270
Panel B: First sub-sample (2018-12-27 to 2020-12-31, $n=736$) RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS $\text{MC}_{0.95}$ $\text{JPBW}_{0.95}$ LARX+N(0, $\sigma$) 7.568 4.790 3.588 1.163 59.095 21.940 120.417 82.117 -0.063 39.127 LARX+N(0, $\Sigma$) 7.568 4.791 3.588 0.912 52.607 21.368 85.142 64.467 -0.080 36.139 LARX+Adaptive CP 7.751 4.959 3.831 -0.525 18.469 LARX+GARCH(1,1) 7.567 4.791 3.601 1.258 65.391 22.099 118.522 81.369 -0.049 42.002 oDistReg 7.849 4.552 3.399 0.994 56.194 21.024 112.649 74.195 -0.006 50.476 oDistReg+GC 7.828 4.551 3.394 0.877 53.687 20.579 76.162 56.599 -0.025 43.055 oDistReg+spGC 7.819 4.548 3.390 0.877 53.706 20.557 75.768 56.561 -0.021 43.102 oMvDistReg(t, CD, OLS, ind) 8.034 4.664 3.557 0.952 56.556 21.777 133.695 71.549 -0.061 44.810 oMvDistReg(t, CD, OLS) 8.050 4.713 3.536 0.916 55.459 21.373 84.530 56.039 -0.045 49.244 oMvDistReg(t, CD, LASSO) 8.050 4.671 3.525 0.927 56.211 21.346 84.322 55.780 -0.012 56.089 oMvDistReg(t, MCD, OLS, ind) 8.056 4.659 3.509 0.961 56.386 21.557 122.110 70.944 -0.017 49.627 oMvDistReg(t, MCD, OLS) 8.059 4.699 3.505 0.900 54.724 21.277 80.581 55.291 -0.036 45.742 oMvDistReg(t, MCD, LASSO) 8.080 4.674 3.500 0.920 55.525 21.281 81.532 55.158 0.007 58.907 oMvDistReg(t, LRA, OLS, ind) 8.047 4.662 3.564 0.941 56.963 21.830 126.458 71.156 -0.041 44.971 oMvDistReg(t, LRA, OLS) 8.057 4.681 3.566 0.952 56.712 21.840 120.469 69.491 -0.034 45.885 oMvDistReg(t, LRA, LASSO) 8.079 4.671 3.552 0.959 57.035 21.798 120.635 69.375 -0.021 48.611
Panel C: Second sub-sample (2021-01-01 to 2023-12-31, $n=1095$) RMSE MAE CRPS $\text{VS}_{p=0.5}$ $\text{VS}_{p=1}$ ES DSS LS $\text{MC}_{0.95}$ $\text{JPBW}_{0.95}$ LARX+N(0, $\sigma$) 39.835 27.237 20.459 7.093 1443.811 121.033 213.973 128.804 -0.182 192.967 LARX+N(0, $\Sigma$) 39.841 27.245 20.460 5.065 1059.817 117.073 150.695 97.225 -0.215 174.550 LARX+Adaptive CP 39.931 27.391 21.507 -0.742 77.465 LARX+GARCH(1,1) 39.840 27.243 19.304 6.739 1378.435 116.095 187.118 115.516 -0.034 206.636 oDistReg 46.157 28.885 21.092 7.228 1712.213 127.073 192.533 116.117 0.029 521.962 oDistReg+GC 43.365 27.894 20.295 5.415 1290.187 117.218 157.607 92.275 0.014 462.839 oDistReg+spGC 43.450 27.915 20.309 5.425 1289.995 117.308 157.499 92.349 0.021 462.912 oMvDistReg(t, CD, OLS, ind) 49.867 30.654 23.377 5.693 1252.714 134.201 201.244 111.508 -0.035 290.222 oMvDistReg(t, CD, OLS) 60.081 36.115 26.825 5.650 1297.036 149.418 154.471 89.439 0.002 424.979 oMvDistReg(t, CD, LASSO) 60.148 36.111 27.081 5.962 1420.631 150.892 156.608 88.918 0.012 474.241 oMvDistReg(t, MCD, OLS, ind) 49.915 30.639 23.357 5.810 1286.812 133.948 211.612 111.628 -0.045 295.708 oMvDistReg(t, MCD, OLS) 59.700 35.929 26.497 5.407 1213.906 147.461 148.778 88.092 -0.038 361.243 oMvDistReg(t, MCD, LASSO) 59.869 35.895 26.510 5.567 1278.824 147.565 152.907 87.985 0.024 460.066 oMvDistReg(t, LRA, OLS, ind) 50.255 30.798 23.874 8.044 >10,000>10,000 142.030 209.106 112.033 -0.054 307.374 oMvDistReg(t, LRA, OLS) 53.441 32.456 25.226 5.949 1904.470 143.167 211.488 110.463 -0.098 263.763 oMvDistReg(t, LRA, LASSO) 53.547 32.524 25.068 5.894 1572.657 142.270 201.539 110.184 -0.064 282.142
Table 4: Scoring Rules. The best score in each column is marked bold. Note that the $\operatorname{LARX}+\mathcal{N}(0,\sigma)$, the oDistReg and the oMvDistreg(…, ind) models do not model the dependence structure.
Table 4 gives the results for the scoring rules for each model. Figure 12 gives daily skills scores of the multivariate distributional regression models over the Copula-based benchmarks. Figure 13 provides one-sided Diebold-Mariano-tests for all pairwise model comparisons. As a mental guidance for the increasing complexity of the online regression models, remember Section 3: The first three models have an adaptive, but unconditional estimation for the scale parameter/matrix under a Gaussian assumption. The Copula-based models employ online, conditional estimation for all marginal distributional parameters and an adaptive, but unconditional estimation for the dependence structure, while the multivariate distributional regression models yield an online estimated conditional multivariate distribution.
Let us note a few main results here, before we discuss the results in more detail.
  • 1.
    Calibration: In terms of the calibration of the joint prediction bands, the multivariate distributional regression models and the Copula-based models yield the best performance. The univariate models yield poor performance with up to 10-16% under-coverage the 95%-JPB for the Gaussian-based models and, given we can only use marginal, quantile-based prediction bands for conformal prediction, 65% under-coverage for the LARX-ACP model. For the models yielding calibrated JPBs, the multivariate distributional regression models yield the narrowest JPBs.
  • 2.
    Marginal Scores: Here, the LARX models yield the best performance in terms of RMSE, MAE and CRPS. These models yield sharp predictions, but at the cost of poor coverage of the joint prediction bands. This is also visible in the very small JPBW.
  • 3.
    Multivariate Scores: The Copula-based models yield the best performance in terms of the ES. The MCD-based multivariate distributional regression model yields the best overall performance in terms of the LS and DSS as well as a second-best performance in the VS.
The $p$-values of the DM-test in Figure 13 largely confirm the statistical significance of the aforementioned results. We note the strong performance of the Copula-based models for the ES and the statistically significant superior performance of the multivariate distributional regression for the DSS, LS, and the VS. We note the disconnect between the results of the ES on one hand and the VS, DSS and LS on the other hand, which is in line with the findings of Pinson and Tastu (2013); Marcotte et al. (2023) and Alexander et al. (2024).
Looking at Figure 12, we see that skill scores are also well correlated with the current market regime. Skill scores are defined as
$\operatorname{SS}_{\text{model}}=1-\frac{\operatorname{LS}_{\text{model}}}{\operatorname{LS}_{\text{baseline}}},$
where we use the oDist+GC model as baseline. During the period of high and volatile prices in 2022, the skill scores of the multivariate distributional regression models over the Copula-based benchmarks are at their highest, while during the low-price period in 2020, the skill scores are at their lowest. Additionally, we see that the models assuming independence have have an increasingly better performance in recent years. This can be interpreted as a sign that the dependence structure has both weakened and changed with the changing market conditions.
Figure 12: Rolling 182-day average of skill scores. The skill score is defined as $\operatorname{SS}=1-\operatorname{LS}_{\text{model}}/\operatorname{LS}_{\text{baseline}}$, where the baseline is the oDist+GC model. A positive skill score indicates an improvement over the baseline. The secondary $y$-axis shows the daily average day-ahead price.
A common theme for distributional regression models is the trade-off between accuracy in the point predictions and the distributional error metrics. Marcjasz et al. (2023) and Hirsch et al. (2024) note this observation in the univariate case, trading RMSE for CRPS. This behavior is rooted in the fact that observations with predicted high variance are weighted down in the estimation of the mean parameters, which leads to slightly worse point predictions, but better distributional forecasts. At the same time, these observations with high variance are especially costly in terms of the RMSE. We see that the behaviour extends in the multivariate case as well, trading off marginal accuracy somewhat better modeling of the dependence structure in otherwise equal models. So far, this issue has not been discussed for multivariate distributional regressions and yet deserves further attention in future research (Gioia et al., 2022; Muschinski et al., 2022).
For the multivariate online distributional regression models, we see that the models using the Cholesky-based parametrizations provide better performance. This is likely due to the fact that these parametrizations are closer to the natural, time-based structure of the (conditional) covariance than the LRA and therefore the path-based regularization allows to to select a parsimounous model, which still reflects the shape of the dependence structure. We further discuss the regularization in the following Subsection 4.3. On the other hand, the estimation using LASSO yields slight improvements in the scoring rules. The rather marginal gain might be explained by the fact that all fundamental variables are known to be relevant for electricity price formation, and hence the regularization does not remove many variables. However, the regularization might help to stabilize the estimates in the online learning setting.
Figure 13: Diebold-Mariano Test Matrix. A $p$-value $p<0.05$ implies that the forecasts given by a model on the column are significantly better than forecasts by a model on the row.
Additionally, we analyze the performance of the models for extreme price spikes and for days with large spreads during the day. These days are interesting from a risk-management respectively from a battery optimization perspective. We define a price spike as a day with a minimum or maximum price exceeding the 5% resp. 95%-quantile of all minimum resp. maximum prices. A large spread event is defined as the min-to-max spread exceeding the 90%-quantile of all min-to-max spreads. Formally:
$\operatorname{Spike}$ $=\{t\mid\min(\bm{\mathrm{y}}_{t})<\operatorname{Q}_{0.05}(\min(\bm{\mathrm{y}}_{t}))\lor\max(\bm{\mathrm{y}}_{t})>\operatorname{Q}_{0.05}(\max(\bm{\mathrm{y}}_{t}))\},\quad\text{and}$
$\operatorname{Spread}$ $=\{t\mid\max(\bm{\mathrm{y}}_{t})-\min(\bm{\mathrm{y}}_{t})>\operatorname{Q}_{0.90}(\max(\bm{\mathrm{y}}_{t})-\min(\bm{\mathrm{y}}_{t}))\}.$
The results for the CRPS, LS and the miscoverage are given in Table 5. For the CRPS, we see that the ordering of the models is similar to the overall results, and remains largely unchanged between spike and no-spike events. For the miscoverage, we see that the multivariate distributional regression models yield the best results for both regimes for large spread events, and the best results for the price spike events, while the Copula-based models yield the best results for the no-spike events. Overall, these results highlight the robustness of the distributional regression models also for extreme events.
CRPS LS $\text{MC}_{0.95}$
Spread Spike Spread Spike Spread Spike
Low High Low High Low High Low High Low High Low High
LARX+N(0, $\sigma$) 3.53 (05) 37.85 (02) 10.93 (03) 29.02 (03) 80.43 (15) 169.63 (15) 100.06 (15) 157.46 (15) -0.06 (13) -0.37 (14) -0.10 (14) -0.45 (13)
LARX+N(0, $\Sigma$) 3.53 (06) 37.85 (03) 10.92 (02) 29.02 (02) 63.70 (07) 123.39 (07) 77.75 (07) 119.74 (07) -0.06 (15) -0.43 (15) -0.12 (15) -0.52 (15)
LARX+Adaptive CP 3.67 (16) 38.27 (04) 11.25 (06) 30.16 (04) (17) (17) (17) (17) -0.56 (16) -0.79 (16) -0.56 (16) -0.86 (16)
LARX+GARCH(1,1) 3.55 (07) 35.11 (01) 10.41 (01) 27.27 (01) 79.99 (14) 135.37 (08) 95.51 (14) 144.52 (14) -0.03 (03) -0.10 (07) -0.07 (13) -0.28 (09)
oDistReg 3.40 (03) 49.54 (07) 11.55 (07) 39.18 (07) 74.13 (13) 143.14 (14) 91.18 (13) 132.34 (13) -0.03 (06) 0.03 (05) 0.00 (01) -0.10 (01)
oDistReg+GC 3.40 (02) 46.18 (05) 11.02 (04) 35.96 (05) 56.46 (06) 114.50 (05) 71.92 (05) 102.66 (06) -0.04 (09) 0.02 (03) -0.02 (06) -0.15 (04)
oDistReg+spGC 3.39 (01) 46.22 (06) 11.04 (05) 36.01 (06) 56.44 (05) 114.55 (06) 71.96 (06) 102.65 (05) -0.03 (06) 0.03 (05) -0.02 (05) -0.13 (03)
oMvDistReg(t, CD, OLS, ind) 3.55 (08) 60.09 (09) 13.23 (09) 48.49 (09) 71.37 (12) 139.14 (11) 88.52 (12) 126.91 (11) -0.05 (11) -0.15 (08) -0.04 (09) -0.32 (11)
oMvDistReg(t, CD, OLS) 3.62 (13) 67.75 (15) 14.95 (14) 54.82 (15) 56.16 (04) 111.61 (04) 70.22 (04) 100.64 (04) -0.03 (06) -0.03 (04) -0.01 (03) -0.18 (06)
oMvDistReg(t, CD, LASSO) 3.62 (14) 67.69 (14) 15.09 (16) 54.60 (12) 56.03 (03) 111.15 (03) 69.96 (03) 100.29 (03) -0.03 (03) -0.00 (01) -0.00 (02) -0.16 (05)
oMvDistReg(t, MCD, OLS, ind) 3.52 (04) 60.06 (08) 13.14 (08) 48.35 (08) 70.90 (10) 139.46 (12) 88.10 (10) 126.46 (10) -0.03 (03) -0.18 (11) -0.02 (06) -0.27 (08)
oMvDistReg(t, MCD, OLS) 3.61 (12) 68.64 (16) 15.00 (15) 56.01 (16) 55.39 (02) 110.49 (02) 69.26 (02) 99.20 (02) -0.06 (13) -0.15 (09) -0.05 (11) -0.26 (07)
oMvDistReg(t, MCD, LASSO) 3.62 (15) 67.62 (13) 14.91 (13) 54.65 (13) 55.32 (01) 110.23 (01) 69.20 (01) 98.83 (01) -0.02 (01) -0.01 (02) 0.01 (04) -0.12 (02)
oMvDistReg(t, LRA, OLS, ind) 3.57 (09) 60.27 (10) 13.60 (10) 48.83 (10) 71.12 (11) 140.08 (13) 88.50 (11) 126.99 (12) -0.05 (11) -0.17 (10) -0.05 (11) -0.31 (10)
oMvDistReg(t, LRA, OLS) 3.58 (10) 66.68 (12) 14.42 (12) 54.65 (14) 69.45 (09) 138.95 (10) 86.85 (09) 125.42 (09) -0.04 (10) -0.33 (13) -0.05 (11) -0.45 (13)
oMvDistReg(t, LRA, LASSO) 3.58 (11) 66.32 (11) 14.28 (11) 54.29 (11) 69.41 (08) 138.54 (09) 86.73 (08) 125.26 (08) -0.03 (06) -0.21 (12) -0.03 (08) -0.35 (12)
Table 5: Price spike and spread analysis for the CRPS and the miscoverage $\text{MC}_{0.95}$. A spike event is defined if the min/max price of a day exceeds the 5% resp. 95%-quantile of all min/max prices. A large spread event is defined as the min-to-max spread exceeding 90%-quantile of all min-to-max spreads. Numbers in the brackets give the model ranking.
Before concluding the results, we want to emphasize some limitations this study.
  • 1.
    We acknowledge that the proposed models yield a trade-off between accuracy in the marginal distributions and correct modeling of the dependence structure. This trade-off is reflected also in the disconnect between marginally dominated scoring rules (CRPS and ES) and the LS, DSS and VS and therefore, we yield somewhat ambiguous results.
  • 2.
    While the results of the spike analysis are encouraging, the parametric assumption on the distributional family might be too restrictive in some cases and could be complemented with non-parametric approaches, or with multivariate extreme value models to capture the tail behavior. This could also help to improve the forecasting performance in the marginal distributions and hence CRPS and ES.
Overall, our results highlight that neglecting the dependence structure by relying solely on marginal, univariate models yields subpar probabilistic forecasting performance. We note that for the truly multivariate approaches, using both, Copula-based combinations of univariate models and the fully multivariate distributional regression yield statistically significant performance improvements. In this regard, this paper is the first to carry out a comprehensive, multivariate probabilistic forecasting study on the day-ahead market, including also the challenging years of the COVID-19 pandemic and the energy crisis following the Russian invasion of Ukraine in the test set.
4.3 Overfitting and Regularization
Figure 14: Analysis of Overfitting. We estimate the oMvDistReg($t$,MCD,OLS) in a 8-fold cross-validation setting with 100 out-of-sample days without early stopping and monitor the in-sample and out-of-sample LS. Confidence bands are 95%-confidence intervals based on the standard deviation of the LS. The number of model coefficients increases with the amount of modeled elements. The information-criterion based early stopping suggest to include one (BIC, HQC) respectively four (AIC) off-diagonals.
Modeling all elements of the scale matrix can lead to overfitting, especially in the high dimensional setting of energy markets. To counter this, we have proposed a path-based regularization approach in Section 2.6. Figure 14 shows the results of a small experiment on the initial training set. We estimate the oMvDistReg($t$,MCD,OLS) in a 8-fold cross-validation setting with 100 out-of-sample days without early stopping. We monitor the in-sample and out-of-sample LS for the number of off-diagonals of $\bm{\mathrm{L}}$ modeled in the path-wise estimation of $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{D}}\bm{\mathrm{L}}$. We see that the out-of-sample LS barely increases after modelling the first off-diagonal and starts to degrade after the 4th to 5th off-diagonal is included in the model for $\bm{\mathrm{\Omega}}$. At the same time, the number of model coefficients increases with the amount of modeled elements. The information-criterion based early stopping suggest to include one (BIC and HQC) respectively four (AIC) off-diagonals. Figure 15 shows the out-of-sample LS for the oMvDistReg($t$,MCD,OLS) model for different levels of path-based regularization, which are well-aligned with the in-sample analysis, as the LS plateaus after including the first off-diagonal. In combination with the results from the previous section, we see that our approach allows for parsimonious, yet interpretable time-varying modeling of the dependence structure by exploiting the natural ordering of the hours in the day-ahead market.
Figure 15: Results for fitting the path-based regularization for the oMvDistReg($t$,MCD,OLS) model.
4.4 Computation Time
Model Initial Fit Avg. Update Total Time Est. Speedup
LARX+N(0, $\sigma$) 1.23 0.06 103.80 $\times$ 21
LARX+N(0, $\Sigma$) 1.23 0.06 103.80 $\times$ 21
LARX+Adaptive CP 1.23 0.06 103.80 $\times$ 21
LARX+GARCH(1,1) 1.44 0.27 501.31 $\times$ 5
oDistReg 55.31 0.30 598.78 $\times$ 169
oDistReg+GC 55.32 0.30 599.61 $\times$ 168
oDistReg+spGC 55.46 0.45 870.45 $\times$ 116
oMvDistReg(t, CD, OLS, ind) 38.26 0.07 173.88 $\times$ 402
oMvDistReg(t, CD, OLS) 123.11 0.18 460.38 $\times$ 489
oMvDistReg(t, CD, LASSO) 460.46 2.34 4736.02 $\times$ 178
oMvDistReg(t, MCD, OLS, ind) 41.72 0.06 147.07 $\times$ 519
oMvDistReg(t, MCD, OLS) 132.20 0.15 403.89 $\times$ 599
oMvDistReg(t, MCD, LASSO) 1484.75 1.79 4754.88 $\times$ 571
oMvDistReg(t, LRA, OLS, ind) 21.82 0.03 83.58 $\times$ 477
oMvDistReg(t, LRA, OLS) 289.85 0.77 1702.47 $\times$ 311
oMvDistReg(t, LRA, LASSO) 5654.00 3.73 12473.45 $\times$ 829
Table 6: Computation times. All timings are in seconds. The out-of-sample data for the forecasting study consists of 1831 days. We update the 24-dimensional distributional regression model each model on each day. All experiments are run on a standard laptop (Intel Core i7 (16 Threads, 4.9 GHz), 32GB RAM). Estimated speed-ups are calculated by taking $\text{Speedup}=(\text{Initial Fit}\times T)/\text{Total Time}$. The figure on the right shows the share of the time spend on the initial fit and on the out-of-sample updating.
Table 6 gives computation times for all experiments. The initial fit for the multivariate distributional regression model takes a few minutes, the update algorithm can be executed in seconds. Using online estimation methods, the experiments can be run in about a few hours on a standard laptop. An estimate for the benefit of online vs. repeated batch fitting can be achieved by multiplying the initial fit duration with 1831 days of out of sample and comparing this to the total time of the online study:
$\text{Speedup}=\frac{\text{Initial Fit}\times T}{\text{Total Time}}.$
By this (albeit simple) measure, the online learning improves computation by a factor of 80 to 600. These estimates are in line with benefits reported in Hirsch et al. (2024) for the univariate online distributional regression case in direct comparison between online estimation and repeated batch estimation.
Figure 16: Efficient frontier of computation time against average log-score for the online multivariate distributional regression model for different update frequencies in the batch and online case. Note that the $x$-axis is in log-scale.
To further analyze the trade-off between computation time and accuracy, we take the OLS-estimated distributional regression model and estimate the model every 1, 7, 14, 30, 60, 180, 365 days online (using a mini-batch online update) and, vice versa, re-estimate the full model in a repeated batch fitting every 7, 14, …, 365 days, using the first subsample of the test data. [^4] Note that already estimating the batch model every 7 days takes more than 10 hours and hence we did not run the experiment for daily estimation in the batch case. We compare the results for online learning, rolling and expanding window batch estimation in Figure 16. For online and batch estimation, we see that increasing update frequency increases forecasting accuracy, but also increases computation times. Crucially, the “efficient frontiers“ of both approaches never intersect. We see that for low computation time budgets, online learning approaches give strictly better results than repeated batch fitting approaches and only for large computation time budgets, expanding window batch estimation takes the lead. A second important observation is that the online learning algorithm is almost constant in the time per update, since we always process a fixed amount of data points. On the other hand the number of data points in the expanding window batch estimation increases, which increases the computation time per fit. Interestingly, the rolling window estimation scheme, which is often favored in the EPF literature, delivers worse predictive accuracy than an expanding window scheme. These results further confirm the estimated speed-up in Table 6.
Summarizing the results on the computational effort, our online update algorithm makes the approach practically viable for researchers and data scientists without access to specialized high-performance computation centers. Taking into account the trade-off between accuracy and computational effort is of practical importance in many industrial settings, where analysts work under time pressure and data arrives at high velocity. Furthermore, time saved in raw computation can be used for better data exploration and feature engineering.
5 Discussion and Conclusion
Distributional learning algorithms such as GAMLSS and deep distributional networks have been used successfully for probabilistic electricity price forecasting (PEPF, see e.g. Muniain and Ziel, 2020; Hirsch et al., 2024; Marcjasz et al., 2023). However, even for univariate distributions, these models are computationally expensive. At the same time, the literature on probabilistic electricity price forecasting has largely focused on modeling the hourly marginal distributions only, leaving the dependence structure neglected. Against this background, we develop an online estimation algorithm for multivariate distributional regression models, making the use of these algorithms feasible even for high-dimensional problems such as the 24-dimensional distribution of electricity prices on a standard laptop. We benchmark our implementation in a forecasting study for the German day-ahead electricity market and thereby provide the first study exclusively focused on online learning for multivariate PEPF.
Our results show that modeling the dependence structure in the day-ahead market improves probabilistic forecasting performance significantly. First, we see that calibration for prediction bands of the 24-dimensional price path is improved significantly by modeling the dependence structure and moving towards proper joint prediction bands (JPB, see Staszewska-Bystrova, 2011) instead of marginal, quantile-based prediction bands. The online, multivariate distributional regression models deliver strong predictive accuracy across a range of multivariate scoring rules. Additionally, we like to highlight the importance of regularization to avoid overfitting in a high-dimensional setting and conduct two experiments to validate the regularization of the scale matrix. Distributional regression models are interpretable and therefore allow us to discuss the economic interpretation of the time-varying dependence structure.
We analyze the trade-off between computation time and forecasting accuracy. Building an efficient frontier between accuracy and computation time we show that online learning, compared to repeated batch fitting, yields better results for given computation budget — with speed-ups of 2-3 orders of magnitude. Our algorithm is implemented in a fairly generic manner, allowing e.g. for different distributional assumptions and keeping a familiar, sklearn-like API to facilitate the usage by other researchers and data scientists (Pedregosa et al., 2011) and contributed the implementation to the ondil package (Hirsch et al., 2024). [^5] See: https://github.com/simon-hirsch/ondil Reproduction code for all experiments is available on GitHub. [^6] See: https://github.com/simon-hirsch/online-mv-distreg.
Our research opens multiple avenues for future work. First, further research on the driving forces of the dependence structure in the German electricity market is necessary to improve the forecasting performance and guide decision-making processes in electricity trading. Modeling the dependence structure in electricity markets is a rather open field and has implications beyond forecasting, concerning also risk and portfolio management and asset optimization (Peña et al., 2024; Löhndorf and Wozabal, 2023; Beykirch et al., 2022, 2024). From an algorithmic perspective, we note that while our algorithm is already quite fast, further improvements in the computation speed might be possible by using a CG-type scoring algorithm (Rigby and Stasinopoulos, 2005; Green, 1984; Cole and Green, 1992) and parallelizing over the elements of the distribution parameter. A further open issue is model selection - while the regularized online estimation is fast, the models are still quite complex and can be prone to overfitting. Lastly, due to the generic nature of our implementation, the usage for other high-dimensional forecasting problems such as probabilistic wind, solar, and load forecasting can be explored.
Acknowledgments
Simon Hirsch is employed as an industrial PhD student by Statkraft Trading GmbH and gratefully acknowledges the support and funding received. This work contains the author’s opinions and does not necessarily reflect Statkraft’s position. Simon Hirsch is grateful for helpful discussions with Florian Ziel, Daniel Gruhlke, Christoph Hanck and the participants of the IWSM 2025 in Limerick.
Declaration of Interest
The author declare that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Statement
The data used in this paper has been provided by Lipiecki et al. (2024) and cannot be shared further by the author. Therefore, for any requests with respect to the data, please contact Lipiecki et al. (2024) directly. A shorter data set of the same explanatory variables is available at the GitHub repository of Marcjasz et al. (2023). [^7] See: https://github.com/gmarcjasz/distributionalnn. For the differences between both data sets please consult Lipiecki et al. (2024), Section 2 and Section 5.1.3.
Generative AI Statement
During the preparation of this work the author used generative AI tools such as ChatGTP and GitHub Copilot in order to improve the quality of the language and of the code (esp. automatic PR reviews, documentation, creation of figures). After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication
Appendix A Appendix
A.1 Abbreviations
ACI Adaptive Conformal Prediction
AIC Akaike Information Criterion
APS Average Pinball Score
BIC Bayesian Information Criterion
CD Cholesky-Decomposition
CDF Cumulative Density Function
CP Conformal Prediction
CRPS Continuous Ranked Probability Score
DDNN Distributional Deep Neural Networks
DSS Dawid-Sebastiani Score
EPF Electricity Price Forecasting
ES Energy Score
GAMLSS Generalized Additive Models for Location, Scale and Shape
GARCH Generalized AutoRegressive Conditional Heteroscedasticity
GLM Generalized Linear Model
HQC Hannan-Quinn Criterion
IC Information Criterion
IRLS Iteratively Reweighted Least Squares
JPB Joint Prediction Band
LASSO Least Absolute Shrinkage and Selection Operator
LARX LASSO-estimated AutoRegressive Model with eXogenous variables
LRA Low-Rank Approximation
LS Log-Score (= negative log-likelihood)
MAE Mean Absolute Error
MCD Modified Cholesky-Decomposition
OCD Online Coordinate Descent
OLS Ordinary Least Squares
PDF Probability Density Function
PEPF Probabilistic Electricity Price Forecasting
PIT Probability Integral Transformation
RMSE Root Mean Squared Error
RS Rigby & Stasinopolous (Algorithm)
SCP Split Conformal Prediction
VS Variogram Score
Table 7: Abbreviations used in the Paper.
A.2 Derivation of Equation 10 and 11 for Newton-Raphson Scoring
We aim to calculate ${\partial\ell}/{\partial\eta}$ and ${\partial^{2}\ell}/{\partial\eta^{2}}$ for the calculation of the score and weight vectors (see Eq. 10 and Eq. 11) using the partial derivatives of the log-likelihood with respect to the distribution parameter (resp. the coordinate of the distribution parameter in case of matrix-valued parameters), ${\partial\ell}/{\partial\theta}$ and ${\partial^{2}\ell}/{\partial\theta^{2}}$. For continuous, twice differentiable link functions $\eta=g(\theta)$, we have
$\frac{\partial\ell}{\partial\eta}$ $=\frac{\partial\ell}{\partial\theta}\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-1}$ (31)
and
$\frac{\partial^{2}\ell}{\partial\eta^{2}}$ $=\frac{\partial\left(\dfrac{\partial\ell}{\partial\theta}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}\right)}{\partial\eta}=\frac{\partial\left(\dfrac{\partial\ell}{\partial\theta}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}\right)}{\partial\theta}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}$
and by the quotient rule, we have
$\frac{\partial^{2}\ell}{\partial\eta^{2}}=\left(\frac{\dfrac{\partial^{2}\ell}{\partial\theta^{2}}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)-\dfrac{\partial\ell}{\partial\theta}\left(\dfrac{\partial^{2}g(\theta)}{\partial\theta^{2}}\right)}{\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{2}}\left(\dfrac{\partial g(\theta)}{\partial\theta}\right)^{-1}\right)$ (32)
and the simplification
$\frac{\partial^{2}\ell}{\partial\eta^{2}}=\left(\frac{\partial^{2}\ell}{\partial\theta^{2}}\frac{\partial g(\theta)}{\partial\theta}-\frac{\partial\ell}{\partial\theta}\frac{\partial^{2}g(\theta)}{\partial\theta^{2}}\right)\left(\frac{\partial g(\theta)}{\partial\theta}\right)^{-3}$ (33)
concludes the derivation $\blacksquare$
A.3 Partial derivatives of the multivariate Gaussian Distribution
The probability density function of the multivariate normal distribution of dimension $D$ is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Sigma}})=\frac{1}{(2\pi)^{D/2}|\bm{\mathrm{\Sigma}}|^{1/2}}\exp\left(-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)$ (34)
with the location or mean vector $\bm{\mathrm{\mu}}$ and the scale respectively covariance matrix $\bm{\mathrm{\Sigma}}$. We parameterize the PDF in terms of the inverse scale matrix $\bm{\mathrm{\Omega}}=\bm{\mathrm{\Sigma}}^{-1}$:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}})=\frac{1}{(2\pi)^{D/2}}|\bm{\mathrm{\Omega}}|^{1/2}\exp\left(-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)$ (35)
and calculate the log-likelihood as $\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}})=\log(f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}}))$, which reads:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}})=-\frac{D}{2}\log(2\pi)-\frac{1}{2}\log(|\bm{\mathrm{\Omega}}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (36)
and parameterize the inverse covariance matrix through the Cholesky-decomposition $\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$, which yields:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1})=-\frac{D}{2}\log(2\pi)-\log(|\bm{\mathrm{A}}^{-1}|)-\frac{1}{2}\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$ (37)
where $\bm{\mathrm{z}}=\bm{\mathrm{A}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ and $z=\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$. The first derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{A}}^{-1}$ are given in Muschinski et al. (2022) and read
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Omega}}_{ik}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})$ (38)
$\frac{\partial\ell}{\partial(\bm{\mathrm{A}}^{-1})_{ij}}$ $=\frac{1}{(\bm{\mathrm{A}}^{-1})_{ij}}-(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{A}}^{-1})_{kj}$ (39)
and the second derivatives are given by
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\bm{\mathrm{\Omega}}_{ii}$ (40)
$\frac{\partial\ell^{2}}{\partial(\bm{\mathrm{A}}^{-1})_{ij}^{2}}$ $=-\frac{1}{(\bm{\mathrm{A}}^{-1})^{2}_{ij}}-(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (41)
For the modifed Cholesky-decomposition $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}\bm{\mathrm{D}}^{-1}\bm{\mathrm{L}}$ , the log-likelihood reads:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{D}}^{-1})=-\frac{D}{2}\log(2\pi)+\log(|\bm{\mathrm{D}}^{-1}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\left(\bm{\mathrm{L}}^{\top}\bm{\mathrm{D}}^{-1}\bm{\mathrm{L}}\right)(\bm{\mathrm{y}}-\bm{\mathrm{\mu}}).$ (42)
The first derivatives with respect to the elements of $\bm{\mathrm{\mu}}$, $\bm{\mathrm{L}}$ and $\bm{\mathrm{D}}^{-1}$ are given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Omega}}_{ik}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})$ (43)
$\frac{\partial\ell}{\partial\bm{\mathrm{L}}_{ij}}$ $=2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}$ (44)
$\frac{\partial\ell}{\partial(\bm{\mathrm{D}}^{-1}_{ii})}$ $=\frac{1}{2}\frac{1}{(\bm{\mathrm{D}}^{-1}_{ii})}+\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)^{2}=\frac{1}{2}\bm{\mathrm{D}}_{ii}+\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)^{2}$ (45)
and the second derivatives are given by:
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\bm{\mathrm{\Omega}}_{ii}$ (46)
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{L}}_{ij}^{2}}$ $=-\bm{\mathrm{D}}^{-2}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})^{2}$ (47)
$\frac{\partial\ell^{2}}{\partial(\bm{\mathrm{D}}^{-1})_{ii}^{2}}$ $=-\frac{1}{2}\frac{1}{(\bm{\mathrm{D}}^{-1}_{ii})^{2}}=-\frac{1}{2}\bm{\mathrm{D}}_{ii}^{2}$ (48)
keeping in mind that the inverse of a diagonal matrix is given by the inverse of the diagonal elements. For the low-rank approximation, we parameterize Equation 35 in terms of the low-rank approximation $\bm{\mathrm{\Omega}}=\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}$, which yields:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{U}},\bm{\mathrm{V}})=-\frac{D}{2}\log(2\pi)-\log(|(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})^{-1}|)-\frac{1}{2}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (49)
we note that the derivatives with respect to the elements of the mean vector $\bm{\mathrm{\mu}}_{i}$ remain the same:
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Omega}}_{ik}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})$ (50)
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\bm{\mathrm{\Omega}}_{ii}$ (51)
and the derivatives with respect to the elements of $\bm{\mathrm{D}}_{ii}$ are given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{1}{2}\left(\bm{\mathrm{\Sigma}}_{ii}-(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})^{2}\right)$ (52)
$\frac{\partial\ell^{2}}{\partial\bm{\mathrm{D}}_{ii}^{2}}$ $=-\frac{1}{2}\Sigma_{ii}^{2}.$ (53)
The partial derivatives with respect to the elements of $\bm{\mathrm{V}}$ are given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{V}}_{ij}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}$ (54)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{V}}_{ij}^{2}}$ $=\bm{\mathrm{\Sigma}}_{ij}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{ii}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{qk}\bm{\mathrm{V}}_{kj}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{iq}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}$ (55)
$-\left((\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}-\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}\right)^{2}\right)$
which concludes the derivation of the partial derivatives $\blacksquare$
A.4 Partial derivatives of the multivariate t-distribution
The probability density function (PDF) of the multivariate $t$-distribution of dimension $D$ is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\Sigma,\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}|\bm{\mathrm{\Sigma}}|^{1/2}}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Sigma}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}$
with the location vector $\mu$, the shape matrix $\bm{\mathrm{\Sigma}}$ and the degrees of freedom $\nu$. We parameterize the PDF in terms of the inverse shape matrix $\bm{\mathrm{\Omega}}=\bm{\mathrm{\Sigma}}^{-1}$:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{\Omega}},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|\bm{\mathrm{\Omega}}|^{1/2}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{\Omega}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}.$ (56)
We start with the partial derivatives for the CD-based parametrization. We have the Choleksy-decomposition $\bm{\mathrm{\Sigma}}=\bm{\mathrm{A}}\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{\Sigma}}^{-1}=\bm{\mathrm{\Omega}}=(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})$, which yields:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|(\bm{\mathrm{A}}^{-1})|\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{A}}^{-1})^{\top}(\bm{\mathrm{A}}^{-1})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}.$
Let us introduce some notation to simply the following derivatives. Define:
$\bm{\mathrm{z}}$ $=\bm{\mathrm{A}}^{-1}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})$ (57)
$z$ $=\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$ (58)
The log-likelihood is given by $\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\log(f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu))$ and reads:
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{A}}^{-1},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|(\bm{\mathrm{A}}^{-1})|\right)+\log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{z}}^{T}\bm{\mathrm{z}})\right)^{-(\nu+D)/2}\right)$ (59)
For the partial derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{A}}^{-1}$, we notice that $\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}}$ can be treated as a function of these elements and employ the chain rule. We see that:
$\frac{\partial(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})}{\partial\bm{\mathrm{\mu}}_{i}}$ $=2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})$ (60)
$\frac{\partial^{2}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-2\bm{\mathrm{\Omega}}_{ij}$ (61)
$\frac{\partial(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})}{\partial(\bm{\mathrm{A}}^{-1})_{ij}}$ $=2(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\sum_{m=1}^{M=j}(\bm{\mathrm{y}}_{m}-\bm{\mathrm{\mu}}_{m})(\bm{\mathrm{A}}^{-1})_{mj}$ (62)
$\frac{\partial(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{2}}{\partial(\bm{\mathrm{A}}^{-1})_{ij}^{2}}$ $=(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (63)
The chain rule for the last term of Equation 59 yields:
$\left[\log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})\right)^{-(\nu+D)/2}\right)\right]^{\prime}$ $=\frac{(D+\nu)}{2((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{\prime}$ (64)
$\left[\log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})\right)^{-(\nu+D)/2}\right)\right]^{\prime\prime}$ $=-\frac{(D+\nu)(((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)(\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{\prime\prime}-((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})^{\prime})^{2})}{2((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)^{2}}$ (65)
and plugging in the according partial derivatives in Equations 60 to 63 and applying integration by parts for the remainder of Equation 59, we have:
$\frac{\partial l}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)$ (66)
$\frac{\partial^{2}}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)(-2\bm{\mathrm{\Omega}}_{ij})-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2(z+\nu)^{2}}$ (67)
$\frac{\partial\ell}{\partial(\bm{\mathrm{A}}^{-1})_{ij}}$ $=\frac{1}{(\bm{\mathrm{A}}^{-1})_{ij}}\bm{1}_{i=j}+\frac{(D+\nu)}{2(z+\nu)}\left(2(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\sum_{m=1}^{M=i}(\bm{\mathrm{y}}_{m}-\bm{\mathrm{\mu}}_{m})(\bm{\mathrm{A}}^{-1})_{mj}\right)$ (68)
$\frac{\partial^{2}l}{\partial(\bm{\mathrm{A}}^{-1}){ij}^{2}}$ $=-\frac{1}{(\bm{\mathrm{A}}^{-1})_{ij}^{2}}\bm{1}_{i=j}-\frac{(D+\nu)\left((z+\nu)(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2((\bm{\mathrm{z}}^{\top}\bm{\mathrm{z}})+\nu)^{2}}$ (69)
Where $\bm{1}$ is the indicator function for $i=j$, since the partial derivative of $\log\left(|\bm{\mathrm{A}}^{-1}|\right)$ are only relevant for the partial derivatives of the diagonal elements of $\bm{\mathrm{A}}^{-1}$. For the partial derivatives with respect to the degrees of freedom $\nu$, integration by parts yields:
$\frac{\partial l}{\partial\nu}$ $=-\frac{-\nu\operatorname{digamma(\frac{D+\nu}{2})}+D+\nu\operatorname{digamma}(\frac{\nu}{2})}{2\nu}+\frac{1}{2}\left(\frac{z(D+\nu)}{\nu(\nu+z)}-\log\left(\frac{(\nu+z)}{\nu}\right)\right)$ (70)
$\frac{\partial^{2}l}{\partial\nu^{2}}$ $=\frac{1}{4}\left(\frac{2k}{\nu^{2}}+\operatorname{trigamma(\frac{D+\nu}{2})}-\operatorname{trigamma}(\frac{\nu}{2})\right)+\frac{z(\nu z-D(2\nu+z))}{2\nu^{2}(\nu+z)^{2}}.$ (71)
For the MCD, we parameterize $\bm{\mathrm{\Omega}}=\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}$ and hence the PDF is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{D}}^{-1},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|\bm{\mathrm{D}}^{-1}|^{1/2}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}$ (72)
and the log-likelihood is given by
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{L}},\bm{\mathrm{D}}^{-1},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|\bm{\mathrm{D}}^{-1}|\right)+\\ \log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}\right)$ (73)
we follow a similar strategy as above and see that the partial derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and with respect to the degrees of freedom $\nu$ are the same as above:
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)$ (74)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)(-2\bm{\mathrm{\Omega}}_{ij})-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2(z+\nu)^{2}}$ (75)
$\frac{\partial l}{\partial\nu}$ $=-\frac{-\nu\operatorname{digamma(\frac{D+\nu}{2})}+D+\nu\operatorname{digamma}(\frac{\nu}{2})}{2\nu}+\frac{1}{2}\left(\frac{z(D+\nu)}{\nu(\nu+z)}-\log\left(\frac{(\nu+z)}{\nu}\right)\right)$ (76)
$\frac{\partial^{2}l}{\partial\nu^{2}}$ $=\frac{1}{4}\left(\frac{2k}{\nu^{2}}+\operatorname{trigamma(\frac{D+\nu}{2})}-\operatorname{trigamma}(\frac{\nu}{2})\right)+\frac{z(\nu z-D(2\nu+z))}{2\nu^{2}(\nu+z)^{2}}.$ (77)
where $z$ is now defined as
$z=(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}\bm{\mathrm{L}}^{\top}(\bm{\mathrm{D}}^{-1})\bm{\mathrm{L}}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}}).$ (78)
For the partial derivatives with respect to the elements of $\bm{\mathrm{D}}^{-1}$, we note that the partial derivatives of the second term are given by:
$\frac{\partial}{\partial\bm{\mathrm{D}}^{-1}_{ii}}$ $=-\frac{1}{2}\bm{\mathrm{D}}_{ii}$ (79)
$\frac{\partial}{\partial(\bm{\mathrm{D}}^{-1}_{ii})^{2}}$ $=\frac{1}{2}\bm{\mathrm{D}}_{ii}^{2}$ (80)
and the partial deriviatives of the third term are given by
$\frac{\partial z}{\partial\bm{\mathrm{D}}^{-1}_{ii}}$ $=\left(\sum_{j=1}^{D}\bm{\mathrm{L}}_{ji}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}$ (81)
$\frac{\partial z}{\partial(\bm{\mathrm{D}}^{-1}_{ii})^{2}}$ $=0$ (82)
$\frac{\partial z}{\partial\bm{\mathrm{L}}_{ij}}$ $=2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}$ (83)
$\frac{\partial z}{\partial(\bm{\mathrm{L}}_{ij})^{2}}$ $=2\bm{\mathrm{D}}_{ii}^{-1}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})^{2}$ (84)
Plugging these results in Equations 64 and 65, we have for the partial derivatives of the log-likelihood with respect to the elements of $\bm{\mathrm{D}}$:
$\frac{\partial l}{\partial\bm{\mathrm{D}}^{-1}_{ii}}$ $=-\frac{1}{2}\bm{\mathrm{D}}_{ii}+\frac{(D+\nu)}{2(z+\nu)}\left(\left(\sum_{j=1}^{D}\bm{\mathrm{L}}_{ji}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)$ (85)
$\frac{\partial^{2}l}{\partial(\bm{\mathrm{D}}^{-1}_{ii})^{2}}$ $=\frac{1}{2}\bm{\mathrm{D}}_{ii}^{2}-\frac{(D+\nu)\left(-\left(\left(\sum_{j=1}^{D}\bm{\mathrm{L}}_{ji}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)^{2}\right)}{2(z+\nu)^{2}}$ (86)
and with respect to the elements of $\bm{\mathrm{L}}$:
$\frac{\partial l}{\partial\bm{\mathrm{L}}_{ij}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)$ (87)
$\frac{\partial^{2}l}{\partial(\bm{\mathrm{L}}_{ij})^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)2\bm{\mathrm{D}}_{ii}^{-1}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})^{2}\left(2\bm{\mathrm{D}}^{-1}_{ii}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})\bm{\mathrm{L}}_{ki}\right)^{2}\right)}{2(z+\nu)^{2}}.$ (88)
For the low-rank approximation, we follow a similar notation. The LRA is given by $\bm{\mathrm{\Omega}}=\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}$ and hence the PDF is given by:
$f(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{D}},\bm{\mathrm{V}},\nu)=\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}|\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}}|^{1/2}\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}$ (89)
and the log-likelihood is given by
$\ell(\bm{\mathrm{y}}\mid\bm{\mathrm{\mu}},\bm{\mathrm{D}},\bm{\mathrm{V}},\nu)=\log\left(\frac{\Gamma((\nu+D)/2)}{\Gamma(\nu/2)\nu^{D/2}\pi^{D/2}}\right)+\log\left(|(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})|\right)+\\ \log\left(\left(1+\frac{1}{\nu}(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})\right)^{-(\nu+D)/2}\right)$ (90)
we follow a similar strategy as above and see that the partial derivatives with respect to the elements of $\bm{\mathrm{\mu}}$ and with respect to the degrees of freedom $\nu$ are the same as above:
$\frac{\partial\ell}{\partial\bm{\mathrm{\mu}}_{i}}$ $=\frac{(D+\nu)}{2(z+\nu)}\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)$ (91)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{\mu}}_{i}^{2}}$ $=-\frac{(D+\nu)\left((z+\nu)(-2\bm{\mathrm{\Omega}}_{ij})-\left(2\sum_{j=1}^{D}\bm{\mathrm{\Omega}}_{ij}(\bm{\mathrm{y}}_{j}-\bm{\mathrm{\mu}}_{j})\right)^{2}\right)}{2(z+\nu)^{2}}$ (92)
$\frac{\partial l}{\partial\nu}$ $=-\frac{-\nu\operatorname{digamma(\frac{D+\nu}{2})}+D+\nu\operatorname{digamma}(\frac{\nu}{2})}{2\nu}+\frac{1}{2}\left(\frac{z(D+\nu)}{\nu(\nu+z)}-\log\left(\frac{(\nu+z)}{\nu}\right)\right)$ (93)
$\frac{\partial^{2}l}{\partial\nu^{2}}$ $=\frac{1}{4}\left(\frac{2k}{\nu^{2}}+\operatorname{trigamma(\frac{D+\nu}{2})}-\operatorname{trigamma}(\frac{\nu}{2})\right)+\frac{z(\nu z-D(2\nu+z))}{2\nu^{2}(\nu+z)^{2}}.$ (94)
where $z$ is now defined as
$z=(\bm{\mathrm{y}}-\bm{\mathrm{\mu}})^{\top}(\bm{\mathrm{D}}+\bm{\mathrm{V}}^{\top}\bm{\mathrm{V}})(\bm{\mathrm{y}}-\bm{\mathrm{\mu}}).$ (95)
For the partial derivatives with respect to the elements of $\bm{\mathrm{D}}$, we note that the partial derivatives of the second term are given by:
$\frac{\partial}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{1}{2}\bm{\mathrm{\Sigma}}_{ii}$ (96)
$\frac{\partial}{\partial(\bm{\mathrm{D}}_{ii})^{2}}$ $=-\frac{1}{2}\left(\bm{\mathrm{\Sigma}}_{ii}\right)^{2}$ (97)
and the partial deriviatives of the third term are given by
$\frac{\partial}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{D+\nu}{2(z+\nu)}(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (98)
$\frac{\partial}{\partial\bm{\mathrm{D}}_{ii}^{2}}$ $=0$ (99)
and the second defaults to 0. The partial derivatives of the log-likelihood with respect to the elements of $\bm{\mathrm{D}}$ are hence given by:
$\frac{\partial\ell}{\partial\bm{\mathrm{D}}_{ii}}$ $=\frac{1}{2}\bm{\mathrm{\Sigma}}_{ii}-\frac{D+\nu}{2(z+\nu)}(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (100)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{D}}_{ii}^{2}}$ $=-\frac{1}{2}\left(\bm{\mathrm{\Sigma}}_{ii}\right)^{2}-\frac{D+\nu}{2(z+\nu)^{2}}\left((\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}\right)^{2}.$ (101)
For the partial derivatives with respect to the elements of $\bm{\mathrm{V}}$, we have a more complex formulation for the second term involving the determinant:
$\frac{\partial}{\partial V_{ij}}$ $=$ (102)
$\frac{\partial}{\partial V_{ij}^{2}}$ $=\bm{\mathrm{\Sigma}}_{ij}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{ii}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{qk}\bm{\mathrm{V}}_{kj}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{iq}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}$ (103)
and the partial derivatives of the third term are given by
$\frac{\partial}{\partial V_{ij}}$ $=\frac{D+\nu}{(z+\nu)}\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}$ (104)
$\frac{\partial}{\partial V_{ij}^{2}}$ $=\frac{D+\nu}{(z+\nu)^{2}}(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}$ (105)
and hence integration by parts again gives us, similiar to the partial derivatives for the multivariate normal distribution in Equation 54 and 55:
$\frac{\partial\ell}{\partial\bm{\mathrm{V}}_{ij}}$ $=\sum_{k=0}^{D}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}+\frac{D+\nu}{(z+\nu)}\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}$ (106)
$\frac{\partial^{2}\ell}{\partial\bm{\mathrm{V}}_{ij}^{2}}$ $=\bm{\mathrm{\Sigma}}_{ij}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{ii}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{qk}\bm{\mathrm{V}}_{kj}-\sum_{k=0}^{D}\sum_{q=0}^{D}\bm{\mathrm{\Sigma}}_{iq}\bm{\mathrm{V}}_{qj}\bm{\mathrm{\Sigma}}_{ik}\bm{\mathrm{V}}_{kj}$ (107)
$-\frac{D+\nu}{(z+\nu)^{2}}\left((\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})^{2}-\left(\sum_{k=0}^{D}(\bm{\mathrm{y}}_{k}-\bm{\mathrm{\mu}}_{k})(\bm{\mathrm{y}}_{i}-\bm{\mathrm{\mu}}_{i})\bm{\mathrm{V}}_{kj}\right)^{2}\right)$
which concludes the derivation of the partial derivatives for the multivariate $t$-distribution $\blacksquare$
A.5 Hyperparameters for the Multivariate Distributional Regression Model
To align with the reproducible research best practices, as described in e.g Lago et al. (2021), we publish the reproduction code on GitHub at https://github.com/simon-hirsch/online-mv-distreg, allowing for full reproducibility of all experiments. Additionally, we take the following paragraph to describe the hyperparameters of the model:
  • 1.
    Information criteria and model selection: We use the BIC for the multivariate distributional regression model and run the online coordinate descent on an exponential grid of 100 $\lambda$ values. We employ fast model selection based on the first derivatives for the CD-based models.
  • 2.
    Link functions: We use the identity link for the location for all models. For the Cholesky-based distributional models, we use the log-link. For the LRA-based models, we employ the square root link for the diagonal matrix $\bm{\mathrm{A}}$ as initial experiments showed a more robust convergence behavior and the identity for the matrix $\bm{\mathrm{V}}$. For the degrees of freedom $\nu$, we employ an inverse softplus shifted to 2.1, which ensures that $\nu>2$ and hence the covariance matrix is positive definite. We have found the shift to be important to avoid numerical instabilities for $\nu$ close to 2. These links also apply to the univariate distributional regression models.
  • 3.
    Early stopping: We employ early stopping for the path-based regularization of the scale matrix if the AIC does not improve, as described in Section 2.6. We limit the number of off-diagonals for the CD-based parameterization to max 6, however note that the algorithm breaks after fitting 1-2 off-diagonals. We do not limit the number of columns fitted in the LRA-based model and note that the algorithm breaks after fitting the full rank-2 matrix $\bm{\mathrm{V}}$.
  • 4.
    Number of iterations, step-size and dampening: We dampen the estimation in the first iteration for the scale parameters only. We generally allow for a maximum of 30 inner and 10 outer iterations in the initial fit and the update steps.